### 11 Mar 2024 v1

# Representing rational integers by generalized quadratic forms over quadratic fields

Abstract.
We investigate generalized quadratic forms with values in the set of rational integers over quadratic fields. We characterize the real quadratic fields which admit a positive definite binary generalized form of this type representing every positive integer. We also show that there are only finitely many such fields where a ternary generalized form with these properties exists.
Key words and phrases:
Generalized quadratic form, universal quadratic form, Hermitian form, real quadratic field
2020 Mathematics Subject Classification:
11E12, 11E16, 11E20, 11E25, 11E39, 11R11, 11R80
We acknowledge support by Czech Science Foundation (GAČR) grant 21-00420M and the Charles University Research Centre program UNCE/24/SCI/022.
1. Introduction
A positive definite quadratic form over $\mathbb{Z}$ is called universal if it represents every positive integer $n$. The history of universal quadratic forms begins with Lagrange’s proof of the Four Square Theorem in 1770: Every positive integer $n$ is of the form $x^{2}+y^{2}+z^{2}+w^{2}$. Diagonal quaternary universal forms were classified by Ramanujan [Ra], who stated that there exist $55$ such forms. Dickson [Di] corrected Ramanujan’s statement by pointing out that the form $x^{2}+2y^{2}+5z^{2}+5u^{2}$ does not represent $15$. Thus there exist exactly $54$ diagonal universal forms. Dickson [Di2] also proved various results on representation by non-diagonal quaternary forms.
The 90’s saw a proof of the celebrated Conway-Schneeberger 15-Theorem. A quadratic form having an integer matrix is called classical.
Theorem 1.1 (15-Theorem, Conway-Schneeberger).
A classical quadratic form $Q$ over $\mathbb{Z}$ is universal if and only if it represents the integers
[…]
The proof was simplified by Bhargava [Bh]. As a corollary, one obtains all classical quaternary forms – there are exactly $207$ of them. Bhargava and Hanke [BH] extended the result to non-classical forms by proving the 290-Theorem: If a positive definite quadratic form $Q$ over $\mathbb{Z}$ represents every positive integer $n\leq 290$, then $Q$ is universal. In fact, for $Q$ to be universal, it is enough that it represents every integer from a list of $29$ numbers, the largest number on the list being $290$. A corollary is that there are exactly $6436$ universal quaternary forms [BH, Theorem 4].
It is natural to extend the notion of universality from $\mathbb{Z}$ to the ring of integers $\mathcal{O}_{K}$ in a number field. Assume that the number field $K$ is totally real. A positive integral quadratic form over $K$ is called universal if it represents all totally positive integers in $K$. A version of the local-global principal over number fields proved by Hsia, Kitaoka, and Kneser [HKK] implies that a universal form over $K$ always exists. Maaß [Ma] showed that the sum of three squares $x^{2}+y^{2}+z^{2}$ is universal over $\mathbb{Q}(\sqrt{5})$. By a result of Siegel [Si], if a sum of squares is universal over $K$, then $K=\mathbb{Q}$ or $\mathbb{Q}(\sqrt{5})$. In 1993, Chan, Kim, and Raghavan [CKR] proved that a ternary universal quadratic form exists over a real quadratic field $K=\mathbb{Q}(\sqrt{D})$ if and only if $D$ equals $2$, $3$, or $5$. Moreover, they identified all ternary universal forms over these fields up to equivalence. For forms of higher ranks, B. M. Kim [Ki] proved that there exist only finitely many $D$’s such that $\mathbb{Q}(\sqrt{D})$ admits a universal form in $7$ variables (see also [KKP2] for an explicit estimate on the size of $D$). On the other hand, a $8$-ary universal form exists over each $\mathbb{Q}(\sqrt{n^{2}-1})$ (when $n^{2}-1$ is squarefree) [Ki2]. Blomer and Kala [BK] showed: For any positive integer $r$, there are infinitely many quadratic fields $\mathbb{Q}(\sqrt{D})$ that do not have a universal lattice of rank $\leq r$. Lower bounds for the rank of a universal form depending on $D$ were obtained by Kala and Tinková [KT].
One of the principal open problems in this area is Kitaoka’s conjecture: There are only finitely many totally real number fields $K$ having a ternary universal form. The conjecture was proved by Kala and Yatsyna [KY] for fields of an arbitrary but fixed degree $d$.
For a more comprehensive account of the theory of universal quadratic forms, especially over number fields, we refer the reader to the survey article [Ka2].
In the same vein, one can study Hermitian forms over imaginary quadratic fields. We call a positive definite Hermitian form over an imaginary quadratic field $\mathbb{Z}$-universal if it represents every $n\in\mathbb{Z}_{\geq 1}$. Earnest and Khosravani [EK] proved that universal binary Hermitian forms exist over $K$ for only finitely many $m$, and identified all such forms in the case when $K$ has class number $1$. B. M. Kim, J. Y. Kim, and P.-S. Park [KKP] proved an analogy of the 15-Theorem for Hermitian quadratic forms: If a positive definite Hermitian form $H$ represents $1$, $2$, $3$, $5$, $6$, $7$, $10$, $13$, $14$, and $15$, then $H$ is universal. They also completely determined the minimal rank of a positive Hermitian form over all imaginary quadratic fields.
The object of study of our paper are generalized quadratic forms introduced recently by Browning, Pierce, and Schindler [BPS, Definition 1.1]. They can be defined for an arbitrary Galois extension $K$ of $\mathbb{Q}$ but we will consider them only over quadratic fields. A quadratic field $K=\mathbb{Q}(\sqrt{D})$ has two $\mathbb{Q}$-automorphisms: the identity and the conjugation map
$\tau:K\hookrightarrow K,a+b\sqrt{D}\mapsto a-b\sqrt{D}.$
Informally speaking, a generalized quadratic form may contain also the conjugates of the variables. For example,
(1) $G(z,w)=z^{2}-z\tau(z)+\tau(z)^{2}+w^{2}+w\tau(w)+\tau(w)^{2}$
is a generalized quadratic form (and we will show later that it represents every $a\in\mathbb{Z}_{\geq 1}$ over $\mathbb{Q}(\sqrt{2})$). Thus, the concept generalizes both quadratic and Hermitian forms.
A generalized quadratic form is a special case of a generalized polynomial (see our Definition 3.1). The values of a generalized polynomial $g$ in $n$ variables with coefficients in $K$ also lie in $K$. However, we restrict our attention to generalized polynomials with values in $\mathbb{Q}$ and find a minimal set of generators for the ring of such polynomials in Theorem 3.4.
If $G(z_{1},z_{2},\dots,z_{n})$ is a generalized form in $n$ variables and
$G(\alpha_{1},\alpha_{2},\dots,\alpha_{n})\in\mathbb{Z},\quad\alpha_{1},\alpha_{2},\dots,\alpha_{n}\in\mathcal{O}_{K},$
then we say that $G$ is $\mathbb{Z}$-valued. The properties of being positive definite, integral, and classical also readily extend to this setting. A positive definite integral $\mathbb{Z}$-valued generalized form which represents every $a\in\mathbb{Z}_{\geq 1}$ will be called $\mathbb{Z}$-universal. Our main theorem is the following.
Theorem 1.2.
Let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}_{\geq 2}$ is squarefree.
  1. i)
    Assume that $D\equiv 2,3\pmod{4}$. A binary $\mathbb{Z}$-universal generalized quadratic form exists over $K$ if and only if $D\in\{2,3,6,7,10\}$.
  2. ii)
    Assume that $D\equiv 1\pmod{4}$. A classical binary $\mathbb{Z}$-universal generalized quadratic form exists over $K$ if and only if $D=5$.
We further extend this result to ternary generalized forms by proving the following.
Theorem 1.3.
Let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}_{\geq 2}$ is squarefree. Assume that one of the following two conditions is satisfied:
  1. i)
    $D\equiv 2,3\pmod{4}$ and there exists a ternary $\mathbb{Z}$-universal generalized quadratic form over $K$,
  2. ii)
    $D\equiv 1\pmod{4}$ and there exists a classical ternary $\mathbb{Z}$-universal generalized quadratic form over $K$.
Then $D\leq 110$.
The proofs heavily rely on the 15-Theorem. A natural question is what happens when we remove the assumption that the form is classical in the case $D\equiv 1\pmod{4}$. By an application of the 290-Theorem, we can find the answer.
Theorem 1.4.
There are only finitely many real quadratic fields which admit a binary or ternary $\mathbb{Z}$-universal generalized quadratic form.
We will prove these as Theorems 5.10, 5.11 and 5.12 in Section 5. In contrast with the situation for binary and ternary forms, it is not very difficult to construct a generalized form in four variables which is $\mathbb{Z}$-universal over every real quadratic field (and we do this later in Proposition 5.13). This provides a fairly complete picture of the minimal rank required for a generalized form to be $\mathbb{Z}$-universal.
The preceding discussion relates only to forms which are positive definite. As regards indefinite forms, we prove the following in Theorem 6.6: If $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}_{\geq 2}$ is squarefree, then the binary generalized form
$G(z,w)=z\tau(z)+w\tau(w)$
represents every $a\in\mathbb{Z}$.
If we start with an integral $\mathbb{Z}$-valued generalized form and express the variables in the integral basis, then we end up with an integral quadratic form over $\mathbb{Z}$ in twice as many variables (this will be proved formally later). For example, if we let $z=x_{1}+y_{1}\sqrt{2}$ and $w=x_{2}+y_{2}\sqrt{2}$ in (1), then we obtain the quadratic form
$Q(x_{1},y_{1},x_{2},y_{2})=x_{1}^{2}+6y_{1}^{2}+3x_{2}^{2}+2y_{2}^{2}.$
(It can be checked that $Q$ is universal by the 15-Theorem.) We call $Q$ the quadratic form associated to $G$. A natural question is which universal quaternary quadratic forms arise in this way. In other words, we ask which quaternary quadratic forms are associated to some $\mathbb{Z}$-universal binary generalized form. We find all such forms (up to equivalence) for $D\in\{2,3,6,7,10\}$. We also find all quadratic forms associated to classical $\mathbb{Z}$-universal binary generalized forms (again, up to equivalence) for $D=5$.
The rest of the paper is organized as follows. In Section 2, we collect the preliminaries about quadratic forms and quadratic fields. In Section 3, we find a minimal set of generators of the ring of $\mathbb{Q}$-valued generalized polynomials. The theory of generalized quadratic forms is developed in Section 4. We prove Theorems 1.2, 1.3 and 1.4 at the end of Section 5 and Theorem 6.6 in Section 6. We find all quaternary forms associated to $\mathbb{Z}$-universal binary generalized forms for $D\in\{2,3,6,7,10\}$ (and to classical ones for $D=5$) in Section 7.
Acknowledgments
This research project was accomplished at the Student Number Theory Seminar at Charles University. We thank Vítězslav Kala for organizing the seminar. We also thank Nicolas Daans for suggesting the proof of Proposition 6.7 and Pavlo Yatsyna for his recommendation of the reference [Ca].
2. Preliminaries
Let $K$ be a number field and let $\mathcal{O}_{K}$ denote the ring of integers of $K$. Consider a polynomial $p$ in $n$ variables over $K$, i.e.,
$p(x_{1},x_{2},\dots,x_{n})=\sum_{\begin{subarray}{c}i_{1},\dots,i_{n}\geq 0\\ i_{1}+\cdots+i_{n}\leq k\end{subarray}}a_{i_{1},i_{2},\dots,i_{n}}x_{1}^{i_{1}}x_{2}^{i_{2}}\cdots x_{n}^{i_{n}}$
where $k\in\mathbb{Z}_{\geq 0}$ and $a_{i_{1},i_{2},\dots,i_{n}}\in K$ for $i_{1},i_{2},\dots,i_{n}\in\{0,1,\dots,k\}$. We will make use of the multi-index notation. If $\mathbf{i}=(i_{1},i_{2},\dots,i_{n})\in\mathbb{Z}_{\geq 0}^{n}$ is a multi-index, then we let $|\mathbf{i}|=\sum_{r=1}^{n}i_{r}$ be the sum of its components, and if $\mathbf{x}=(x_{1},x_{2},\dots,x_{n})$, then we let $\mathbf{x}^{\mathbf{i}}=x_{1}^{i_{1}}x_{2}^{i_{2}}\cdots x_{n}^{i_{n}}$. The polynomial $p$ can be succinctly written as
$p(\mathbf{x})=\sum_{|\mathbf{i}|\leq k}a_{\mathbf{i}}\mathbf{x}^{\mathbf{i}}.$
The lexicographic order on $\mathbb{Z}_{\geq 0}^{n}$ is defined as follows: Let $\mathbf{i},\mathbf{j}\in\mathbb{Z}_{\geq 0}^{n}$, $\mathbf{i}=(i_{1},i_{2},\dots,i_{n})$ and $\mathbf{j}=(j_{1},j_{2},\dots,j_{n})$. Then $\mathbf{i}<\mathbf{j}$ if and only if there exists $s\in\{1,2,\dots,n\}$ such that $i_{r}=j_{r}$ for $1\leq r<s$ and $i_{s}<j_{s}$.
A quadratic form $Q$ in $n$ variables over $K$ is a homogeneous polynomial of degree $2$, i.e.,
$Q(x_{1},x_{2},\dots,x_{n})=\sum_{1\leq i\leq j\leq n}a_{ij}x_{i}x_{j}$
where $a_{ij}\in K$. The matrix $M_{Q}$ of the form $Q$ is defined as
$M_{Q}=\begin{pmatrix}a_{11}&\frac{a_{12}}{2}&\dots&\frac{a_{1n}}{2}\\ \frac{a_{12}}{2}&a_{22}&\dots&\frac{a_{2n}}{2}\\ \vdots&\vdots&\ddots&\vdots\\ \frac{a_{1n}}{2}&\frac{a_{2n}}{2}&\dots&a_{nn}\end{pmatrix}.$
The rank of $Q$ is the rank of the matrix $M_{Q}$ over $K$. If the coefficients $a_{ij}$ belong to $\mathcal{O}_{K}$, then $Q$ is called integral. If the matrix $M_{Q}$ has entries in $\mathcal{O}_{K}$, i.e., $a_{ij}\in 2\mathcal{O}_{K}$ for $i\neq j$, then $Q$ is called classical.
Let $Q$ be a quadratic form over $\mathbb{Q}$. We say that $Q$ represents $b\in\mathbb{Z}$ over $\mathbb{Z}$ if there exists $\mathbf{a}\in\mathbb{Z}^{n}$ such that $Q(\mathbf{a})=b$. In particular, according to our definition, every quadratic form represents 00. If there exists $\mathbf{a}\in\mathbb{Z}^{n}\setminus\{(0,0,\dots,0)\}$ such that $Q(\mathbf{a})=0$, then we say that $Q$ represents 00 non-trivially. The form $Q$ is called positive definite if $M_{Q}$ is a positive definite matrix and $Q$ is called universal if it is a positive definite integral form such that it represents every $b\in\mathbb{Z}_{\geq 1}$.
In the rest of the paper, we consider quadratic forms over $K$ only when $K=\mathbb{Q}$ or $K$ is a quadratic field. Let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}\setminus\{0,1\}$ is squarefree. If we let
$\omega_{D}=\begin{cases}\sqrt{D},&\text{if }D\equiv 2,3\pmod{4},\\ \frac{1+\sqrt{D}}{2},&\text{if }D\equiv 1\pmod{4},\end{cases}$
then $(1,\omega_{D})$ is a basis of $\mathcal{O}_{K}$. The quadratic field $K$ has two embeddings into $\mathbb{C}$: the identity and the conjugation map
$\tau:K\hookrightarrow\mathbb{C},a+b\sqrt{D}\mapsto a-b\sqrt{D}.$
For $\alpha\in K$, we let $\operatorname{Tr}(\alpha)=\alpha+\tau(\alpha)$ and $\operatorname{N}(\alpha)=\alpha\tau(\alpha)$ be the trace and norm of $\alpha$, respectively. More generally, if $K$ is a number field with $r$ real and $2s$ complex embeddings
$\sigma_{1},\sigma_{2},\dots,\sigma_{r}:K\hookrightarrow\mathbb{R},\sigma_{r+1},\bar{\sigma}_{r+1},\dots,\sigma_{r+s},\bar{\sigma}_{r+s}:K\hookrightarrow\mathbb{C},$
then we let $V=\mathbb{R}^{r}\times\mathbb{C}^{s}$ and call $V$ the Minkowski space. The field $K$ embeds into $V$ via the map
$\sigma:K\hookrightarrow V,\alpha\mapsto(\sigma_{1}(\alpha),\dots,\sigma_{r+s}(\alpha)).$
We have $K\otimes_{\mathbb{Q}}\mathbb{R}\simeq V$ (see [Mi, p. 70]). In particular, $\sigma(K)$ is dense in $V$.
3. Generalized polynomials with rational values
We begin by defining generalized polynomials over a quadratic field.
Definition 3.1.
Let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}\setminus\{-1,0\}$ is squarefree. A generalized polynomial in $n$ variables over $K$ is an expression of the form
(2) $g(z_{1},z_{2},\dots,z_{n})=\sum_{\begin{subarray}{c}i_{1},j_{1},\dots,i_{n},j_{n}\geq 1\\ i_{1}+j_{1}+\cdots+i_{n}+j_{n}\leq k\end{subarray}}\alpha_{i_{1},j_{1},\dots i_{n},j_{n}}z_{1}^{i_{1}}\tau(z_{1})^{j_{1}}\cdots z_{n}^{i_{n}}\tau(z_{n})^{j_{n}}$
where $k\in\mathbb{Z}_{\geq 0}$ and $\alpha_{i_{1},j_{1},\dots,i_{n},j_{n}}\in K$ for every $i_{1},j_{1},\dots,i_{n},j_{n}\in\{0,1,\dots,k\}$.
If $\mathbf{z}=(z_{1},z_{2},\dots,z_{n})$, then we let $\tau(\mathbf{z})=(\tau(z_{1}),\tau(z_{2}),\dots,\tau(z_{n}))$. Using the multi-index notation, we can write $g$ as
$g(\mathbf{z})=\sum_{|\mathbf{i}|+|\mathbf{j}|\leq k}\alpha_{\mathbf{i}\mathbf{j}}\mathbf{z}^{\mathbf{i}}\tau(\mathbf{z})^{\mathbf{j}}.$
If $g$ is a generalized polynomial of the form (2), then we define the degree of $g$ as
$\deg(g)=\max\{|\mathbf{i}|+|\mathbf{j}|:\;\alpha_{\mathbf{i}\mathbf{j}}\neq 0\}.$
We see that $g$ is a generalized polynomial in $n$ variables if and only if there exists a polynomial $\tilde{g}$ in $2n$ variables such that
$g(z_{1},z_{2},\dots,z_{n})=\tilde{g}(z_{1},\tau(z_{1}),\dots,z_{n},\tau(z_{n})).$
We are particularly interested in generalized polynomials $g$ with values in $\mathbb{Q}$. They are characterized by the property $g(\mathbf{a})=\tau(g(\mathbf{a}))$ for every $\mathbf{a}\in K^{n}$. We prove below that the coefficients of the two conjugate terms $\mathbf{z}^{\mathbf{i}}\tau(\mathbf{z})^{\mathbf{j}}$ and $\tau(\mathbf{z})^{\mathbf{j}}\mathbf{z}^{\mathbf{i}}$ in such a polynomial must be conjugate, i.e., $\alpha_{\mathbf{j}\mathbf{i}}=\tau(\alpha_{\mathbf{i}\mathbf{j}})$.
Lemma 3.2.
Let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}\setminus\{0,1\}$ is squarefree and let $g$ be a generalized polynomial in $n$ variables over $K$. If $g(\mathbf{a})=0$ for every $\mathbf{a}\in K^{n}$, then $g$ is the zero polynomial.
Proof.
Let $\tilde{g}$ be a polynomial in $2n$ variables such that
$g(z_{1},z_{2},\dots,z_{n})=\tilde{g}(z_{1},\tau(z_{1}),\dots,z_{n},\tau(z_{n})).$
First, suppose that $K$ is real ($r=2$ and $s=0$). We know that $\sigma(K)=\{(\alpha,\tau(\alpha)):\;\alpha\in K\}$ is dense in $V=\mathbb{R}^{2}$. Consequently, $\sigma(K)^{n}$ is dense in $\mathbb{R}^{2n}$. Because the polynomial $\tilde{g}$ vanishes on $\sigma(K)^{n}$, it vanishes on the whole set $\mathbb{R}^{2n}$, and hence $\tilde{g}$ is the zero polynomial.
Secondly, suppose that $K$ is imaginary ($r=0$ and $s=1$). In this case, $\sigma(K)=K$ is dense in $V=\mathbb{C}$. Thus $K^{n}$ is dense in $\mathbb{C}^{n}$ and the polynomial $\tilde{g}$ is zero on the set
$\{(z_{1},\overline{z_{1}},\dots,z_{n},\overline{z_{n}}):\;z_{1},\dots,z_{n}\in K\}.$
Let $h$ be another polynomial in $2n$ variables defined by
$h(x_{1},y_{1},\dots,x_{n},y_{n})=\tilde{g}(x_{1}+iy_{1},x_{1}-iy_{1},\dots,x_{n}+iy_{n},x_{n}-iy_{n}).$
Since $h$ vanishes on $\mathbb{R}^{2n}$ and
$g(z_{1},z_{2},\dots,z_{n})=h\left(\frac{z_{1}+\overline{z_{1}}}{2},\frac{z_{1}-\overline{z_{1}}}{2i},\dots,\frac{z_{n}+\overline{z_{n}}}{2},\frac{z_{n}-\overline{z_{n}}}{2i}\right),$
$g$ is the zero polynomial. ∎
Proposition 3.3.
Let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}\setminus\{0,1\}$ is squarefree and let $g$ be a generalized polynomial in $n$ variables over $K$ given by (2). We have $g(\mathbf{a})\in\mathbb{Q}$ for every $\mathbf{a}\in K^{n}$ if and only if $\alpha_{\mathbf{j}\mathbf{i}}=\tau(\alpha_{\mathbf{i}\mathbf{j}})$ for every $\mathbf{i},\mathbf{j}\in\mathbb{Z}_{\geq 0}^{n}$.
Proof.
Consider the two generalized polynomials
$g(\mathbf{z})$ $=\sum_{|\mathbf{i}|+|\mathbf{j}|\leq k}\alpha_{\mathbf{i}\mathbf{j}}\mathbf{z}^{\mathbf{i}}\tau(\mathbf{z})^{\mathbf{j}}=\sum_{|\mathbf{i}|+|\mathbf{j}|\leq k}\alpha_{\mathbf{j}\mathbf{i}}\mathbf{z}^{\mathbf{j}}\tau(\mathbf{z})^{\mathbf{i}}$
and
$\tau(g)(\mathbf{z})=\sum_{|\mathbf{i}|+|\mathbf{j}|\leq k}\tau(\alpha_{\mathbf{i}\mathbf{j}})\tau(\mathbf{z})^{\mathbf{i}}\mathbf{z}^{\mathbf{j}}.$
The condition $g(\mathbf{a})\in\mathbb{Q}$ for every $\mathbf{a}\in K$ is equivalent to the condition $g(\mathbf{a})=\tau(g)(\mathbf{a})$ for every $\mathbf{a}\in K$. By Lemma 3.2, this is equivalent to $g-\tau(g)$ being the zero polynomial, which is in turn equivalent to $\alpha_{\mathbf{j}\mathbf{i}}=\tau(\alpha_{\mathbf{i}\mathbf{j}})$ for every $\mathbf{i},\mathbf{j}\in\mathbb{Z}_{\geq 0}^{n}$. ∎
Let $K$ be a quadratic field and let $\Gamma_{K}^{n}$ be the ring of $\mathbb{Q}$-valued generalized polynomials in $n$ variables, i.e.,
$\Gamma_{K}^{n}=\left\{g(z_{1},\dots,z_{n})\in K[z_{1},\tau(z_{1}),\dots,z_{n},\tau(z_{n})]:\;g(\mathbf{a})\in\mathbb{Q}\text{ for every }\mathbf{a}\in K^{n}\right\}.$
For $d\in\mathbb{Z}_{\geq 0}$, let
$\Gamma_{K}^{n,d}=\left\{g\in\Gamma_{K}^{n}:\;\deg(g)\leq d\right\}.$
Clearly, $\Gamma_{K}^{n,d}$ is a vector space over $\mathbb{Q}$.
The main theorem of this section is the following one.
Theorem 3.4.
Let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}\setminus\{0,1\}$ is squarefree. If $n\in\mathbb{Z}_{\geq 1}$, then
$\Gamma_{K}^{n}=\mathbb{Q}\left[z_{1}+\tau(z_{1}),\sqrt{D}(z_{1}-\tau(z_{1})),\dots,z_{n}+\tau(z_{n}),\sqrt{D}(z_{n}-\tau(z_{n}))\right].$
The first step is to prove the theorem for $n=1$. In this case, we denote the single variable by $z$ instead of $z_{1}$, thus
$\Gamma_{K}^{1}=\{g(z)\in K[z,\tau(z)]:\;g(\alpha)\in\mathbb{Q}\text{ for every}\alpha\in K\}.$
Lemma 3.5.
If $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}\setminus\{0,1\}$ is squarefree, then $\Gamma_{K}^{1}=\mathbb{Q}\left[z+\tau(z),\sqrt{D}(z-\tau(z))\right]$.
Proof.
Let $S_{1}=\mathbb{Q}\left[z+\tau(z),\sqrt{D}(z-\tau(z))\right]$. We have $S_{1}\subset\Gamma_{K}^{1}$ because $z+\tau(z)$ and $\sqrt{D}(z-\tau(z))$ are $\mathbb{Q}$-valued, and we want to prove $\Gamma_{K}^{1}\subset S_{1}$.
We claim that $z\tau(z)\in S_{1}$. Since
$D(z-\tau(z))^{2}=\left(\sqrt{D}(z-\tau(z))\right)^{2}\in S_{1},$
we have also $(z-\tau(z))^{2}\in S_{1}$, and hence
$z\tau(z)=\frac{1}{2}\left((z+\tau(z))^{2}-(z-\tau(z))^{2}\right)\in S_{1}.$
This proves the claim.
Let $g\in\Gamma_{K}^{1}$. We prove $g\in S_{1}$ by induction on $d=\deg(g)$. If $d=0$, then $g(z)=a\in\mathbb{Q}$ is a constant polynomial. Thus we assume that $d\geq 1$ and that the $\mathbb{Q}$-valued generalized polynomials of degree $<d$ are contained in $S_{1}$. We write $g$ as
$g(z)=\sum_{\begin{subarray}{c}i,j\geq 0\\ i+j\leq d\end{subarray}}\alpha_{i,j}z^{i}\tau(z)^{j}=\sum_{\begin{subarray}{c}i,j\geq 1\\ i+j\leq d\end{subarray}}\alpha_{i,j}z^{i}\tau(z)^{j}+\sum_{i=0}^{d}\alpha_{i,0}z^{i}+\sum_{j=0}^{d}\alpha_{0,j}\tau(z)^{j}.$
By Proposition 3.3, $\alpha_{j,i}=\tau(\alpha_{i,j})$. Let
$g_{1}(z)$ $=\sum_{\begin{subarray}{c}i,j\geq 1\\ i+j\leq d\end{subarray}}\alpha_{i,j}z^{i}\tau(z)^{j},$
$g_{2}(z)$ $=\sum_{i=0}^{d}\alpha_{i,0}z^{i}+\tau(\alpha_{i,0})\tau(z)^{i},$
so that $g(z)=g_{1}(z)+g_{2}(z)$. We have $g_{1}(z)=z\tau(z)\cdot h_{1}(z)$ where
$h_{1}(z)=\sum_{\begin{subarray}{c}i,j\geq 1\\ i+j\leq d\end{subarray}}\alpha_{i,j}z^{i-1}\tau(z)^{j-1}$
is a $\mathbb{Q}$-valued generalized polynomial of degree $<d$. By induction assumption, $h_{1}\in S_{1}$, and because we also know that $z\tau(z)\in S_{1}$, we get $g_{1}\in S_{1}$.
Next, we have $g_{2}(z)=\alpha_{d,0}z^{d}+\tau(\alpha_{d,0})\tau(z)^{d}+h_{2}(z)$ where
$h_{2}(z)=\sum_{i=0}^{d-1}\alpha_{i,0}z^{i}+\tau(\alpha_{i,0})\tau(z)^{i}$
is a $\mathbb{Q}$-valued generalized polynomial of degree $<d$. By induction assumption, $h_{2}\in S_{2}$. We show $\alpha_{d,0}z^{d}+\tau(\alpha_{d,0})\tau(z)^{d}\in S_{1}$. If we let $\alpha_{d,0}=a+b\sqrt{D}$ for $a,b\in\mathbb{Q}$, then
$\alpha_{d,0}z^{d}+\tau(\alpha_{d,0})\tau(z)^{d}=a(z^{d}+\tau(z)^{d})+b\sqrt{D}(z^{d}-\tau(z)^{d}).$
We have $z^{d}+\tau(z)^{d}=(z+\tau(z))^{d}-z\tau(z)\cdot r_{1}(z)$ where
$r_{1}(z)=\sum_{i=1}^{d-1}\binom{d}{i}z^{i-1}\tau(z)^{d-i-1}$
is a $\mathbb{Q}$-valued generalized polynomial of degree $<d$, hence $r_{1}\in S_{1}$. Because $(z+\tau(z))^{d}$ and $z\tau(z)$ also belong to $S_{1}$, we get $z^{d}+\tau(z)^{d}\in S_{1}$.
We also have $\sqrt{D}(z^{d}-\tau(z)^{d})=\sqrt{D}(z-\tau(z))\cdot r_{2}(z)$ where
$r_{2}(z)=\sum_{i=0}^{d-1}z^{d-1-i}\tau(z)^{i}$
is a $\mathbb{Q}$-valued generalized polynomial of degree $<d$, hence $r_{2}\in S_{1}$. Because $\sqrt{D}(z-\tau(z))\in S_{1}$, we get $\sqrt{D}(z^{d}-\tau(z)^{d})\in S_{1}$. This proves $\alpha_{d,0}z^{d}+\tau(\alpha_{d,0})\tau(z)^{d}\in S_{1}$, thus $g_{2}\in S_{1}$.
Finally, because $g(z)=g_{1}(z)+g_{2}(z)$, we get $g\in S_{1}$. ∎
Next, we find a basis of the vector space $\Gamma_{K}^{n,d}$ over $\mathbb{Q}$. Let $n\in\mathbb{Z}_{\geq 1}$ and $d\in\mathbb{Z}_{\geq 0}$. We define the polynomials $g_{\mathbf{i}\mathbf{j}}$ for $\mathbf{i},\mathbf{j}\in\mathbb{Z}_{\geq 0}^{n}$, $|\mathbf{i}|+|\mathbf{j}|\leq d$ as follows: If $\mathbf{i}<\mathbf{j}$, then
$g_{\mathbf{i}\mathbf{j}}(\mathbf{z})$ $=\mathbf{z}^{\mathbf{i}}\tau(\mathbf{z})^{\mathbf{j}}+\tau(\mathbf{z})^{\mathbf{i}}\mathbf{z}^{\mathbf{j}},$
$g_{\mathbf{j}\mathbf{i}}(\mathbf{z})$ $=\sqrt{D}\left(\mathbf{z}^{\mathbf{i}}\tau(\mathbf{z})^{\mathbf{j}}-\tau(\mathbf{z})^{\mathbf{i}}\mathbf{z}^{\mathbf{j}}\right)$
and if $\mathbf{i}=\mathbf{j}$, then
$g_{\mathbf{i}\mathbf{i}}=\mathbf{z}^{\mathbf{i}}\tau(\mathbf{z})^{\mathbf{i}}.$
Lemma 3.6.
Let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}\setminus\{0,1\}$ is squarefree, $n\in\mathbb{Z}_{\geq 1}$, and $d\in\mathbb{Z}_{\geq 0}$. The elements $g_{\mathbf{i}\mathbf{j}}$ for $\mathbf{i},\mathbf{j}\in\mathbb{Z}_{\geq 0}^{n}$, $|\mathbf{i}|+|\mathbf{j}|\leq d$ form a basis of the vector space $\Gamma_{K}^{n,d}$ over $\mathbb{Q}$.
Proof.
Let
$g(\mathbf{z})=\sum_{|\mathbf{i}|+|\mathbf{j}|\leq d}\alpha_{\mathbf{i}\mathbf{j}}\mathbf{z}^{\mathbf{i}}\tau(\mathbf{z})^{\mathbf{j}}\in\Gamma_{K}^{n,d}.$
By Proposition 3.3, $\alpha_{\mathbf{j}\mathbf{i}}=\tau(\alpha_{\mathbf{i}\mathbf{j}})$ for every $\mathbf{i},\mathbf{j}\in\mathbb{Z}_{\geq 0}^{n}$, hence
$g(\mathbf{z})=\sum_{\begin{subarray}{c}\mathbf{i}<\mathbf{j}\\ |\mathbf{i}|+|\mathbf{j}|\leq d\end{subarray}}\alpha_{\mathbf{i}\mathbf{j}}\mathbf{z}^{\mathbf{i}}\tau(\mathbf{z})^{\mathbf{j}}+\tau(\alpha_{\mathbf{i}\mathbf{j}})\tau(\mathbf{z})^{\mathbf{i}}\mathbf{z}^{\mathbf{j}}+\sum_{2|\mathbf{i}|\leq d}\alpha_{\mathbf{i}\mathbf{i}}\mathbf{z}^{\mathbf{i}}\tau(\mathbf{z})^{\mathbf{i}}.$
Let $\alpha_{\mathbf{i}\mathbf{j}}=a_{\mathbf{i}\mathbf{j}}+b_{\mathbf{i}\mathbf{j}}\sqrt{D}$ where $a_{\mathbf{i}\mathbf{j}},b_{\mathbf{i}\mathbf{j}}\in\mathbb{Q}$. In particular, $\alpha_{\mathbf{i}\mathbf{i}}=\tau(\alpha_{\mathbf{i}\mathbf{i}})$ implies $\alpha_{\mathbf{i}\mathbf{i}}=a_{\mathbf{i}\mathbf{i}}$. We get
$g(\mathbf{z})$ $=\sum_{\begin{subarray}{c}\mathbf{i}<\mathbf{j}\\ |\mathbf{i}|+|\mathbf{j}|\leq d\end{subarray}}a_{\mathbf{i}\mathbf{j}}\left(\mathbf{z}^{\mathbf{i}}\tau(\mathbf{z})^{\mathbf{j}}+\tau(\mathbf{z})^{\mathbf{i}}\mathbf{z}^{\mathbf{j}}\right)+b_{\mathbf{i}\mathbf{j}}\sqrt{D}\left(\mathbf{z}^{\mathbf{i}}\tau(\mathbf{z})^{\mathbf{j}}-\tau(\mathbf{z})^{\mathbf{i}}\mathbf{z}^{\mathbf{j}}\right)+\sum_{2|\mathbf{i}|\leq d}a_{\mathbf{i}\mathbf{i}}\mathbf{z}^{\mathbf{i}}\tau(\mathbf{z})^{\mathbf{i}},$
hence $g$ is a $\mathbb{Q}$-linear combination of the elements $g_{\mathbf{i}\mathbf{j}}$.
Let
$B=\{g_{\mathbf{i}\mathbf{j}}:\;\mathbf{i},\mathbf{j}\in\mathbb{Z}_{\geq 0}^{n},|\mathbf{i}|+|\mathbf{j}|\leq d\}.$
If $\mathbf{i}\neq\mathbf{j}$, then the only generalized polynomials in $B$ containing the term $\mathbf{z}^{\mathbf{i}}\tau(\mathbf{z})^{\mathbf{j}}$ are $g_{\mathbf{i}\mathbf{j}}$ and $g_{\mathbf{j}\mathbf{i}}$. Since they are linearly independent, neither can be expressed as a linear combination of the remaining elements in $B$. Similarly, if $\mathbf{i}=\mathbf{j}$, then $\mathbf{z}^{\mathbf{i}}\tau(\mathbf{z})^{\mathbf{i}}$ is contained only in $g_{\mathbf{i}\mathbf{i}}$, and thus $g_{\mathbf{i}\mathbf{i}}$ cannot be expressed as a linear combination of the other elements in $B$. This proves that the $g_{\mathbf{i}\mathbf{j}}$ are linearly independent. ∎
Proof of Theorem 3.4.
Let
$S_{n}=\mathbb{Q}\left[z_{1}+\tau(z_{1}),\sqrt{D}(z_{1}-\tau(z_{1})),\dots,z_{n}+\tau(z_{n}),\sqrt{D}(z_{n}-\tau(z_{n}))\right].$
We have $S_{n}\subset\Gamma_{K}^{n}$ because the generators of $S_{n}$ are $\mathbb{Q}$-valued, and we want to prove $\Gamma_{K}^{n}\subset S_{n}$.
We proceed by induction on $n$. For $n=1$, this is Lemma 3.5. Suppose that $n\geq 2$ and the statement is true for polynomials with $<n$ variables.
Let $g\in\Gamma_{K}^{n}$. We show $g\in S_{n}$ by induction on $d=\deg(g)$. If $d=0$, then $g(\mathbf{z})=a\in\mathbb{Q}$ is a constant polynomial. Thus we assume that $d\geq 1$ and that the $\mathbb{Q}$-valued generalized polynomials of degree $<d$ in $n$ variables are contained in $S_{n}$. The polynomial $g$ is a $\mathbb{Q}$-linear combination of the elements in the basis from Lemma 3.6. Thus it is enough to show that the basis elements of degree $d$ belong to $S_{n}$.
We show that if $\mathbf{i},\mathbf{j}\in\mathbb{Z}_{\geq 0}^{n}$, $|\mathbf{i}|+|\mathbf{j}|=d$, and $\mathbf{i}<\mathbf{j}$, then
$g_{\mathbf{i}\mathbf{j}}(\mathbf{z})$ $=\mathbf{z}^{\mathbf{i}}\tau(\mathbf{z})^{\mathbf{j}}+\tau(\mathbf{z})^{\mathbf{i}}\mathbf{z}^{\mathbf{j}}\in S_{n},$
$g_{\mathbf{j}\mathbf{i}}(\mathbf{z})$ $=\sqrt{D}\left(\mathbf{z}^{\mathbf{i}}\tau(\mathbf{z})^{\mathbf{j}}-\tau(\mathbf{z})^{\mathbf{i}}\mathbf{z}^{\mathbf{j}}\right)\in S_{n}.$
If $i_{n}=0$ and $j_{n}=0$, then the variable $z_{n}$ does not appear in $g_{\mathbf{i}\mathbf{j}}$ and $g_{\mathbf{j}\mathbf{i}}$, hence $g_{\mathbf{i}\mathbf{j}},g_{\mathbf{j}\mathbf{i}}\in S_{n}$ by induction assumption.
If $i_{n}\geq 1$, then we let
$h(\mathbf{z})$ $=z_{1}^{i_{1}}\tau(z_{1})^{j_{1}}z_{2}^{i_{2}}\tau(z_{2})^{j_{2}}\cdots z_{n}^{i_{n}-1}\tau(z_{n})^{j_{n}},$
$g^{1}_{\mathbf{i}\mathbf{j}}(\mathbf{z})$ $=(z_{n}+\tau(z_{n}))\cdot(h(\mathbf{z})+\tau(h(\mathbf{z}))),$
$g^{2}_{\mathbf{i}\mathbf{j}}(\mathbf{z})$ $=(z_{n}-\tau(z_{n}))\cdot(h(\mathbf{z})-\tau(h(\mathbf{z}))),$
$g^{1}_{\mathbf{j}\mathbf{i}}(\mathbf{z})$ $=\sqrt{D}(z_{n}-\tau(z_{n}))\cdot(h(\mathbf{z})+\tau(h(\mathbf{z}))),$
$g^{2}_{\mathbf{j}\mathbf{i}}(\mathbf{z})$ $=(z_{n}+\tau(z_{n}))\cdot\sqrt{D}(h(\mathbf{z})-\tau(h(\mathbf{z}))).$
so that
$g_{\mathbf{i}\mathbf{j}}(\mathbf{z})$ $=z_{n}h(\mathbf{z})+\tau(z_{n})\tau(h(\mathbf{z}))=\frac{1}{2}\left(g^{1}_{\mathbf{i}\mathbf{j}}+g^{2}_{\mathbf{i}\mathbf{j}}\right),$
$g_{\mathbf{j}\mathbf{i}}(\mathbf{z})$ $=\sqrt{D}\left(z_{n}h(\mathbf{z})-\tau(z_{n})\tau(h(\mathbf{z}))\right)=\frac{1}{2}\left(g^{1}_{\mathbf{j}\mathbf{i}}+g^{2}_{\mathbf{j}\mathbf{i}}\right).$
Because $h(\mathbf{z})$ has degree $d-1$, we can apply induction on $h(\mathbf{z})+\tau(h(\mathbf{z}))$ and $\sqrt{D}(h(\mathbf{z})-\tau(h(\mathbf{z})))$. Since
$z_{n}+\tau(z_{n}),\ \sqrt{D}(z_{n}-\tau(z_{n})),\ h(\mathbf{z})+\tau(h(\mathbf{z})),\ \sqrt{D}(h(\mathbf{z})-\tau(h(\mathbf{z})))\in S_{n},$
we immediately obtain $g^{1}_{\mathbf{i}\mathbf{j}},g^{1}_{\mathbf{j}\mathbf{i}},g^{2}_{\mathbf{j}\mathbf{i}}\in S_{n}$. We also get
$g^{2}_{\mathbf{i}\mathbf{j}}=\frac{1}{D}\cdot\sqrt{D}(z_{n}-\tau(z_{n}))\cdot\sqrt{D}(h(\mathbf{z})-\tau(h(\mathbf{z})))\in S_{n},$
and $g_{\mathbf{i}\mathbf{j}},g_{\mathbf{j}\mathbf{i}}\in S_{n}$ follows.
If $i_{n}=0$ and $j_{n}\geq 1$, then the proof is analogous to the case $i_{n}\geq 1$.
Next, we show that if $\mathbf{i}\in\mathbb{Z}_{\geq 0}^{n}$ and $2|\mathbf{i}|=d$, then $g_{\mathbf{i}\mathbf{i}}(\mathbf{z})=\mathbf{z}^{\mathbf{i}}\tau(\mathbf{z})^{\mathbf{i}}\in S_{n}$.
If $i_{n}=0$, then $\mathbf{g}_{ii}$ is a polynomial in $n-1$ variables, hence $\mathbf{g}_{ii}\in S_{n}$ by induction assumption.
If $i_{n}\geq 1$, then we let
$h(\mathbf{z})=z_{1}^{i_{1}}\tau(z_{1})^{i_{1}}z_{2}^{i_{2}}\tau(z_{2})^{i_{2}}\cdots z_{n}^{i_{n}-1}\tau(z_{n})^{i_{n}-1}.$
We have $g_{\mathbf{i}\mathbf{i}}(\mathbf{z})=z_{n}\tau(z_{n})\cdot h(\mathbf{z})$. Since the theorem holds for generalized polynomials in one variable, we get
$z_{n}\tau(z_{n})\in\mathbb{Q}\left[z_{n}+\tau(z_{n}),\sqrt{D}\left(z_{n}-\tau(z_{n})\right)\right]\subset S_{n}.$
The generalized polynomial $h$ has degree $d-2$, hence $h\in S_{n}$ by induction assumption, and $g_{\mathbf{i}\mathbf{i}}\in S_{n}$ follows. ∎
4. Generalized quadratic forms
Generalized quadratic forms are defined in [BPS, Definition 1.1] over a Galois extension $K/\mathbb{Q}$ of degree $d$ (which is assumed to be totally real). We need the definition only over quadratic fields.
Definition 4.1.
Let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}\setminus\{0,1\}$ is squarefree. A generalized quadratic form in $n$ variables over $K$ is a generalized polynomial
(3) $G(z_{1},z_{2},\dots,z_{n})=\sum_{1\leq i\leq j\leq n}\alpha_{ij}z_{i}z_{j}+\sum_{1\leq i,j\leq n}\beta_{ij}z_{i}\tau(z_{j})+\sum_{1\leq i\leq j\leq n}\gamma_{ij}\tau(z_{i})\tau(z_{j})$
where $\alpha_{ij},\gamma_{ij}\in K$ for $1\leq i\leq j\leq n$ and $\beta_{ij}\in K$ for $1\leq i,j\leq n$.
A generalized quadratic form $G$ given by (3) will be called integral if $\alpha_{ij},\gamma_{ij}\in\mathcal{O}_{K}$ for $1\leq i\leq j\leq n$ and $\beta_{ij}\in\mathcal{O}_{K}$ for $1\leq i,j\leq n$. An integral generalized quadratic form will be called classical if $\alpha_{ij},\gamma_{ij}\in 2\mathcal{O}_{K}$ for $i<j$ and $\beta_{ij}\in 2\mathcal{O}_{K}$ for $1\leq i,j\leq n$.
If $G$ is a generalized quadratic form in $n$ variables over $K$, then we can define a quadratic form $Q$ in $2n$ variables by
(4) $Q(x_{1},y_{1},\dots,x_{n},y_{n})=G(x_{1}+y_{1}\omega_{D},\dots,x_{n}+y_{n}\omega_{D}).$
In other words, $Q$ is obtained from $G$ by expressing the variables $z_{1},z_{2},\dots,z_{n}$ in the integral basis $(1,\omega_{D})$. We call $Q$ the quadratic form associated to $G$. It has coefficients in $K$ and the variables $x_{1},y_{1},\dots,x_{n},y_{n}$ take values in $\mathbb{Q}$.
If $G$ is a generalized form in $n$ variables over $K$ given by (3), then we define the matrix of $G$ as the $2n\times 2n$ matrix
$M_{G}=\begin{pmatrix}A&B\\ B^{\top}&C\end{pmatrix}$
where
$A=\begin{pmatrix}\alpha_{11}&\frac{\alpha_{12}}{2}&\dots&\frac{\alpha_{1n}}{2}\\ \frac{\alpha_{12}}{2}&\alpha_{22}&\dots&\frac{\alpha_{2n}}{2}\\ \vdots&\vdots&\ddots&\vdots\\ \frac{\alpha_{1n}}{2}&\frac{\alpha_{2n}}{2}&\dots&\alpha_{nn}\end{pmatrix},\ B=\begin{pmatrix}\frac{\beta_{11}}{2}&\frac{\beta_{12}}{2}&\dots&\frac{\beta_{1n}}{2}\\ \frac{\beta_{21}}{2}&\frac{\beta_{22}}{2}&\dots&\frac{\beta_{2n}}{2}\\ \vdots&\vdots&\ddots&\vdots\\ \frac{\beta_{n1}}{2}&\frac{\beta_{n2}}{2}&\dots&\frac{\beta_{nn}}{2}\end{pmatrix},\ C=\begin{pmatrix}\gamma_{11}&\frac{\gamma_{12}}{2}&\dots&\frac{\gamma_{1n}}{2}\\ \frac{\gamma_{12}}{2}&\gamma_{22}&\dots&\frac{\gamma_{2n}}{2}\\ \vdots&\vdots&\ddots&\vdots\\ \frac{\gamma_{1n}}{2}&\frac{\gamma_{2n}}{2}&\dots&\gamma_{nn}\end{pmatrix}.$
Thus $M_{G}$ is the matrix such that
$G(\mathbf{z})=(\mathbf{z},\tau(\mathbf{z}))M_{G}\begin{pmatrix}\mathbf{z}\\ \tau(\mathbf{z})\end{pmatrix}$
for $\mathbf{z}\in K^{n}$.
The $2n\times 2n$ matrix
$T=\begin{pmatrix}1&\omega_{D}&0&0&\dots&0&0\\ 0&0&1&\omega_{D}&\dots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots&\vdots\\ 0&0&\dots&0&0&1&\omega_{D}\\ 1&\tau(\omega_{D})&0&0&\dots&0&0\\ 0&0&1&\tau(\omega_{D})&\dots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots&\vdots\\ 0&0&\dots&0&0&1&\tau(\omega_{D})\\ \end{pmatrix}$
determines the change of variables $(x_{1},y_{1},\dots,x_{n},y_{n})\mapsto(z_{1},\dots,z_{n},\tau(z_{1}),\dots,\tau(z_{n}))$. Thus, if $M_{Q}$ is the matrix of the associated quadratic form $Q$ defined by (4), then
$M_{Q}=T^{\top}M_{G}T.$
Lemma 4.2.
Let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}\setminus\{0,1\}$ is squarefree, $G$ be a generalized form in $n$ variables over $K$ given by (3), and $M_{G}$ be the matrix of $G$. If $Q$ is the quadratic form in $2n$ variables associated to $G$ and $M_{Q}$ is the matrix of $Q$, then
$\det(M_{Q})=D^{n}\cdot\begin{cases}\det(2M_{G}),&\text{if }D\equiv 2,3\pmod{4},\\ \det(M_{G}),&\text{if }D\equiv 1\pmod{4}.\end{cases}$
Proof.
Since $M_{G}=T^{\top}M_{Q}T$, we get
$\det(M_{Q})=(\det T)^{2}\det(M_{G}).$
Let us compute the determinant of $T$. If $\pi$ is the permutation
$\pi=\begin{pmatrix}1&2&3&4&\dots&2n-1&2n\\ 1&n+1&2&n+2&\dots&n&2n\end{pmatrix}$
(the out-shuffle), then $\operatorname{sgn}(\pi)=(-1)^{\frac{n(n-1)}{2}}$. Applying this permutation to the rows of $T$, we obtain a block-diagonal matrix where each block equals
$M=\begin{pmatrix}1&\omega_{D}\\ 1&\tau(\omega_{D})\end{pmatrix}.$
If $D\equiv 2,3\pmod{4}$, then
$\det M=\det\begin{pmatrix}1&\sqrt{D}\\ 1&-\sqrt{D}\end{pmatrix}=-2\sqrt{D},$
while if $D\equiv 1\pmod{4}$, then
$\det M=\det\begin{pmatrix}1&\frac{1+\sqrt{D}}{2}\\ 1&\frac{1-\sqrt{D}}{2}\end{pmatrix}=-\sqrt{D}.$
Thus
$\det T=(-1)^{\frac{n(n-1)}{2}}(\det M)^{n}=\begin{cases}(-1)^{\frac{n(n+1)}{2}}D^{\frac{n}{2}}2^{n},&\text{if }D\equiv 2,3\pmod{4},\\ (-1)^{\frac{n(n+1)}{2}}D^{\frac{n}{2}},&\text{if }D\equiv 1\pmod{4}.\end{cases}$
In the case $D\equiv 2,3\pmod{4}$, we get
$\det(M_{Q})=D^{n}2^{2n}\det(M_{G})=D^{n}\det(2M_{G}),$
and in the case $D\equiv 1\pmod{4}$, we get
$\det(M_{Q})=D^{n}\det(M_{G}).\qed$
Corollary 4.3.
Let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}\setminus\{0,1\}$ is squarefree, $G$ be a generalized form in $n$ variables over $K$ given by (3), $Q$ be the quadratic form in $2n$ variables associated to $G$, and $M_{Q}$ be the matrix of $Q$. If one of the following conditions is satisfied:
  1. i)
    $D\equiv 2,3\pmod{4}$ and $G$ is integral,
  2. ii)
    $D\equiv 1\pmod{4}$ and $G$ is classical,
then $D^{n}\mid\det(M_{Q})$.
Proof.
First, assume that $D\equiv 2,3\pmod{4}$ and $G$ is integral. Since the elements of $M_{G}$ are in $\frac{1}{2}\mathcal{O}_{K}$, we must have $\det(2M_{G})\in\mathcal{O}_{K}$. Thus $D^{n}\mid\det(M_{Q})$ by Lemma 4.2.
Secondly, assume that $D\equiv 1\pmod{4}$ and $G$ is classical. The elements of $M_{G}$ are in $\mathcal{O}_{K}$, hence $\det(M_{G})\in\mathcal{O}_{K}$, and $D^{n}\mid\det(M_{Q})$ by Lemma 4.2. ∎
Lemma 4.4.
Let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}\setminus\{0,1\}$ is squarefree and let $G$ be a generalized form in $n$ variables over $K$ given by (3). Assume that $G(\mathbf{a})\in\mathbb{Q}$ for every $\mathbf{a}\in K^{n}$. Let $Q$ be the quadratic form in $2n$ variables associated to $G$. If $1\leq i\leq j\leq n$, then we let $a_{ij}$, $b_{ij}$, $b_{ji}$, and $c_{ij}$ be the coefficients of the terms $x_{i}x_{j}$, $x_{i}y_{j}$, $x_{j}y_{i}$, and $y_{i}y_{j}$ in $Q$, respectively, so that the matrix $M_{Q}$ of the quadratic form $Q$ is
$M_{Q}=\begin{pmatrix}\begin{matrix}a_{11}&\frac{b_{11}}{2}\\ \frac{b_{11}}{2}&c_{11}\end{matrix}&\begin{matrix}\frac{a_{12}}{2}&\frac{b_{12}}{2}\\ \frac{b_{21}}{2}&\frac{c_{12}}{2}\end{matrix}&\dots&\begin{matrix}\frac{a_{1n}}{2}&\frac{b_{1n}}{2}\\ \frac{b_{n1}}{2}&\frac{c_{1n}}{2}\end{matrix}\\ \begin{matrix}\frac{a_{12}}{2}&\frac{b_{21}}{2}\\ \frac{b_{12}}{2}&\frac{c_{12}}{2}\end{matrix}&\begin{matrix}a_{22}&\frac{b_{22}}{2}\\ \frac{b_{22}}{2}&c_{22}\end{matrix}&\dots&\begin{matrix}\frac{a_{2n}}{2}&\frac{b_{2n}}{2}\\ \frac{b_{n2}}{2}&\frac{c_{2n}}{2}\end{matrix}\\ \vdots&\vdots&\ddots&\vdots\\ \begin{matrix}\frac{a_{1n}}{2}&\frac{b_{n1}}{2}\\ \frac{b_{1n}}{2}&\frac{c_{1n}}{2}\end{matrix}&\begin{matrix}\frac{a_{2n}}{2}&\frac{b_{n2}}{2}\\ \frac{b_{2n}}{2}&\frac{c_{2n}}{2}\end{matrix}&\dots&\begin{matrix}a_{nn}&\frac{b_{nn}}{2}\\ \frac{b_{nn}}{2}&c_{nn}\end{matrix}\end{pmatrix}.$
If $1\leq i<j\leq n$, then
$a_{ij}$ $=\operatorname{Tr}(\alpha_{ij})+\operatorname{Tr}(\beta_{ij}),$
$b_{ij}$ $=\operatorname{Tr}(\alpha_{ij}\omega_{D})+\operatorname{Tr}(\beta_{ij}\tau(\omega_{D})),$
$b_{ji}$ $=\operatorname{Tr}(\alpha_{ij}\omega_{D})+\operatorname{Tr}(\beta_{ij}\omega_{D}),$
$c_{ij}$ $=\operatorname{Tr}(\alpha_{ij}\omega_{D}^{2})+\operatorname{Tr}(\beta_{ij})\operatorname{N}(\omega_{D}).$
Moreover, if $1\leq i\leq n$, then
$a_{ii}$ $=\operatorname{Tr}(\alpha_{ii})+\beta_{ii},$
$b_{ii}$ $=\operatorname{Tr}(2\alpha_{ii}\omega_{D})+\beta_{ii}\operatorname{Tr}(\omega_{D}),$
$c_{ii}$ $=\operatorname{Tr}(\alpha_{ii}\omega_{D}^{2})+\beta_{ii}\operatorname{N}(\omega_{D}).$
Proof.
The quadratic form $Q$ is given by
$Q(x_{1},y_{1},\dots,x_{n},y_{n})$ $=\sum_{1\leq i\leq j\leq n}\alpha_{ij}(x_{i}+y_{i}\omega)(x_{j}+y_{j}\omega)+\sum_{1\leq i,j\leq n}\beta_{ij}(x_{i}+y_{i}\omega)(x_{j}+y_{j}\tau(\omega))$
$+\sum_{1\leq i\leq j\leq n}\gamma_{ij}(x_{i}+y_{i}\tau(\omega))(x_{j}+y_{j}\tau(\omega)).$
By Proposition 3.3, $G(\mathbf{a})\in\mathbb{Q}$ for every $\mathbf{a}\in K^{n}$ if and only if the coefficients of conjugate terms are conjugate, i.e., $\gamma_{ij}=\tau(\alpha_{ij})$ for $1\leq i\leq j\leq n$ and $\beta_{ji}=\tau(\beta_{ij})$ for $1\leq i,j\leq n$. If $i<j$, then
$a_{ij}$ $=\alpha_{ij}+\beta_{ij}+\beta_{ji}+\gamma_{ij}=\operatorname{Tr}(\alpha_{ij})+\operatorname{Tr}(\beta_{ij}),$
$b_{ij}$ $=\alpha_{ij}\omega+\beta_{ij}\tau(\omega)+\beta_{ji}\omega+\gamma_{ij}\tau(\omega)=\operatorname{Tr}(\alpha_{ij}\omega)+\operatorname{Tr}(\beta_{ij}\tau(\omega)),$
$b_{ji}$ $=\alpha_{ij}\omega+\beta_{ij}\omega+\beta_{ji}\tau(\omega)+\gamma_{ij}\tau(\omega)=\operatorname{Tr}(\alpha_{ij}\omega)+\operatorname{Tr}(\beta_{ij}\omega),$
$c_{ij}$ $=\alpha_{ij}\omega^{2}+(\beta_{ij}+\beta_{ji})\operatorname{N}(\omega)+\gamma_{ij}\tau(\omega)^{2}=\operatorname{Tr}(\alpha_{ij}\omega^{2})+\operatorname{Tr}(\beta_{ij})\operatorname{N}(\omega).$
Moreover,
$a_{ii}$ $=\alpha_{ii}+\beta_{ii}+\gamma_{ii}=\operatorname{Tr}(\alpha_{ii})+\beta_{ii},$
$b_{ii}$ $=2\alpha_{ii}\omega+\beta_{ii}\operatorname{Tr}(\omega)+2\gamma_{ii}\tau(\omega)=\operatorname{Tr}(2\alpha_{ii}\omega)+\beta_{ii}\operatorname{Tr}(\omega),$
$c_{ii}$ $=\alpha_{ii}\omega^{2}+\beta_{ii}\operatorname{N}(\omega)+\gamma_{ii}\tau(\omega)^{2}=\operatorname{Tr}(\alpha_{ii}\omega^{2})+\beta_{ii}\operatorname{N}(\omega).\qed$
Lemma 4.5.
Let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}\setminus\{0,1\}$ is squarefree and let $G$ be a generalized form in $n$ variables over $K$ given by (3). Assume that $G$ is $\mathbb{Z}$-valued. Let $Q$ be the quadratic form in $2n$ variables associated to $G$. If $G$ is integral, then $Q$ is integral. Moreover, if one of the following conditions holds:
  1. (i)
    $D\equiv 2,3\pmod{4}$ and $G$ is integral,
  2. (ii)
    $D\equiv 1\pmod{4}$ and $G$ is classical,
then $Q$ is classical.
Proof.
As before, we let $a_{ij}$, $b_{ij}$, $b_{ji}$, and $c_{ij}$ be the coefficients of the terms $x_{i}x_{j}$, $x_{i}y_{j}$, $x_{j}y_{i}$, and $y_{i}y_{j}$ in $Q$, respectively, for $1\leq i\leq j\leq n$. If $G$ is integral, then $\alpha_{ij},\beta_{ij}\in\mathcal{O}_{K}$ for $i\leq j$. The elements of $\mathcal{O}_{K}$ have trace in $\mathbb{Z}$, hence $a_{ij},b_{ij},b_{ji},c_{ij}\in\mathbb{Z}$ by Lemma 4.4 and $Q$ is integral.
If $G$ is classical, then $\alpha_{ij},\beta_{ij}\in 2\mathcal{O}_{K}$ for $i<j$ and $\beta_{ii}\in 2\mathbb{Z}$. It follows from Lemma 4.4 that $a_{ij},b_{ij},b_{ji},c_{ij}\in 2\mathbb{Z}$ for $i<j$ and $b_{ii}\in 2\mathbb{Z}$, hence $Q$ is classical.
If $D\equiv 2,3\pmod{4}$, then $\operatorname{Tr}(\alpha)$ is even for every $\alpha\in\mathcal{O}_{K}$. Thus it is sufficient to assume that $G$ is integral to conclude that $Q$ is classical. ∎
Lemma 4.6.
Let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}\setminus\{0,1\}$ is squarefree and let $G$ be a generalized form in $n$ variables given by (3). Assume that $G(\mathbf{a})\in\mathbb{Q}$ for every $\mathbf{a}\in K^{2}$. If $1\leq i\leq j\leq n$, then we let $\alpha_{ij}=r_{ij}+s_{ij}\omega_{D}$ and $\beta_{ij}=t_{ij}+u_{ij}\omega_{D}$ where $r_{ij},s_{ij},t_{ij},u_{ij}\in\mathbb{Q}$. Let $Q$ be the quadratic form in $2n$ variables associated to $G$. If $1\leq i\leq j\leq n$, then we let $a_{ij}$, $b_{ij}$, $b_{ji}$, and $c_{ij}$ be the coefficients of the terms $x_{i}x_{j}$, $x_{i}y_{j}$, $x_{j}y_{i}$, and $y_{i}y_{j}$ in $Q$, respectively.
If $D\equiv 2,3\pmod{4}$ and $1\leq i<j\leq n$, then
$a_{ij}$ $=2r_{ij}+2t_{ij},$
$b_{ij}$ $=2Ds_{ij}-2Du_{ij},$
$b_{ji}$ $=2Ds_{ij}+2Du_{ij},$
$c_{ij}$ $=2Dr_{ij}-2Dt_{ij}.$
Moreover, if $1\leq i\leq n$, then
$a_{ii}$ $=2r_{ii}+t_{ii},$
$b_{ii}$ $=4Ds_{ii},$
$c_{ii}$ $=2Dr_{ii}-Dt_{ii}.$
If $D\equiv 1\pmod{4}$ and $1\leq i<j\leq n$, then
$a_{ij}$ $=2r_{ij}+s_{ij}+2t_{ij}+u_{ij},$
$b_{ij}$ $=r_{ij}+\frac{1+D}{2}s_{ij}+t_{ij}+\frac{1-D}{2}u_{ij},$
$b_{ji}$ $=r_{ij}+\frac{1+D}{2}s_{ij}+t_{ij}+\frac{1+D}{2}u_{ij},$
$c_{ij}$ $=\frac{1+D}{2}r_{ij}+\frac{1+3D}{4}s_{ij}+\frac{1-D}{2}t_{ij}+\frac{1-D}{4}u_{ij}.$
Moreover, if $1\leq i\leq n$, then
$a_{ii}$ $=2r_{ii}+s_{ii}+t_{ii},$
$b_{ii}$ $=2r_{ii}+(1+D)s_{ii}+t_{ii},$
$c_{ii}$ $=\frac{1+D}{2}r_{ii}+\frac{1+3D}{4}s_{ii}+\frac{1-D}{4}t_{ii}.$
Proof.
We express the coefficients of $Q$ using Lemma 4.4. If $1\leq i<j\leq n$, then
$a_{ij}$ $=\operatorname{Tr}(\alpha_{ij})+\operatorname{Tr}(\beta_{ij})=\operatorname{Tr}(r_{ij}+s_{ij}\omega_{D})+\operatorname{Tr}(t_{ij}+u_{ij}\omega_{D})$
$=2r_{ij}+s_{ij}\operatorname{Tr}(\omega_{D})+2t_{ij}+u_{ij}\operatorname{Tr}(\omega_{D}),$
$b_{ij}$ $=\operatorname{Tr}(\alpha_{ij}\omega_{D})+\operatorname{Tr}(\beta_{ij}\tau(\omega_{D}))=\operatorname{Tr}(r_{ij}\omega_{D}+s_{ij}\omega_{D}^{2})+\operatorname{Tr}(t_{ij}\tau(\omega_{D})+u_{ij}\operatorname{N}(\omega_{D}))$
$=r_{ij}\operatorname{Tr}(\omega_{D})+s_{ij}\operatorname{Tr}(\omega_{D}^{2})+t_{ij}\operatorname{Tr}(\omega_{D})+2u_{ij}\operatorname{N}(\omega_{D}),$
$b_{ji}$ $=\operatorname{Tr}(\alpha_{ij}\omega_{D})+\operatorname{Tr}(\beta_{ij}\omega_{D})=\operatorname{Tr}(r_{ij}\omega_{D}+s_{ij}\omega_{D}^{2})+\operatorname{Tr}(t_{ij}\omega_{D}+u_{ij}\omega_{D}^{2})$
$=r_{ij}\operatorname{Tr}(\omega_{D})+s_{ij}\operatorname{Tr}(\omega_{D}^{2})+t_{ij}\operatorname{Tr}(\omega_{D})+u_{ij}\operatorname{Tr}(\omega_{D}^{2}),$
$c_{ij}$ $=\operatorname{Tr}(\alpha_{ij}\omega_{D}^{2})+\operatorname{Tr}(\beta_{ij})\operatorname{N}(\omega_{D})=\operatorname{Tr}(r_{ij}\omega_{D}^{2}+s_{ij}\omega_{D}^{3})+\operatorname{Tr}(t_{ij}+u_{ij}\omega_{D})\operatorname{N}(\omega_{D})$
$=r_{ij}\operatorname{Tr}(\omega_{D}^{2})+s_{ij}\operatorname{Tr}(\omega_{D}^{3})+2t_{ij}\operatorname{N}(\omega_{D})+u_{ij}\operatorname{Tr}(\omega_{D})\operatorname{N}(\omega_{D}).$
Moreover, if $1\leq i\leq n$, then
$a_{ii}$ $=\operatorname{Tr}(\alpha_{ii})+\beta_{ii}=\operatorname{Tr}(r_{ii}+s_{ii}\omega_{D})+t_{ii}=2r_{ii}+s_{ii}\operatorname{Tr}(\omega_{D})+t_{ii},$
$b_{ii}$ $=\operatorname{Tr}(2\alpha_{ii}\omega_{D})+\beta_{ii}\operatorname{Tr}(\omega_{D})=\operatorname{Tr}(2r_{ii}\omega_{D}+2s_{ii}\omega_{D}^{2})+t_{ii}\operatorname{Tr}(\omega_{D})=2r_{ii}\operatorname{Tr}(\omega_{D})+2s_{ii}\operatorname{Tr}(\omega_{D}^{2})+t_{ii}\operatorname{Tr}(\omega_{D}),$
$c_{ii}$ $=\operatorname{Tr}(\alpha_{ii}\omega_{D}^{2})+\beta_{ii}\operatorname{N}(\omega_{D})=\operatorname{Tr}(r_{ii}\omega_{D}^{2}+s_{ii}\omega_{D}^{3})+t_{ii}\operatorname{N}(\omega_{D})=r_{ii}\operatorname{Tr}(\omega_{D}^{2})+s_{ii}\operatorname{Tr}(\omega_{D}^{3})+t_{ii}\operatorname{N}(\omega_{D}).$
If $D\equiv 2,3\pmod{4}$, then $\omega_{D}=\sqrt{D}$, hence $\operatorname{Tr}(\omega_{D})=0$, $\operatorname{Tr}(\omega_{D}^{2})=\operatorname{Tr}(D)=2D$, $\operatorname{Tr}(\omega_{D}^{3})=\operatorname{Tr}(D\sqrt{D})=0$, and $\operatorname{N}(\omega_{D})=-D$.
If $D\equiv 1\pmod{4}$, then $\omega_{D}=\frac{1+\sqrt{D}}{2}$, hence $\operatorname{Tr}(\omega_{D})=1$, $\operatorname{Tr}(\omega_{D}^{2})=\operatorname{Tr}\left(\frac{1+D}{4}+\frac{\sqrt{D}}{2}\right)=\frac{1+D}{2}$, $\operatorname{Tr}(\omega_{D}^{3})=\operatorname{Tr}\left(\frac{1+3D}{8}+\frac{3+D}{8}\sqrt{D}\right)=\frac{1+3D}{4}$, and $\operatorname{N}(\omega_{D})=\frac{1-D}{4}$.
After substituting these expressions into the formulas above, we get the lemma. ∎
If $p$ is a prime and $M$ a matrix with entries in $\mathbb{Z}$, then we let $M\bmod p$ denote the matrix with entries in $\mathbb{Z}/p\mathbb{Z}$ obtained by reducing the elements of $M$ modulo $p$.
Proposition 4.7.
Let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}\setminus\{0,1\}$ is squarefree, $p\mid D$ be a prime number, and $G$ an integral generalized quadratic form in $n$ variables over $K$. Assume that $G$ is $\mathbb{Z}$-valued. If $Q$ is the quadratic form associated to $G$ and $M_{Q}$ is the matrix of $Q$, then $\operatorname{rank}(M_{Q}\bmod p)\leq n$.
Proof.
We introduce the following notation: if $M$ is a $2n\times 2n$ matrix and $1\leq k\leq 2n$, let $[M]_{k}$ denote the $k$-th row of $M$.
If $1\leq i\leq j\leq n$, then we let $\alpha_{ij}=r_{ij}+s_{ij}\omega_{D}$ and $\beta_{ij}=t_{ij}+u_{ij}\omega_{D}$ where $r_{ij},s_{ij},t_{ij},u_{ij}\in\mathbb{Q}$.
First assume that $D\equiv 2,3\pmod{4}$. If $1\leq k\leq n$, then the $2k$-th row of $M_{Q}$ equals
$[M_{Q}]_{2k}=\begin{pmatrix}\frac{b_{1k}}{2}&\frac{c_{1k}}{2}&\dots&\frac{b_{kk}}{2}&c_{kk}&\dots&\frac{b_{nk}}{2}&\frac{c_{kn}}{2}\end{pmatrix}.$
By Lemma 4.6, if $1\leq i<j\leq n$, then
$\frac{b_{ij}}{2}=D(s_{ij}-u_{ij}),\ \frac{b_{ji}}{2}=D(s_{ij}+u_{ij}),\ \frac{c_{ij}}{2}=D(r_{ij}-t_{ij}),$
and if $1\leq i\leq n$, then
$\frac{b_{ii}}{2}=2Ds_{ii},\ c_{ii}=D(2r_{ii}-t_{ii}).$
Since we assume that $G$ is integral, $r_{ij},s_{ij},t_{ij},u_{ij}\in\mathbb{Z}$. Thus $D$ divides every element of $[M_{Q}]_{2k}$. It follows that $M_{Q}\bmod{p}$ contains $n$ zero rows, hence $\operatorname{rank}(M_{Q}\bmod{p})\leq n$.
Secondly, assume that $D\equiv 1\pmod{4}$. Let $N_{Q}$ be the matrix obtained from $M_{Q}$ by subtracting two times the $2k$-th row of $M_{Q}$ from the $(2k-1)$-th row of $M_{Q}$ for every $1\leq k\leq n$. The $(2k-1)$-th row of $N_{Q}$ equals
$[M_{Q}]_{2k-1}-2[M_{Q}]_{2k}=\begin{pmatrix}\frac{a_{1k}}{2}-b_{1k}&\frac{b_{k1}}{2}-c_{1k}&\dots&a_{kk}-b_{kk}&\frac{b_{kk}}{2}-2c_{kk}&\dots&\frac{a_{kn}}{2}-b_{nk}&\frac{b_{kn}}{2}-c_{kn}\end{pmatrix}.$
By Lemma 4.6, if $1\leq i<j\leq n$, then
$\frac{a_{ij}}{2}-b_{ij}$ $=\frac{1}{2}\left(2r_{ij}+s_{ij}+2t_{ij}+u_{ij}\right)-\left(r_{ij}+\frac{1+D}{2}s_{ij}+t_{ij}+\frac{1-D}{2}u_{ij}\right)=D\cdot\frac{u_{ij}+s_{ij}}{2},$
$\frac{b_{ji}}{2}-c_{ij}$ $=\frac{1}{2}\left(r_{ij}+\frac{1+D}{2}s_{ij}+t_{ij}+\frac{1+D}{2}u_{ij}\right)-\left(\frac{1+D}{2}r_{ij}+\frac{1+3D}{4}s_{ij}+\frac{1-D}{2}t_{ij}+\frac{1-D}{4}u_{ij}\right)$
$=D\cdot\frac{t_{ij}+u_{ij}-r_{ij}-s_{ij}}{2},$
and if $1\leq i\leq n$, then
$a_{ii}-b_{ii}$ $=2r_{ii}+s_{ii}+t_{ii}-\left(2r_{ii}+(1+D)s_{ii}+t_{ii}\right)=-Ds_{ii},$
$\frac{b_{ii}}{2}-2c_{ii}$ $=\frac{1}{2}\left(2r_{ii}+(1+D)s_{ii}+t_{ii}\right)-2\left(\frac{1+D}{2}r_{ii}+\frac{1+3D}{4}s_{ii}+\frac{1-D}{4}t_{ii}\right)=-Dr_{ii}-Ds_{ii}+\frac{D}{2}t_{ii}.$
We have $r_{ij},s_{ij},t_{ij},u_{ij}\in\frac{1}{2}\mathbb{Z}$ because $G$ is integral. Since we assume $D\equiv 1\pmod{4}$, $p$ is an odd prime and $2$ is invertible modulo $p$. We see that every element of $[N_{Q}]_{2k-1}$ vanishes modulo $p$, hence $N_{Q}\bmod{p}$ contains $n$ zero rows and $\operatorname{rank}(M_{Q}\bmod{p})=\operatorname{rank}(N_{Q}\bmod{p})\leq n$. ∎
5. Binary and ternary generalized quadratic forms
Let us formally state a definition from the introduction.
Definition 5.1.
Let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}_{\geq 2}$ is squarefree and let $G$ be an integral generalized form in $n$ variables over $K$. Assume that $G$ is $\mathbb{Z}$-valued. If for every $a\in\mathbb{Z}_{\geq 1}$, there exist $\alpha_{1},\alpha_{2},\dots,\alpha_{n}\in\mathcal{O}_{K}$ such that
$G(\alpha_{1},\alpha_{2},\dots,\alpha_{n})=a,$
then we say that $G$ is $\mathbb{Z}$-universal.
The purpose of this section is to investigate binary and ternary $\mathbb{Z}$-universal forms. We begin by proving one implication in statements (i) and (ii) of Theorem 1.2.
Proposition 5.2.
Let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}_{\geq 2}$ is squarefree.
  1. (i)
    If $D\equiv 2,3\pmod{4}$ and there exists a positive definite binary generalized form over $K$ which is $\mathbb{Z}$-universal, then $D\in\{2,3,6,7,10\}$.
  2. (ii)
    If $D\equiv 1\pmod{4}$ and there exists a classical positive definite binary generalized form over $K$ which is $\mathbb{Z}$-universal, then $D=5$.
Proof.
Suppose that $G$ is a positive definite binary generalized form over $K$ which is $\mathbb{Z}$-universal. If $D\equiv 1\pmod{4}$, we further assume that $G$ is classical. If $Q$ is the associated quaternary quadratic form defined by (4), then $Q$ is positive definite and universal.
By Lemma 4.5, $Q$ is also classical. Bhargava obtained, as a corollary of his proof of the 15-Theorem, a complete list of classical quaternary positive definite universal quadratic forms up to equivalence [Bh, Table 5]. Every form on this list has determinant $\leq 112$. By Corollary 4.3, $D^{2}\mid\det(M_{Q})$, thus $D^{2}\leq 112$. This yields the proof. ∎
Lemma 5.3.
Let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}\setminus\{0,1\}$ is squarefree and let $G$ be a binary generalized form over $K$ given by the matrix
(5) $M_{G}=\begin{pmatrix}a&\frac{b}{2}&\frac{d}{2}&\frac{e}{2}\\ \frac{b}{2}&c&\frac{\tau(e)}{2}&\frac{f}{2}\\ \frac{d}{2}&\frac{\tau(e)}{2}&\tau(a)&\frac{\tau(b)}{2}\\ \frac{e}{2}&\frac{f}{2}&\frac{\tau(b)}{2}&\tau(c)\end{pmatrix}$
where $a,b,c,e\in K$ and $d,f\in\mathbb{Q}$. Let $a=a_{1}+a_{2}\omega_{D}$, $b=b_{1}+b_{2}\omega_{D}$, $c=c_{1}+c_{2}\omega_{D}$, and $e=e_{1}+e_{2}\omega_{D}$. Let $Q$ be the quaternary quadratic form associated to $G$ and $M_{Q}$ be the matrix of $Q$.
If $D\equiv 2,3\pmod{4}$, then
$M_{Q}=\begin{pmatrix}2a_{1}+d&2Da_{2}&b_{1}+e_{1}&D(b_{2}-e_{2})\\ 2Da_{2}&D(2a_{1}-d)&D(b_{2}+e_{2})&D(b_{1}-e_{1})\\ b_{1}+e_{1}&D(b_{2}+e_{2})&2c_{1}+f&2Dc_{2}\\ D(b_{2}-e_{2})&D(b_{1}-e_{1})&2Dc_{2}&D(2c_{1}-f)\end{pmatrix}.$
If $D\equiv 1\pmod{4}$, then
$M_{Q}=\begin{pmatrix}M_{Q,1}&M_{Q,2}\\ M_{Q,2}^{\top}&M_{Q,3}\end{pmatrix}$
where
$M_{Q,1}$ $=\begin{pmatrix}2a_{1}+a_{2}+d&a_{1}+\frac{1+D}{2}a_{2}+\frac{d}{2}\\ a_{1}+\frac{1+D}{2}a_{2}+\frac{d}{2}&\frac{1+D}{2}a_{1}+\frac{1+3D}{4}a_{2}+\frac{1-D}{4}d\end{pmatrix},$
$M_{Q,2}$ $=\begin{pmatrix}b_{1}+\frac{b_{2}}{2}+e_{1}+\frac{e_{2}}{2}&\frac{b_{1}}{2}+\frac{1+D}{4}b_{2}+\frac{e_{1}}{2}+\frac{1-D}{4}e_{2}\\ \frac{b_{1}}{2}+\frac{1+D}{4}b_{2}+\frac{e_{1}}{2}+\frac{1+D}{4}e_{2}&\frac{1+D}{4}b_{1}+\frac{1+3D}{8}b_{2}+\frac{1-D}{4}e_{1}+\frac{1-D}{8}e_{2}\end{pmatrix},$
$M_{Q,3}$ $=\begin{pmatrix}2c_{1}+c_{2}+f&c_{1}+\frac{1+D}{2}c_{2}+\frac{f}{2}\\ c_{1}+\frac{1+D}{2}c_{2}+\frac{f}{2}&\frac{1+D}{2}c_{1}+\frac{1+3D}{4}c_{2}+\frac{1-D}{4}f\end{pmatrix}.$
Proof.
The matrix of $G$ is
$M_{G}=\begin{pmatrix}\alpha_{11}&\frac{\alpha_{12}}{2}&\frac{\beta_{11}}{2}&\frac{\beta_{12}}{2}\\ \frac{\alpha_{12}}{2}&\alpha_{22}&\frac{\tau(\beta_{12})}{2}&\frac{\beta_{22}}{2}\\ \frac{\beta_{11}}{2}&\frac{\tau(\beta_{12})}{2}&\tau(\alpha_{11})&\frac{\tau(\alpha_{12})}{2}\\ \frac{\beta_{12}}{2}&\frac{\beta_{22}}{2}&\frac{\tau(\alpha_{12})}{2}&\tau(\alpha_{22})\end{pmatrix}$
where $\alpha_{11}=a$, $\alpha_{12}=b$, $\alpha_{22}=c$, $\beta_{11}=d$, $\beta_{12}=e$, and $\beta_{22}=f$. In the notation of Lemma 4.6, $\alpha_{ij}=r_{ij}+s_{ij}\omega_{D}$ and $\beta_{ij}=t_{ij}+u_{ij}\omega_{D}$, hence $r_{11}=a_{1}$, $s_{11}=a_{2}$, $r_{12}=b_{1}$, $s_{12}=b_{2}$, $r_{22}=c_{1}$, $s_{22}=c_{2}$, $t_{11}=d$, $t_{12}=e_{1}$, $u_{12}=e_{2}$, and $t_{22}=f$.
The matrix of $Q$ is
$M_{Q}=\begin{pmatrix}a_{11}&\frac{b_{11}}{2}&\frac{a_{12}}{2}&\frac{b_{12}}{2}\\ \frac{b_{11}}{2}&c_{11}&\frac{b_{21}}{2}&\frac{c_{12}}{2}\\ \frac{a_{12}}{2}&\frac{b_{21}}{2}&a_{22}&\frac{b_{22}}{2}\\ \frac{b_{12}}{2}&\frac{c_{12}}{2}&\frac{b_{22}}{2}&c_{22}\end{pmatrix}.$
If $D\equiv 2,3\pmod{4}$, then by Lemma 4.6,
$a_{11}$ $=2r_{11}+t_{11}=2a_{1}+d,$
$b_{11}$ $=4Ds_{11}=4Da_{2},$
$c_{11}$ $=2Dr_{11}-Dt_{11}=D(2a_{1}-d),$
$a_{12}$ $=2r_{12}+2t_{12}=2(b_{1}+e_{1}),$
$b_{12}$ $=2Ds_{12}-2Du_{12}=2D(b_{2}-e_{2}),$
$b_{21}$ $=2Ds_{12}+2Du_{12}=2D(b_{2}+e_{2}),$
$c_{12}$ $=2Dr_{12}-2Dt_{12}=2D(b_{1}-e_{1}),$
$a_{22}$ $=2r_{22}+t_{22}=2c_{1}+f,$
$b_{22}$ $=4Ds_{22}=4Dc_{2},$
$c_{22}$ $=2Dr_{22}-Dt_{22}=D(2c_{1}-f).$
If $D\equiv 1\pmod{4}$, then by Lemma 4.6,
$a_{11}$ $=2r_{11}+s_{11}+t_{11}=2a_{1}+a_{2}+d,$
$b_{11}$ $=2r_{11}+(1+D)s_{11}+t_{11}=2a_{1}+(1+D)a_{2}+d,$
$c_{11}$ $=\frac{1+D}{2}r_{11}+\frac{1+3D}{4}s_{11}+\frac{1-D}{4}t_{11}=\frac{1+D}{2}a_{1}+\frac{1+3D}{4}a_{2}+\frac{1-D}{4}d,$
$a_{12}$ $=2r_{12}+s_{12}+2t_{12}+u_{12}=2b_{1}+b_{2}+2e_{1}+e_{2},$
$b_{12}$ $=r_{12}+\frac{1+D}{2}s_{12}+t_{12}+\frac{1-D}{2}u_{12}=b_{1}+\frac{1+D}{2}b_{2}+e_{1}+\frac{1-D}{2}e_{2},$
$b_{21}$ $=r_{12}+\frac{1+D}{2}s_{12}+t_{12}+\frac{1+D}{2}u_{12}=b_{1}+\frac{1+D}{2}b_{2}+e_{1}+\frac{1+D}{2}e_{2},$
$c_{12}$ $=\frac{1+D}{2}r_{12}+\frac{1+3D}{4}s_{12}+\frac{1-D}{2}t_{12}+\frac{1-D}{4}u_{12}=\frac{1+D}{2}b_{1}+\frac{1+3D}{4}b_{2}+\frac{1-D}{2}e_{1}+\frac{1-D}{4}e_{2},$
$a_{22}$ $=2r_{22}+s_{22}+t_{22}=2c_{1}+c_{2}+f,$
$b_{22}$ $=2r_{22}+(1+D)s_{22}+t_{22}=2c_{1}+(1+D)c_{2}+f,$
$c_{22}$ $=\frac{1+D}{2}r_{22}+\frac{1+3D}{4}s_{22}+\frac{1-D}{4}t_{22}=\frac{1+D}{2}c_{1}+\frac{1+3D}{4}c_{2}+\frac{1-D}{4}f.\qed$
In Lemmas 5.4 to 5.9, we find an example of a $\mathbb{Z}$-universal generalized form over $\mathbb{Q}(\sqrt{D})$ for $D\in\{2,3,6,7,10\}$ and of a classical $\mathbb{Z}$-universal generalized form over $\mathbb{Q}(\sqrt{5})$.
Lemma 5.4.
Let $K=\mathbb{Q}(\sqrt{2})$ and
$G_{2}(z,w)=z^{2}-\sqrt{2}zw+w^{2}-z\tau(z)-w\tau(w)+\tau(z)^{2}+\sqrt{2}\tau(z)\tau(w)+\tau(w)^{2}.$
The generalized form $G_{2}$ is $\mathbb{Z}$-universal over $K$.
Proof.
The coefficients of $G_{2}$ are $a=1$, $b=-\sqrt{2}$, $c=1$, $d=-1$, $e=0$, $f=-1$. Let $Q_{2}$ be the quaternary quadratic form $Q_{2}(x_{1},y_{1},x_{2},y_{2})=G_{2}(x_{1}+y_{1}\sqrt{2},x_{2}+y_{2}\sqrt{2})$ and let $M_{Q_{2}}$ be the matrix of $Q_{2}$. By Lemma 5.3 with $D=2$,
$M_{Q_{2}}=\begin{pmatrix}2a_{1}+d&4a_{2}&b_{1}+e_{1}&2(b_{2}-e_{2})\\ 4a_{2}&2(2a_{1}-d)&2(b_{2}+e_{2})&2(b_{1}-e_{1})\\ b_{1}+e_{1}&2(b_{2}+e_{2})&2c_{1}+f&4c_{2}\\ 2(b_{2}-e_{2})&2(b_{1}-e_{1})&4c_{2}&2(2c_{1}-f)\end{pmatrix}=\begin{pmatrix}1&0&0&-2\\ 0&6&-2&0\\ 0&-2&1&0\\ -2&0&0&6\end{pmatrix}.$
We get
$M_{Q_{2}}\sim\begin{pmatrix}1&0&0&0\\ 0&6&-2&0\\ 0&-2&1&0\\ 0&0&0&2\end{pmatrix}\sim\begin{pmatrix}1&0&0&0\\ 0&2&0&0\\ 0&0&1&0\\ 0&0&0&2\end{pmatrix}.$
Since the last matrix is positive definite, $G_{2}$ is also positive definite. Moreover, the diagonal quadratic form $x^{2}+2y^{2}+z^{2}+2w^{2}$ is universal by the 15-Theorem, hence $G_{2}$ is $\mathbb{Z}$-universal. ∎
Lemma 5.5.
Let $K=\mathbb{Q}(\sqrt{3})$ and
$G_{3}(z,w)=2z^{2}-2\sqrt{3}zw+2w^{2}-3z\tau(z)-\sqrt{3}z\tau(w)+\sqrt{3}w\tau(z)+2\tau(z)^{2}+2\sqrt{3}\tau(z)\tau(w)+2\tau(w)^{2}.$
The generalized form $G_{3}$ is $\mathbb{Z}$-universal over $K$.
Proof.
The coefficients of $G_{3}$ are $a=2$, $b=-2\sqrt{3}$, $c=2$, $d=-3$, $e=-\sqrt{3}$, $f=0$. Let $Q_{3}$ be the quaternary quadratic form $Q_{3}(x_{1},y_{1},x_{2},y_{2})=G_{3}(x_{1}+y_{1}\sqrt{3},x_{2}+y_{2}\sqrt{3})$. By Lemma 5.3 with $D=3$,
$M_{Q_{3}}=\begin{pmatrix}2a_{1}+d&6a_{2}&b_{1}+e_{1}&3(b_{2}-e_{2})\\ 6a_{2}&3(2a_{1}-d)&3(b_{2}+e_{2})&3(b_{1}-e_{1})\\ b_{1}+e_{1}&3(b_{2}-e_{2})&2c_{1}+f&6c_{2}\\ 3(b_{2}-e_{2})&3(b_{1}-e_{1})&6c_{2}&3(2c_{1}-f)\end{pmatrix}=\begin{pmatrix}1&0&0&-3\\ 0&21&-9&0\\ 0&-9&4&0\\ -3&0&0&12\end{pmatrix}.$
We get
$M_{Q_{3}}\sim\begin{pmatrix}1&0&0&0\\ 0&21&-9&0\\ 0&-9&4&0\\ 0&0&0&3\end{pmatrix}\sim\begin{pmatrix}1&0&0&0\\ 0&1&-1&0\\ 0&-1&4&0\\ 0&0&0&3\end{pmatrix}\sim\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&3&0\\ 0&0&0&3\end{pmatrix},$
The diagonal quadratic form $x^{2}+y^{2}+3z^{2}+3w^{2}$ is positive definite and universal by the 15-Theorem, hence $G_{3}$ is also positive definite and $\mathbb{Z}$-universal. ∎
Lemma 5.6.
Let $K=\mathbb{Q}(\sqrt{5})$ and
$G_{5}(z,w)=(3-\omega_{5})z^{2}+w^{2}-4z\tau(z)+(3-\tau(\omega_{5}))\tau(z)^{2}+\tau(w)^{2}$
where $\omega_{5}=\frac{1+\sqrt{5}}{2}$. The generalized form $G_{5}$ is classical and $\mathbb{Z}$-universal over $K$.
Proof.
The coefficients of $G_{5}$ are $a=3-\omega_{5}$, $b=0$, $c=1$, $d=-4$, $e=0$, and $f=0$. Since the coefficients of the off-diagonal terms are even, $G_{5}$ is classical. Let $Q_{5}$ be the quaternary quadratic form $Q_{5}(x_{1},y_{1},x_{2},y_{2})=G_{5}(x_{1}+y_{1}\omega_{5},x_{2}+y_{2}\omega_{5})$. By Lemma 5.3 with $D=5$,
$M_{Q_{5}}=\begin{pmatrix}M_{Q_{5},1}&M_{Q_{5},2}\\ M_{Q_{5},2}^{\top}&M_{Q_{5},3}\end{pmatrix}$
where
$M_{Q_{5},1}$ $=\begin{pmatrix}2a_{1}+a_{2}+d&a_{1}+3a_{2}+\frac{d}{2}\\ a_{1}+3a_{2}+\frac{d}{2}&3a_{1}+4a_{2}-d\end{pmatrix}=\begin{pmatrix}1&-2\\ -2&9\end{pmatrix},$
$M_{Q_{5},2}$ $=\begin{pmatrix}b_{1}+\frac{b_{2}}{2}+e_{1}+\frac{e_{2}}{2}&\frac{b_{1}}{2}+\frac{3}{2}b_{2}+\frac{e_{1}}{2}-e_{2}\\ \frac{b_{1}}{2}+\frac{3}{2}b_{2}+\frac{e_{1}}{2}+\frac{3}{2}e_{2}&\frac{3}{2}b_{1}+2b_{2}-e_{1}-\frac{1}{2}e_{2}\end{pmatrix}=\begin{pmatrix}0&0\\ 0&0\end{pmatrix},$
$M_{Q_{5},3}$ $=\begin{pmatrix}2c_{1}+c_{2}+f&c_{1}+3c_{2}+\frac{f}{2}\\ c_{1}+3c_{2}+\frac{f}{2}&3c_{1}+4c_{2}-f\end{pmatrix}=\begin{pmatrix}2&1\\ 1&3\end{pmatrix},$
thus
$M_{Q_{5}}=\begin{pmatrix}1&-2&0&0\\ -2&9&0&0\\ 0&0&2&1\\ 0&0&1&3\end{pmatrix}\sim\begin{pmatrix}1&0&0&0\\ 0&5&0&0\\ 0&0&2&1\\ 0&0&1&3\end{pmatrix}.$
This quadratic form is positive definite and universal by the 15-Theorem. Alternatively, it is possible to consult [Bh, Table 5], which contains the equivalent form
$\begin{pmatrix}1&0&0&0\\ 0&2&1&0\\ 0&1&3&0\\ 0&0&0&5\end{pmatrix}$
encoded as “25: 2 3 5 0 0 2:2523500225:\ 2\ 3\ 5\ 0\ 0\ 225 : 2 3 5 0 0 2”. ∎
Lemma 5.7.
Let $K=\mathbb{Q}(\sqrt{6})$ and $G_{6}$ be the generalized form with the matrix
$M_{G_{6}}=\begin{pmatrix}3&\sqrt{6}&-\frac{5}{2}&\frac{\sqrt{6}}{2}\\ \sqrt{6}&3&-\frac{\sqrt{6}}{2}&-\frac{1}{2}\\ -\frac{5}{2}&-\frac{\sqrt{6}}{2}&3&-\sqrt{6}\\ \frac{\sqrt{6}}{2}&-\frac{1}{2}&-\sqrt{6}&3\end{pmatrix}.$
The generalized form $G_{6}$ is $\mathbb{Z}$-universal over $K$.
Proof.
The coefficients of $G_{6}$ are $a=3$, $b=2\sqrt{6}$, $c=3$, $d=-5$, $e=\sqrt{6}$, and $f=-1$. Let $Q_{6}$ be the quaternary quadratic form $M_{Q_{6}}(x_{1},y_{1},x_{2},y_{2})=G_{6}(x_{1}+y_{1}\sqrt{6},x_{2}+y_{2}\sqrt{6})$. By Lemma 5.3 with $D=6$,
$M_{Q_{6}}=\begin{pmatrix}2a_{1}+d&12a_{2}&b_{1}+e_{1}&6(b_{2}-e_{2})\\ 12a_{2}&6(2a_{1}-d)&6(b_{2}+e_{2})&6(b_{1}-e_{1})\\ b_{1}+e_{1}&6(b_{2}+e_{2})&2c_{1}+f&12c_{2}\\ 6(b_{2}-e_{2})&6(b_{1}-e_{1})&12c_{2}&6(2c_{1}-f)\end{pmatrix}=\begin{pmatrix}1&0&0&6\\ 0&66&18&0\\ 0&18&5&0\\ 6&0&0&42\end{pmatrix}.$
We get
$M_{Q_{6}}\sim\begin{pmatrix}1&0&0&0\\ 0&66&18&0\\ 0&18&5&0\\ 0&0&0&6\end{pmatrix}\sim\begin{pmatrix}1&0&0&0\\ 0&3&3&0\\ 0&3&5&0\\ 0&0&0&6\end{pmatrix}\sim\begin{pmatrix}1&0&0&0\\ 0&3&0&0\\ 0&0&2&0\\ 0&0&0&6\end{pmatrix}.$
The last quadratic form is positive definite and universal by the 15-Theorem. ∎
Lemma 5.8.
Let $K=\mathbb{Q}(\sqrt{7})$ and $G_{7}$ be the generalized quadratic form with the matrix
$M_{G_{7}}=\begin{pmatrix}3&\frac{1+2\sqrt{7}}{2}&-\frac{5}{2}&\frac{-1+2\sqrt{7}}{2}\\ \frac{1+2\sqrt{7}}{2}&4&\frac{-1-2\sqrt{7}}{2}&\frac{3}{2}\\ -\frac{5}{2}&\frac{-1-2\sqrt{7}}{2}&3&\frac{1-2\sqrt{7}}{2}\\ \frac{-1+2\sqrt{7}}{2}&\frac{3}{2}&\frac{1-2\sqrt{7}}{2}&4\end{pmatrix}.$
The generalized form $G_{7}$ is $\mathbb{Z}$-universal over $K$.
Proof.
The coefficients of $G_{7}$ are $a=3$, $b=1+2\sqrt{7}$, $c=4$, $d=-5$, $e=-1+2\sqrt{7}$, and $f=3$. By Lemma 5.3 with $D=7$,
$M_{Q_{7}}=\begin{pmatrix}2a_{1}+d&14a_{2}&b_{1}+e_{1}&7(b_{2}-e_{2})\\ 14a_{2}&7(2a_{1}-d)&7(b_{2}+e_{2})&7(b_{1}-e_{1})\\ b_{1}+e_{1}&7(b_{2}+e_{2})&2c_{1}+f&14c_{2}\\ 7(b_{2}-e_{2})&7(b_{1}-e_{1})&14c_{2}&7(2c_{1}-f)\end{pmatrix}=\begin{pmatrix}1&0&0&0\\ 0&77&28&14\\ 0&28&11&0\\ 0&14&0&35\end{pmatrix}.$
Let
$A_{7}=\begin{pmatrix}1&0&0&0\\ 0&2&-3&3\\ 0&-5&8&-7\\ 0&-1&1&-1\end{pmatrix}.$
It can be verified that $\det(A_{7})=1$, hence $A_{7}\in\operatorname{SL}_{4}(\mathbb{Z})$. The equivalent form with the matrix
$A_{7}^{\top}M_{Q_{7}}A_{7}=\begin{pmatrix}1&0&0&0\\ 0&2&-3&3\\ 0&-5&8&-7\\ 0&-1&1&-1\end{pmatrix}^{\top}\begin{pmatrix}1&0&0&0\\ 0&77&28&14\\ 0&28&11&0\\ 0&14&0&35\end{pmatrix}\begin{pmatrix}1&0&0&0\\ 0&2&-3&3\\ 0&-5&8&-7\\ 0&-1&1&-1\end{pmatrix}=\begin{pmatrix}1&0&0&0\\ 0&2&1&0\\ 0&1&4&0\\ 0&0&0&7\end{pmatrix}$
is positive definite and universal by the 15-Theorem. Alternatively, it can be found in [Bh, Table 5] as “49: 2 4 7 0 0 2”. ∎
Lemma 5.9.
Let $K=\mathbb{Q}(\sqrt{10})$. If $G_{10}$ is the generalized quadratic form with the matrix
$M_{G_{10}}=\begin{pmatrix}3&\frac{\sqrt{10}}{2}&-\frac{5}{2}&\frac{\sqrt{10}}{2}\\ \frac{\sqrt{10}}{2}&5&-\frac{\sqrt{10}}{2}&\frac{7}{2}\\ -\frac{5}{2}&-\frac{\sqrt{10}}{2}&3&\frac{-\sqrt{10}}{2}\\ \frac{\sqrt{10}}{2}&\frac{7}{2}&\frac{\sqrt{10}}{2}&5\end{pmatrix},$
then $G_{10}$ is $\mathbb{Z}$-universal over $K$.
Proof.
The coefficients of $G_{10}$ are $a=3$, $b=\sqrt{10}$, $c=5$, $d=-5$, $e=\sqrt{10}$, and $f=7$. Let $Q_{10}$ be the quaternary quadratic form $Q_{10}(x_{1},y_{1},x_{2},y_{2})=G_{10}(x_{1}+y_{1}\sqrt{10},x_{2}+y_{2}\sqrt{10})$. By Lemma 5.3 with $D=10$,
$M_{Q_{10}}=\begin{pmatrix}2a_{1}+d&20a_{2}&b_{1}+e_{1}&10(b_{2}-e_{2})\\ 20a_{2}&10(2a_{1}-d)&10(b_{2}+e_{2})&10(b_{1}-e_{1})\\ b_{1}+e_{1}&10(b_{2}+e_{2})&2c_{1}+f&20c_{2}\\ 10(b_{2}-e_{2})&10(b_{1}-e_{1})&20c_{2}&10(2c_{1}-f)\end{pmatrix}=\begin{pmatrix}1&0&0&0\\ 0&110&20&0\\ 0&20&17&20\\ 0&0&20&30\end{pmatrix}.$
Let
$A_{10}=\begin{pmatrix}1&0&0&0\\ 0&1&1&2\\ 0&-6&-5&-10\\ 0&4&3&7\end{pmatrix}.$
It can be verified that $\det(A_{10})=1$, hence $A\in\operatorname{SL}_{4}(\mathbb{Z})$. The equivalent form with the matrix
$A_{10}^{\top}M_{Q_{10}}A_{10}=\begin{pmatrix}1&0&0&0\\ 0&1&1&2\\ 0&-6&-5&-10\\ 0&4&3&7\end{pmatrix}^{\top}\begin{pmatrix}1&0&0&0\\ 0&110&20&0\\ 0&20&17&20\\ 0&0&20&30\end{pmatrix}\begin{pmatrix}1&0&0&0\\ 0&1&1&2\\ 0&-6&-5&-10\\ 0&4&3&7\end{pmatrix}=\begin{pmatrix}1&0&0&0\\ 0&2&0&0\\ 0&0&5&0\\ 0&0&0&10\end{pmatrix}$
is positive definite and universal by the 15-Theorem. ∎
Now we are ready to prove Theorem 1.2, which we restate here for the convenience of the reader.
Theorem 5.10.
Let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}_{\geq 2}$ is squarefree.
  1. i)
    Assume that $D\equiv 2,3\pmod{4}$. A binary $\mathbb{Z}$-universal generalized quadratic form exists over $K$ if and only if $D\in\{2,3,6,7,10\}$.
  2. ii)
    Assume that $D\equiv 1\pmod{4}$. A classical binary $\mathbb{Z}$-universal generalized quadratic form exists over $K$ if and only if $D=5$.
Proof.
First we prove (i). We assume that $D\equiv 2,3\pmod{4}$. If a binary $\mathbb{Z}$-universal generalized form exists over $K=\mathbb{Q}(\sqrt{D})$, then $D\in\{2,3,6,7,10\}$ by Proposition 5.2 (i). Conversely, in Lemmas 5.4, 5.5, 5.7, 5.8 and 5.9 we found an example of a generalized form with the required properties over each of these fields.
Next, we prove (ii). We assume that $D\equiv 1\pmod{4}$. If a classical binary $\mathbb{Z}$-universal generalized form exists over $K=\mathbb{Q}(\sqrt{D})$, then $D=5$ by Proposition 5.2 (ii). We found an example of such a form over $\mathbb{Q}(\sqrt{5})$ in Lemma 5.6. ∎
Theorem 1.3 is an immediate consequence of the following slightly refined version.
Theorem 5.11.
Let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}_{\geq 2}$ is squarefree. Assume that one of the following two conditions is satisfied:
  1. i)
    $D\equiv 2,3\pmod{4}$ and there exists a ternary $\mathbb{Z}$-universal generalized quadratic form over $K$,
  2. ii)
    $D\equiv 1\pmod{4}$ and there exists a classical ternary $\mathbb{Z}$-universal generalized quadratic form over $K$.
Then $D\in\{1,\dots,62\}\cup\{66,70,74,77,78,82,85,86,87,93,94,95,102,106,110\}$.
Proof.
Let $G$ be a ternary generalized quadratic form over $K$ which is $\mathbb{Z}$-universal. If $D\equiv 1\pmod{4}$, then we further assume that $G$ is classical. Let $Q$ be the quadratic form in $6$ variables associated to $G$ and let $M_{Q}$ be the matrix of $Q$. By Lemma 4.5, $Q$ is classical.
A classical universal quadratic form has a quaternary subform (not necessarily universal) obtained by the method of escalations [Bh, p. 31]. Let $R$ be such a subform of $Q$ and $M_{R}$ the matrix of $R$. If $p\mid D$ is a prime, then $\operatorname{rank}(M_{Q}\bmod{p})\leq 3$ by Proposition 4.7. The rank of $R$ over $\mathbb{Z}/p\mathbb{Z}$ is less than or equal to the rank of $Q$ over $\mathbb{Z}/p\mathbb{Z}$, hence $\operatorname{rank}(M_{R}\bmod{p})\leq 3$. Since $M_{R}$ is a $4\times 4$ matrix, $\det(M_{R})\equiv 0\pmod{p}$.
This proves $D\mid\det(M_{R})$. Every quaternary form obtained by the method of escalations is listed in [Bh, Table 3] and in particular, has determinant $\leq 112$. It remains to determine which squarefree $D\leq 112$ divide the determinant of some form on the list. Every squarefree $D\in\{1,\dots,62\}$ satisfies this condition. Additionally, the squarefree determinants greater than $62$ on the list are $66,70,74,77,78,82,85,86,87,93,94,95,102,106,110$. ∎
Using the 290-Theorem, we can remove the “classical” assumption in the case $D\equiv 1\pmod{4}$.
Theorem 5.12.
There are only finitely many real quadratic fields which admit a binary or ternary $\mathbb{Z}$-universal generalized quadratic form.
Proof.
We show that there exists a positive constant $C$ such that $D\leq C$ for every such field $K=\mathbb{Q}(\sqrt{D})$. If $D\equiv 2,3\pmod{4}$, then the statement follows from Theorems 5.10 and 5.11. Hence, we can assume that $D\equiv 1\pmod{4}$. Let $G$ be a $\mathbb{Z}$-universal generalized form over $K$ with matrix $M_{G}$ and $Q$ be the quadratic form associated to $G$ with matrix $M_{Q}$.
First, assume that $G$ is binary. If $M_{G}$ is the matrix of $G$ and $M_{Q}$ is the matrix of $Q$, then $\det(M_{Q})=D^{2}\det(M_{G})$ by Lemma 4.2. The matrix $M_{G}$ is a $4\times 4$ matrix with entries in $\frac{1}{2}\mathcal{O}_{K}$, hence $16\det(M_{G})\in\mathcal{O}_{K}\cap\mathbb{Q}=\mathbb{Z}$, and we get $D^{2}\mid 16\det(M_{Q})$.
A corollary of the 290-Theorem is that there are exactly 6436 universal quaternary forms [BH, Theorem 4]. If we let $m$ be the maximum of their determinants, then $D^{2}\leq 16m$.
Secondly, assume that $G$ is ternary. A universal quadratic form contains a quaternary subform equivalent to one of $6560$ basic escalators [BH, p. 4]. Let $R$ be such a subform of $Q$ and $M_{R}$ the matrix of $R$. If $p\mid D$ is a prime, then $\operatorname{rank}(M_{Q}\bmod{p})\leq 3$ by Proposition 4.7, hence $\operatorname{rank}(M_{R}\bmod{p})\leq 3$ as well. Since $M_{R}$ is a $4\times 4$ matrix, $\det(M_{R})\equiv 0\pmod{p}$, and since the entries of $M_{R}$ are in $\frac{1}{2}\mathbb{Z}$, we have $16\det(M_{R})\in\mathbb{Z}$.
This proves $D\mid 16\det(M_{R})$. If we let $m$ be the maximum of the determinants of the basic escalators, then $D\leq 16m$. ∎
In contrast to Theorem 5.12, it is not difficult to construct a $\mathbb{Z}$-universal generalized form in $4$ variables over every real quadratic field.
Proposition 5.13.
Let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}_{\geq 2}$ is squarefree. The generalized quadratic form
$G(z_{1},z_{2},z_{3},z_{4})$ $=z_{1}^{2}-z_{1}\tau(z_{1})+\tau(z_{1})^{2}+z_{2}^{2}-z_{2}\tau(z_{2})+\tau(z_{2})^{2}$
$+z_{3}^{2}-z_{3}\tau(z_{3})+\tau(z_{3})^{2}+z_{4}^{2}-z_{4}\tau(z_{4})+\tau(z_{4})^{2}$
is $\mathbb{Z}$-universal over $K$. Thus, the minimal rank of a $\mathbb{Z}$-universal generalized form over $K$ is at most $4$.
Proof.
Clearly, $G$ is positive definite. It is enough to show that $G$ represents every $a\in\mathbb{Z}_{\geq 1}$ over $\mathbb{Z}$. If $z_{i}\in\mathbb{Z}$ for $i=1,\dots,4$, then $\tau(z_{i})=z_{i}$, hence
$G(z_{1},z_{2},z_{3},z_{4})=z_{1}^{2}+z_{2}^{2}+z_{3}^{2}+z_{4}^{2}.$
This quadratic form is universal by Lagrange’s Four Square Theorem. ∎
We end this section with a proposition that will be used in Section 7.
Proposition 5.14.
Let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}\setminus\{0,1\}$ is squarefree, $D\equiv 2,3\pmod{4}$. Let $G$ be an integral binary generalized form over $K$. Assume that $G$ is $\mathbb{Z}$-valued. If $Q$ is the quaternary quadratic form associated to $G$ and $M_{Q}$ is the matrix of $Q$, then
$\frac{\det(M_{Q})}{D^{2}}\equiv 0,1\pmod{4}.$
Proof.
By Lemma 4.2,
$\det(M_{Q})=D^{2}\cdot\det(2M_{G}).$
The matrix $2M_{G}$ is
$2M_{G}=\begin{pmatrix}2a&b&d&e\\ b&2c&\tau(e)&f\\ d&\tau(e)&2\tau(a)&\tau(b)\\ e&f&\tau(b)&2\tau(c)\end{pmatrix}$
where $a,b,c,e\in\mathcal{O}_{K}$ and $d,f\in\mathbb{Z}$ because $G$ is integral. Expanding $\det(2M_{G})$ along the first column, we get
$\det(2M_{G})=2a\begin{vmatrix}2c&\tau(e)&f\\ \tau(e)&2\tau(a)&\tau(b)\\ f&\tau(b)&2\tau(c)\end{vmatrix}-b\begin{vmatrix}b&d&e\\ \tau(e)&2\tau(a)&\tau(b)\\ f&\tau(b)&2\tau(c)\end{vmatrix}+d\begin{vmatrix}b&d&e\\ 2c&\tau(e)&f\\ f&\tau(b)&2\tau(c)\end{vmatrix}-e\begin{vmatrix}b&d&e\\ 2c&\tau(e)&f\\ \tau(e)&2\tau(a)&\tau(b)\end{vmatrix},$
hence
$\det(2M_{G})\equiv$ $2a\left[\tau(e)\tau(b)f+f\tau(e)\tau(b)\right]-b\left[d\tau(b)f+e\tau(e)\tau(b)-2f\tau(a)e-\tau(b)^{2}b-2\tau(c)\tau(e)d\right]$
$+d\left[2b\tau(e)\tau(c)+df^{2}+2ec\tau(b)-f\tau(e)e-\tau(b)fb\right]$
$-e\left[b\tau(e)\tau(b)+df\tau(e)-\tau(e)^{2}e-2\tau(a)fb-2\tau(b)cd\right]$
$\equiv$ $-\operatorname{N}(b)df-\operatorname{N}(b)\operatorname{N}(e)+2\tau(a)bef+\operatorname{N}(b)^{2}+2b\tau(c)d\tau(e)$
$+2b\tau(c)d\tau(e)+d^{2}f^{2}+2\tau(b)cde-d\operatorname{N}(e)f-\operatorname{N}(b)df$
$-\operatorname{N}(b)\operatorname{N}(e)-d\operatorname{N}(e)f+\operatorname{N}(e)^{2}+2\tau(a)bef+2\tau(b)cde$
$\equiv$ $-2\operatorname{N}(b)df-2\operatorname{N}(b)\operatorname{N}(e)+\operatorname{N}(b)^{2}+d^{2}f^{2}-2d\operatorname{N}(e)f+\operatorname{N}(e)^{2}$
$\equiv$ $\left(\operatorname{N}(b)+\operatorname{N}(e)+df\right)^{2}\pmod{4}.$
Thus,
$\frac{\det(M_{Q})}{D^{2}}=\det(2M_{G})\equiv 0,1\pmod{4}.\qed$
6. Indefinite generalized quadratic forms
This section focuses on proving the universality of $G(z,w)=z\tau(z)+w\tau(w)$ over real quadratic fields, where the form is indefinite. One of the classical methods for studying representations by quadratic forms involves using the local-global principle. We will use the following version of the local-global principle for indefinite forms. If $p$ is a prime number, then $\mathbb{Z}_{p}$ denotes the ring of $p$-adic integers.
Theorem 6.1 ([Ca, Chap. 9, Thm. 1.5]).
Let $Q$ be a regular indefinite integral form in $n\geq 4$ variables and let $a\neq 0$ be an integer. Suppose that $a$ is represented by $Q$ over all $\mathbb{Z}_{p}$. Then $a$ is represented by $Q$ over $\mathbb{Z}$.
Recall that by our definition, every quadratic form $Q$ represents 00 (trivially). To represent nonzero $a$, we initially find the representation modulo $p$, followed by lifting it to a representation over $\mathbb{Z}_{p}$ using the following corollaries of Hensel’s lemma. The tuple $(x_{1},x_{2},\dots,x_{n})\in\mathbb{Z}_{p}^{n}$ is called primitive if at least one of the $x_{i}$ is invertible in $\mathbb{Z}_{p}$.
Lemma 6.2 ([Se, Chap. 2, Cor. 2]).
Suppose $p\neq 2$. Let $Q(x)=\sum a_{ij}x_{i}x_{j}$ with $a_{ij}=a_{ji}$ be a quadratic form with coefficients in $\mathbb{Z}_{p}$ whose determinant $\det(a_{ij})$ is invertible. Let $a\in\mathbb{Z}_{p}$. Every primitive solution of the equation $Q(x)\equiv a\pmod{p}$ lifts to a true solution.
Lemma 6.3 ([Se, Chap. 2, Cor. 3]).
Suppose $p=2$. Let $Q(x)=\sum a_{ij}x_{i}x_{j}$ with $a_{ij}=a_{ji}$ be a quadratic form with coefficients in $\mathbb{Z}_{2}$ and let $a\in\mathbb{Z}_{2}$. Let $x$ be a primitive solution of $Q(x)\equiv a\pmod{8}$. We can lift $x$ to a true solution provided $x$ does not annihilate all the $\frac{\partial Q}{\partial x_{j}}$ modulo $4$; this last condition is fulfilled if $\det(a_{ij})$ is invertible.
We proceed with proving that $x^{2}+y^{2}-Dz^{2}-Dw^{2}$ represents every $a\in\mathbb{Z}$.
Lemma 6.4.
Let $p\neq 2$ be a prime, $D\in\mathbb{Z}$ squarefree, and $a\in\mathbb{Z}_{p}$. The quadratic form $x^{2}+y^{2}-Dz^{2}-Dw^{2}$ represents $a$ over $\mathbb{Z}_{p}$.
Proof.
We can assume $a\neq 0$. First, suppose $a=p^{2k}a_{1}$ where $k\in\mathbb{Z}_{\geq 0}$ and $p\nmid a_{1}$. The congruence
$x^{2}+y^{2}\equiv a_{1}\pmod{p}$
has a solution, hence by Lemma 6.2, there exist $x_{1},y_{1}\in\mathbb{Z}_{p}$ such that
$x_{1}^{2}+y_{1}^{2}=a_{1}.$
The tuple $(x,y,z,w)=(p^{k}x_{1},p^{k}y_{1},0,0)$ is a solution of the equation $x^{2}+y^{2}-Dz^{2}-Dw^{2}=a$.
Secondly, suppose $a=p^{2k+1}a_{1}$ where $k\in\mathbb{Z}_{\geq 0}$ and $p\nmid a_{1}$. We distinguish two cases.
If $p\nmid D$, then the congruence
$x^{2}+y^{2}\equiv pa_{1}+D\pmod{p}$
has a solution, hence by Lemma 6.2, there exist $x_{1},y_{1}\in\mathbb{Z}_{p}$ such that
$x_{1}^{2}+y_{1}^{2}=pa_{1}+D.$
The tuple $(x,y,z,w)=(p^{k}x_{1},p^{k}y_{1},p^{k},0)$ is a solution of the original equation because
$x^{2}+y^{2}-Dz^{2}-Dw^{2}=p^{2k}(x_{1}^{2}+z_{1}^{2}-D)=p^{2k}\cdot pa_{1}=a.$
If $p\mid D$, let $D_{p}=\frac{D}{p}$. We have $p\nmid D_{p}$ because $D$ is squarefree. The congruence
$-D_{p}(z^{2}+w^{2})\equiv a_{1}\pmod{p}$
has a solution, hence by Lemma 6.2, there exist $z_{1},w_{1}\in\mathbb{Z}_{p}$ such that
$-D_{p}(z_{1}^{2}+w_{1}^{2})=a_{1}.$
The tuple $(x,y,z,w)=(0,0,p^{k}z_{1},p^{k}w_{1})$ is a solution because
$-Dz^{2}-Dw^{2}=p^{2k+1}(-D_{p})(z_{1}^{2}+w_{1}^{2})=p^{2k+1}a_{1}=a.\qed$
Lemma 6.5.
Let $D\in\mathbb{Z}$ be squarefree and $a\in\mathbb{Z}_{2}$. The quadratic form $x^{2}+y^{2}-Dz^{2}-Dw^{2}$ represents $a$ over $\mathbb{Z}_{2}$.
Proof.
We can assume $a\neq 0$. First, suppose that $4\nmid a$, hence $a\equiv 1,2,3,5,6,7\pmod{8}$. We find a solution to the congruence
$x^{2}+y^{2}-Dz^{2}-Dw^{2}\equiv a\pmod{8}$
where $x$ is odd, hence $\frac{\partial Q}{\partial x}=2x\neq 0\pmod{4}$. This solution lifts to a true solution in $\mathbb{Z}_{2}$ by Lemma 6.3.
Let
$(x,y)=\begin{cases}(1,0),&\text{if }a\equiv 1\pmod{8},\\ (1,1),&\text{if }a\equiv 2\pmod{8},\\ (1,2),&\text{if }a\equiv 5\pmod{8}.\end{cases}$
We have $x^{2}+y^{2}\equiv a\pmod{8}$ in each case.
If $a\equiv 3\pmod{8}$, let
$(z,w)=\begin{cases}(1,1),&\text{if }D\equiv 1,3,5\pmod{8},\\ (1,0),&\text{if }D\equiv 2,6,7\pmod{8},\end{cases}$
if $a\equiv 6\pmod{8}$, let
$(z,w)=\begin{cases}(2,0),&\text{if }D\equiv 1,3,5,7\pmod{8},\\ (1,1),&\text{if }D\equiv 2,6\pmod{8},\end{cases}$
and if $a\equiv 7\pmod{8}$, let
$(z,w)=\begin{cases}(1,1),&\text{if }D\equiv 1,3,5,7\pmod{8},\\ (1,0),&\text{if }D\equiv 2,6\pmod{8}.\end{cases}$
In each of the above cases, $a+Dz^{2}+Dw^{2}\pmod{8}\in\{1,2,5\}$, hence we can find $(x,y)=(1,y)$ such that
$x^{2}+y^{2}\equiv a+Dz^{2}+Dw^{2}\pmod{8}.$
Secondly, assume $4\mid a$. Let us write $a=4^{k}a_{1}$ where $4\nmid a_{1}$. By the first part of the proof, there exists a tuple $(x_{1},y_{1},z_{1},w_{1})\in\mathbb{Z}_{p}^{4}$ such that
$x_{1}^{2}+y_{1}^{2}+Dz_{1}^{2}+Dw_{1}^{2}=a_{1}$
and we set $(x,y,z,w)=(2^{k}x_{1},2^{k}y_{1},2^{k}z_{1},2^{k}w_{1})$. ∎
Theorem 6.6.
Let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}_{\geq 2}$ is squarefree. The binary generalized form
$G(z,w)=z\tau(z)+w\tau(w)$
represents every $a\in\mathbb{Z}$.
Proof.
It is sufficient to show that $G$ represents every $a\in\mathbb{Z}\setminus\{0\}$ over $\mathbb{Z}[\sqrt{D}]$ because $\mathbb{Z}[\sqrt{D}]\subset\mathcal{O}_{K}$.
Let $z=x_{1}+y_{1}\sqrt{D}$ and $w=x_{2}+y_{2}\sqrt{D}$ where $x_{1},y_{1},x_{2},y_{2}\in\mathbb{Z}$. Let
$Q(x_{1},y_{1},x_{2},y_{2})=G(x_{1}+y_{1}\sqrt{D},x_{2}+y_{2}\sqrt{D}).$
We have
$Q(x_{1},y_{1},x_{2},y_{2})=(x_{1}+y_{1}\sqrt{D})(x_{1}-y_{1}\sqrt{D})+(x_{2}+y_{2}\sqrt{D})(x_{2}-y_{2}\sqrt{D})=x_{1}^{2}+x_{2}^{2}-Dy_{1}^{2}-Dy_{2}^{2}.$
The form $Q$ represents $a$ over $\mathbb{Z}_{p}$ for $p$ odd by Lemma 6.4 and for $p=2$ by Lemma 6.5. By Theorem 6.1, it represents $a$ over $\mathbb{Z}$. ∎
Proposition 6.7.
Let $D\in\mathbb{Z}_{\geq 1}$. The quadratic form
$Q(x,y,z,w)=x^{2}+y^{2}-Dz^{2}-Dw^{2}$
represents 00 non-trivially if and only if $D$ is a sum of two squares.
Proof.
Assume that $D$ is a sum of two squares, say $D=a^{2}+b^{2}$ where $a,b\in\mathbb{Z}$. If $(x,y,z,w)=(a,b,1,0)$, then $Q(x,y,z,w)=0$.
Now we prove the converse under the assumption that $D$ is squarefree. Let $p\mid D$ be an odd prime and $(x,y,z,w)$ be a primitive solution to
$x^{2}+y^{2}-Dz^{2}-Dw^{2}=0.$
In particular, not all of $x$, $y$, $z$, $w$ are divisible by $p$.
First, suppose that $p\nmid x$ or $p\nmid y$. Without loss of generality, we can assume $p\nmid y$. Reducing the equation modulo $p$, we get $x^{2}+y^{2}\equiv 0\pmod{p}$, hence
$\left(\frac{x}{y}\right)^{2}\equiv-1\pmod{p}.$
This shows that $-1$ is a square modulo $p$, hence $p\equiv 1\pmod{4}$.
Secondly, suppose that $p\mid x$ and $p\mid y$. From the equation, we have $p^{2}\mid D(z^{2}+w^{2})$, and because $D$ is squarefree, $p\mid z^{2}+w^{2}$. Since $(x,y,z,w)$ is primitive, we get $p\nmid z$ or $p\nmid w$, and the congruence $z^{2}+w^{2}\equiv 0\pmod{p}$ implies $p\equiv 1\pmod{4}$. We proved that every odd prime $p\mid D$ satisfies $p\equiv 1\pmod{4}$, hence $D$ is a sum of two squares.
To prove the converse for general $D$, write $D=k^{2}d$ where $k,d\in\mathbb{Z}$ and $d$ is squarefree. If $(x_{0},y_{0},z_{0},w_{0})$ is a solution to
$x^{2}+y^{2}-k^{2}dz^{2}-k^{2}dw^{2}=0,$
then $(x_{0},y_{0},kz_{0},kw_{0})$ is a solution to
$x^{2}+y^{2}-dz^{2}-dw^{2}=0.$
Thus, if $Q$ represents 00, then $x^{2}+y^{2}-dz^{2}-dw^{2}$ also represents 00. By the first part of the proof, $d$ is a sum of two squares, hence $D$ is also a sum of two squares. ∎
7. Examples of universal binary generalized forms
7.1. Preliminaries
Let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}_{\geq 2}$ is squarefree, $G$ be a $\mathbb{Z}$-valued generalized form over $K$ in $n$ variables, and $Q$ be the quadratic form in $2n$ variables associated to $G$. By Lemma 4.5, if one of the following conditions is satisfied:
  1. i)
    $D\equiv 2,3\pmod{4}$ and $G$ is integral,
  2. ii)
    $D\equiv 1\pmod{4}$ and $G$ is classical,
then $Q$ is classical. By Theorem 1.2, a $\mathbb{Z}$-universal binary generalized form satisfying the conditions above exists over $K$ if and only if $D\in\{2,3,5,6,7,10\}$.
A complete list of universal classical positive definite quaternary quadratic forms (up to equivalence) is available [Bh, Table 5]. In this section, we answer the following question: Which of these quadratic forms are equivalent to a quadratic form $Q$ associated to $G$ satisfying the conditions above? For each $D\in\{2,3,5,6,7,10\}$ we identify all such forms.
We use Bhargava’s notation $D:\ a\ b\ c\ d\ e\ f$ for a ternary quadratic form with matrix
$M_{L}=\begin{pmatrix}a&\frac{f}{2}&\frac{e}{2}\\ \frac{f}{2}&b&\frac{d}{2}\\ \frac{e}{2}&\frac{d}{2}&c\end{pmatrix}$
of determinant $D$. If $L=D:\ a\ b\ c\ d\ e\ f$, then $1\oplus L$ denotes the quaternary form with matrix
$M_{1\oplus L}=\begin{pmatrix}1&0&0&0\\ 0&a&\frac{f}{2}&\frac{e}{2}\\ 0&\frac{f}{2}&b&\frac{d}{2}\\ 0&\frac{e}{2}&\frac{d}{2}&c\end{pmatrix}.$
We also need a concise way of representing a generalized binary form. Thus we let $[a,b,c,d,e,f]$ represent the binary generalized form with matrix
$\begin{pmatrix}a&\frac{b}{2}&\frac{d}{2}&\frac{e}{2}\\ \frac{b}{2}&c&\frac{\tau(e)}{2}&\frac{f}{2}\\ \frac{d}{2}&\frac{\tau(e)}{2}&\tau(a)&\frac{\tau(b)}{2}\\ \frac{e}{2}&\frac{f}{2}&\frac{\tau(b)}{2}&\tau(c)\end{pmatrix}$
where $a,b,c,e\in\mathcal{O}_{K}$ and $d,f\in\mathbb{Z}$. We also use $[a_{11},a_{12},a_{13},a_{14},a_{22},a_{23},a_{24},a_{33},a_{34},a_{44}]$ to represent the quaternary quadratic form with matrix
$\begin{pmatrix}a_{11}&\frac{a_{12}}{2}&\frac{a_{13}}{2}&\frac{a_{14}}{2}\\ \frac{a_{12}}{2}&a_{22}&\frac{a_{23}}{2}&\frac{a_{24}}{2}\\ \frac{a_{13}}{2}&\frac{a_{23}}{2}&a_{33}&\frac{a_{34}}{2}\\ \frac{a_{14}}{2}&\frac{a_{24}}{2}&\frac{a_{34}}{2}&a_{44}\end{pmatrix}.$
7.2. Universal quadratic forms associated to binary generalized forms for $D=2$
Let $G$ be a $\mathbb{Z}$-universal binary generalized form over $\mathbb{Q}(\sqrt{2})$, $Q$ be the associated quaternary quadratic form, and $M_{Q}$ be the matrix of $Q$. We have $4\mid\det(M_{Q})$ by Lemma 4.2 and $\frac{\det(M_{Q})}{4}\equiv 0,1\pmod{4}$ by Proposition 5.14. From the table of universal classical quaternary forms [Bh, Table 5], we select all entries satisfying these two conditions. We also know that $\operatorname{rank}(M_{Q}\bmod 2)\leq 2$ by Proposition 4.7, hence we compute the rank of each of them. The results are in Table 1. We see that we are left with $23$ possible candidates.
$L$ $\operatorname{rank}(M_{1\oplus L}\bmod{2})$
4: 1 1 4 0 0 0:41140004:\ 1\ 1\ 4\ 0\ 0\ 04 : 1 1 4 0 0 0 $3$
4: 1 2 2 0 0 0:41220004:\ 1\ 2\ 2\ 0\ 0\ 04 : 1 2 2 0 0 0 $2$
4: 2 2 2 2 2 0:42222204:\ 2\ 2\ 2\ 2\ 2\ 04 : 2 2 2 2 2 0 $3$
16: 1 2 8 0 0 0:1612800016:\ 1\ 2\ 8\ 0\ 0\ 016 : 1 2 8 0 0 0 $2$
16: 2 2 4 0 0 0:1622400016:\ 2\ 2\ 4\ 0\ 0\ 016 : 2 2 4 0 0 0 $1$
16: 2 3 3 2 0 0:1623320016:\ 2\ 3\ 3\ 2\ 0\ 016 : 2 3 3 2 0 0 $2$
20: 1 2 10 0 0 0:20121000020:\ 1\ 2\ 10\ 0\ 0\ 020 : 1 2 10 0 0 0 $2$
20: 2 2 5 0 0 0:$\ 2$\ 5\ 0\ 0\ 020 : 2 2 5 0 0 0 $2$
20: 2 2 6 2 2 0:$\ 2$\ 6\ 2\ 2\ 020 : 2 2 6 2 2 0 $3$
20: 2 3 4 0 0 2:2023400220:\ 2\ 3\ 4\ 0\ 0\ 220 : 2 3 4 0 0 2 $3$
20: 2 4 4 4 2 0:2024442020:\ 2\ 4\ 4\ 4\ 2\ 020 : 2 4 4 4 2 0 $3$
32: 2 4 4 0 0 0:3224400032:\ 2\ 4\ 4\ 0\ 0\ 032 : 2 4 4 0 0 0 $1$
32: 2 4 5 4 0 0:3224540032:\ 2\ 4\ 5\ 4\ 0\ 032 : 2 4 5 4 0 0 $2$
36: 2 3 6 0 0 0:3623600036:\ 2\ 3\ 6\ 0\ 0\ 036 : 2 3 6 0 0 0 $2$
36: 2 4 5 0 2 0:3624502036:\ 2\ 4\ 5\ 0\ 2\ 036 : 2 4 5 0 2 0 $3$
36: 2 4 6 4 2 0:3624642036:\ 2\ 4\ 6\ 4\ 2\ 036 : 2 4 6 4 2 0 $3$
36: 2 5 5 4 2 2:3625542236:\ 2\ 5\ 5\ 4\ 2\ 236 : 2 5 5 4 2 2 $3$
48: 2 3 8 0 0 0:4823800048:\ 2\ 3\ 8\ 0\ 0\ 048 : 2 3 8 0 0 0 $2$
48: 2 4 6 0 0 0:4824600048:\ 2\ 4\ 6\ 0\ 0\ 048 : 2 4 6 0 0 0 $1$
48: 2 5 5 2 0 0:4825520048:\ 2\ 5\ 5\ 2\ 0\ 048 : 2 5 5 2 0 0 $2$
52: 2 3 9 2 0 0:5223920052:\ 2\ 3\ 9\ 2\ 0\ 052 : 2 3 9 2 0 0 $2$
52: 2 5 6 2 0 2:5225620252:\ 2\ 5\ 6\ 2\ 0\ 252 : 2 5 6 2 0 2 $3$
52: 2 5 6 4 0 0:5225640052:\ 2\ 5\ 6\ 4\ 0\ 052 : 2 5 6 4 0 0 $2$
64: 2 4 8 0 0 0:6424800064:\ 2\ 4\ 8\ 0\ 0\ 064 : 2 4 8 0 0 0 $1$
68: 2 4 9 0 2 0:6824902068:\ 2\ 4\ 9\ 0\ 2\ 068 : 2 4 9 0 2 0 $3$
68: 2 4 10 4 2 0:68241042068:\ 2\ 4\ 10\ 4\ 2\ 068 : 2 4 10 4 2 0 $3$
68: 2 5 7 2 0 0:6825720068:\ 2\ 5\ 7\ 2\ 0\ 068 : 2 5 7 2 0 0 $2$
80: 2 4 10 0 0 0:80241000080:\ 2\ 4\ 10\ 0\ 0\ 080 : 2 4 10 0 0 0 $1$
80: 2 4 11 4 0 0:80241140080:\ 2\ 4\ 11\ 4\ 0\ 080 : 2 4 11 4 0 0 $2$
80: 2 5 8 0 0 0:8025800080:\ 2\ 5\ 8\ 0\ 0\ 080 : 2 5 8 0 0 0 $2$
96: 2 4 12 0 0 0:96241200096:\ 2\ 4\ 12\ 0\ 0\ 096 : 2 4 12 0 0 0 $1$
96: 2 4 13 4 0 0:96241340096:\ 2\ 4\ 13\ 4\ 0\ 096 : 2 4 13 4 0 0 $2$
100: 2 4 13 0 2 0:1002413020100:\ 2\ 4\ 13\ 0\ 2\ 0100 : 2 4 13 0 2 0 $3$
100: 2 4 14 4 2 0:1002414420100:\ 2\ 4\ 14\ 4\ 2\ 0100 : 2 4 14 4 2 0 $3$
100: 2 5 10 0 0 0:1002510000100:\ 2\ 5\ 10\ 0\ 0\ 0100 : 2 5 10 0 0 0 $2$
112: 2 4 14 0 0 0:1122414000112:\ 2\ 4\ 14\ 0\ 0\ 0112 : 2 4 14 0 0 0 $1$
Table 1. Ternary forms $L$ such that $1\oplus L$ is universal, $\det(M_{1\oplus L})$ is divisible by $4$, and $\frac{\det(M_{1\oplus L})}{4}\equiv 0,1\pmod{4}$.
To determine which of these candidates are equivalent to a quadratic form associated to some $G$, we use the $2$-adic normal form. Let $\operatorname{NF}(M)$ denote the normal form of a symmetric matrix $M$ over $\mathbb{Z}_{2}$. The precise definition is very technical and we do not give it here. It is sufficient for us to be able to compute $\operatorname{NF}(M)$, for which we use SageMath. Details on the implementation of the $p$-adic normal form can be found in the reference manual [Sa]. The important property of the normal form is uniqueness: Let $Q_{1}$ and $Q_{2}$ be two quadratic forms in $n$ variables over $\mathbb{Z}_{2}$ with matrices $M_{Q_{1}}$ and $M_{Q_{2}}$, respectively. Then $Q_{1}$ and $Q_{2}$ are equivalent over $\mathbb{Z}_{2}$ if and only if $\operatorname{NF}(M_{Q_{1}})=\operatorname{NF}(M_{Q_{2}})$.
Lemma 7.1.
Let $K=\mathbb{Q}(\sqrt{2})$ and let $G$ be an integral binary generalized form over $K$ such that $G(\mathbf{a})\in\mathbb{Q}$ for every $\mathbf{a}\in K^{2}$. Let $Q$ be the quaternary quadratic form associated to $G$ and $M_{Q}$ be the matrix of $Q$. We have $\operatorname{NF}(M_{Q})\bmod{8}=\operatorname{NF}(M)\bmod{8}$ for some matrix $M$ such that
(6) $M=\begin{pmatrix}A_{1}&4a_{2}&B_{1}&2E_{1}\\ 4a_{2}&2A_{2}&2E_{2}&2B_{2}\\ B_{1}&2E_{2}&C_{1}&4c_{2}\\ 2E_{1}&2B_{2}&4c_{2}&2C_{2}\end{pmatrix}$
where $A_{1},B_{1},C_{1}\in\{0,\dots,7\}$, $a_{2},c_{2}\in\{0,1\}$, $A_{2},C_{2},B_{2},E_{1},E_{2}\in\{0,\dots,3\}$, $A_{1}+A_{2}\equiv 0\pmod{4}$, $C_{1}+C_{2}\equiv 0\pmod{4}$, $B_{1}+B_{2}\equiv 0\pmod{2}$, and $E_{1}+E_{2}\equiv 0\pmod{2}$.
Proof.
By Lemma 5.3 with $D=2$, the matrix of $Q$ equals
$M_{Q}=\begin{pmatrix}2a_{1}+d&4a_{2}&b_{1}+e_{1}&2(b_{2}-e_{2})\\ 4a_{2}&2(2a_{1}-d)&2(b_{2}+e_{2})&2(b_{1}-e_{1})\\ b_{1}+e_{1}&2(b_{2}+e_{2})&2c_{1}+f&4c_{2}\\ 2(b_{2}-e_{2})&2(b_{1}-e_{1})&4c_{2}&2(2c_{1}-f)\end{pmatrix}.$
Let $A_{1}=2a_{1}+d$, $A_{2}=2a_{1}-d$, $C_{1}=2c_{1}+f$, $C_{2}=2c_{1}-f$, $B_{1}=b_{1}+e_{1}$, $B_{2}=b_{1}-e_{1}$, $E_{1}=b_{2}-e_{2}$, and $E_{2}=b_{2}+e_{2}$. These parameters clearly satisfy the stated congruences. Since $\operatorname{NF}(M_{Q})\bmod{8}$ depends only on $M_{Q}\bmod{8}$, we can reduce $M_{Q}$ modulo $8$ to obtain one of the matrices (6). ∎
We compute $\operatorname{NF}(M_{1\oplus L})\bmod{8}$ for each of the forms in Table 1 satisfying $\operatorname{rank}(1\oplus L\bmod{2})\leq 2$ and check whether it equals $\operatorname{NF}(M)\bmod{8}$ for one of the finitely many matrices $M$ in Lemma 7.1. If not, then we know that $1\oplus L$ is not equivalent to a quadratic form associated to some $G$ and we can discard it. The SageMath code of this computation is available at: https://github.com/zinmik/Generalized-forms.
We are left with $15$ quadratic forms $1\oplus L$ and for each one, we directly find $G$ such that the quadratic form associated to $G$ is equivalent to $1\oplus L$, see Table 2.
$L$ $G$ $Q$
4: 1 2 2 0 0 0:41220004:\ 1\ 2\ 2\ 0\ 0\ 04 : 1 2 2 0 0 0 $[1,-\sqrt{2},1,-1,0,-1]$ $[1,0,0,-4,6,-4,0,1,0,6]$
16: 2 2 4 0 0 0:1622400016:\ 2\ 2\ 4\ 0\ 0\ 016 : 2 2 4 0 0 0 $[1,-1,1,-1,1,0]$ $[1,0,0,0,6,0,-8,2,0,4]$
20: 1 2 10 0 0 0:20121000020:\ 1\ 2\ 10\ 0\ 0\ 020 : 1 2 10 0 0 0 $[1,-\sqrt{2},2,-1,-\sqrt{2},-1]$ $[1,0,0,0,6,-8,0,3,0,10]$
20: 2 2 5 0 0 0:$\ 2$\ 5\ 0\ 0\ 020 : 2 2 5 0 0 0 $[1,-\sqrt{2},2,1,\sqrt{2},1]$ $[3,0,0,-8,2,0,0,5,0,6]$
32: 2 4 4 0 0 0:3224400032:\ 2\ 4\ 4\ 0\ 0\ 032 : 2 4 4 0 0 0 $[1,-1-\sqrt{2},2,-1,1,-2]$ $[1,0,0,-4,6,-4,-8,2,0,12]$
36: 2 3 6 0 0 0:3623600036:\ 2\ 3\ 6\ 0\ 0\ 036 : 2 3 6 0 0 0 $[1,0,1,-1,0,1]$ $[1,0,0,0,6,0,0,3,0,2]$
48: 2 4 6 0 0 0:4824600048:\ 2\ 4\ 6\ 0\ 0\ 048 : 2 4 6 0 0 0 $[1,0,1,-1,0,0]$ $[1,0,0,0,6,0,0,2,0,4]$
52: 2 3 9 2 0 0:5223920052:\ 2\ 3\ 9\ 2\ 0\ 052 : 2 3 9 2 0 0 $[1,-\sqrt{2},2,1,0,-1]$ $[3,0,0,-4,2,-4,0,3,0,10]$
52: 2 5 6 4 0 0:5225640052:\ 2\ 5\ 6\ 4\ 0\ 052 : 2 5 6 4 0 0 $[1,-\sqrt{2},2,-1,0,1]$ $[1,0,0,-4,6,-4,0,5,0,6]$
64: 2 4 8 0 0 0:6424800064:\ 2\ 4\ 8\ 0\ 0\ 064 : 2 4 8 0 0 0 $[1,-\sqrt{2},2,-1,0,-2]$ $[1,0,0,-4,6,-4,0,2,0,12]$
68: 2 5 7 2 0 0:6825720068:\ 2\ 5\ 7\ 2\ 0\ 068 : 2 5 7 2 0 0 $[2,-2\sqrt{2},2,-1,0,1]$ $[3,0,0,-8,10,-8,0,5,0,6]$
80: 2 4 10 0 0 0:80241000080:\ 2\ 4\ 10\ 0\ 0\ 080 : 2 4 10 0 0 0 $[1,-1,2,0,-1,-1]$ $[2,0,-4,0,4,0,0,3,0,10]$
96: 2 4 12 0 0 0:96241200096:\ 2\ 4\ 12\ 0\ 0\ 096 : 2 4 12 0 0 0 $[1,-1+\sqrt{2},3-\sqrt{2},-1,1+\sqrt{2},-2]$ $[1,0,0,0,6,8,-8,4,-8,16]$
100: 2 5 10 0 0 0:1002510000100:\ 2\ 5\ 10\ 0\ 0\ 0100 : 2 5 10 0 0 0 $[2,-\sqrt{2},2,-1,\sqrt{2},1]$ $[3,0,0,-8,10,0,0,5,0,6]$
112: 2 4 14 0 0 0:1122414000112:\ 2\ 4\ 14\ 0\ 0\ 0112 : 2 4 14 0 0 0 $[2,0,1,-3,0,0]$ $[1,0,0,0,14,0,0,2,0,4]$
Table 2. Quadratic forms associated to $\mathbb{Z}$-universal binary generalized forms over $\mathbb{Q}(\sqrt{2})$.
We proved the following proposition.
Proposition 7.2.
Let $K=\mathbb{Q}(\sqrt{2})$ and let $G$ be a $\mathbb{Z}$-universal binary generalized form over $K$. If $Q$ is the quadratic form associated to $G$, then $Q$ is equivalent to $1\oplus L$ where $L$ is one of the $15$ ternary forms in Table 2. Conversely, for each of these forms, there exists an equivalent quadratic form associated to a $\mathbb{Z}$-universal binary form over $K$.
7.3. Universal quadratic forms associated to binary generalized forms for $D=3$
Lemma 7.3.
Let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}\setminus\{0,1\}$ is squarefree. Let $G$ be an integral binary generalized form over $K$ and $Q$ be the quadratic form associated to $G$ with matrix $M_{Q}$. Assume that $G(\mathbf{a})\in\mathbb{Q}$ for every $\mathbf{a}\in K^{2}$. If $D\equiv 3\pmod{4}$, then $\operatorname{rank}(M_{Q}\bmod{2})\in\{0,2,4\}$.
Proof.
Let the matrix of $G$ be given by (5). By Lemma 5.3,
$M_{Q}=\begin{pmatrix}2a_{1}+d&2Da_{2}&b_{1}+e_{1}&D(b_{2}-e_{2})\\ 2Da_{2}&D(2a_{1}-d)&D(b_{2}+e_{2})&D(b_{1}-e_{1})\\ b_{1}+e_{1}&D(b_{2}+e_{2})&2c_{1}+f&2Dc_{2}\\ D(b_{2}-e_{2})&D(b_{1}-e_{1})&2Dc_{2}&D(2c_{1}-f)\end{pmatrix}.$
Let $B=b_{1}+e_{1}$ and $E=b_{2}+e_{2}$. Since $D\equiv 1\pmod{2}$, we get
$M_{Q}\bmod{2}=\begin{pmatrix}d&0&B&E\\ 0&d&E&B\\ B&E&f&0\\ E&B&0&f\end{pmatrix}.$
There are $16$ choices for $(d,B,E,f)$ modulo $2$ and it can be checked that $\operatorname{rank}(M_{Q}\bmod{2})\in\{0,2,4\}$ for each of them. ∎
Let $Q$ be a classical quaternary positive definite form over $\mathbb{Z}$ with matrix $M_{Q}$ and suppose that there exists a binary integral generalized form $G$ over $K=\mathbb{Q}(\sqrt{3})$ such that $Q$ is associated to $G$. We know from Corollary 4.3 that $3^{2}\mid\det(M_{Q})$. Moreover, by Proposition 5.14,
$\frac{\det(M_{Q})}{3^{2}}\equiv 0,1\pmod{4}.$
By Proposition 4.7, $\operatorname{rank}(M_{Q}\bmod{3})\leq 2$ and by Lemma 7.3, $\operatorname{rank}(M_{Q}\bmod{2})$ is even.
In Table 3, we list all universal classical quaternary quadratic forms $1\oplus L$ with $\det(M_{1\oplus L})$ divisible by $9$ and such that $\frac{\det(M_{1\oplus L})}{9}\equiv 0,1\pmod{4}$ from [Bh, Table 5]. For each of them we compute its rank modulo $2$ and $3$.
$L$ $\operatorname{rank}(M_{1\oplus L}\bmod{2})$ $\operatorname{rank}(M_{1\oplus L}\bmod{3})$
9: 1 2 5 2 0 0:91252009:\ 1\ 2\ 5\ 2\ 0\ 09 : 1 2 5 2 0 0 $4$ $3$
9: 1 3 3 0 0 0:91330009:\ 1\ 3\ 3\ 0\ 0\ 09 : 1 3 3 0 0 0 $4$ $2$
9: 2 2 3 0 0 $\ 2$\ 3\ 0\ 0\ 29 : 2 2 3 0 0 2 $4$ $2$
36: 2 3 6 0 0 0:3623600036:\ 2\ 3\ 6\ 0\ 0\ 036 : 2 3 6 0 0 0 $2$ $2$
36: 2 4 5 0 2 0:3624502036:\ 2\ 4\ 5\ 0\ 2\ 036 : 2 4 5 0 2 0 $3$ $3$
36: 2 4 6 4 2 0:3624642036:\ 2\ 4\ 6\ 4\ 2\ 036 : 2 4 6 4 2 0 $3$ $3$
36: 2 5 5 4 2 2:3625542236:\ 2\ 5\ 5\ 4\ 2\ 236 : 2 5 5 4 2 2 $3$ $2$
45: 2 4 7 0 2 2:4524702245:\ 2\ 4\ 7\ 0\ 2\ 245 : 2 4 7 0 2 2 $4$ $3$
45: 2 5 5 0 2 0:4525502045:\ 2\ 5\ 5\ 0\ 2\ 045 : 2 5 5 0 2 0 $4$ $3$
45: 2 5 6 4 2 2:4525642245:\ 2\ 5\ 6\ 4\ 2\ 245 : 2 5 6 4 2 2 $4$ $3$
72: 2 4 9 0 0 0:7224900072:\ 2\ 4\ 9\ 0\ 0\ 072 : 2 4 9 0 0 0 $2$ $3$
72: 2 4 10 4 0 0:72241040072:\ 2\ 4\ 10\ 4\ 0\ 072 : 2 4 10 4 0 0 $1$ $3$
72: 2 5 8 4 0 0:7225840072:\ 2\ 5\ 8\ 4\ 0\ 072 : 2 5 8 4 0 0 $2$ $3$
90: 2 4 12 2 2 0:90241222090:\ 2\ 4\ 12\ 2\ 2\ 090 : 2 4 12 2 2 0 $3$ $3$
90: 2 5 9 0 0 0:9025900090:\ 2\ 5\ 9\ 0\ 0\ 090 : 2 5 9 0 0 0 $3$ $3$
108: 2 4 14 0 2 0:1082414020108:\ 2\ 4\ 14\ 0\ 2\ 0108 : 2 4 14 0 2 0 $3$ $3$
Table 3. Ternary forms $L$ such that $1\oplus L$ is universal, $\det(M_{1\oplus L})$ is divisible by $9$ and $\frac{\det(M_{1\oplus L})}{9}\equiv 0,1\pmod{4}$.
Proposition 7.4.
Let $K=\mathbb{Q}(\sqrt{3})$. If $G$ is a $\mathbb{Z}$-universal binary generalized form over $K$ and $Q$ is the quadratic form associated to $G$, then $Q$ is equivalent to one of the following forms:
$\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&3&0\\ 0&0&0&3\end{pmatrix},\ \begin{pmatrix}1&0&0&0\\ 0&2&1&0\\ 0&1&2&0\\ 0&0&0&3\end{pmatrix},\ \begin{pmatrix}1&0&0&0\\ 0&2&0&0\\ 0&0&3&0\\ 0&0&0&6\end{pmatrix}.$
Conversely, for each of these forms, there exists an equivalent form $Q$ associated to a $\mathbb{Z}$-universal binary generalized form over $K$.
Proof.
Let $M_{Q}$ be the matrix of $Q$. If $Q$ is associated to a $\mathbb{Z}$-universal binary generalized form, then $Q$ must be equivalent to one of the forms in Table 3 by the discussion before the proposition. Moreover, $\operatorname{rank}(M_{Q}\bmod{3})\leq 2$ and $\operatorname{rank}(M_{Q}\bmod{2})$ is even. The only forms $L$ in Table 3 such that $\operatorname{rank}(M_{1\oplus L}\bmod{2})$ is even and $\operatorname{rank}(M_{1\oplus L}\bmod{3})\leq 2$ are 9: 1 3 3 0 0 0:91330009:\ 1\ 3\ 3\ 0\ 0\ 09 : 1 3 3 0 0 0, 9: 2 2 3 0 0 $\ 2$\ 3\ 0\ 0\ 29 : 2 2 3 0 0 2, and 36: 2 3 6 0 0 0:3623600036:\ 2\ 3\ 6\ 0\ 0\ 036 : 2 3 6 0 0 0.
Now we prove the converse. Let $G_{1}$ be the generalized form $[2,-2\sqrt{3},2,-3,-\sqrt{3},0]$ and let $Q_{1}$ be the quadratic form associated to $G_{1}$. By Lemma 5.3, the matrix of $Q_{1}$ equals
$M_{Q_{1}}=\begin{pmatrix}1&0&0&-3\\ 0&21&-9&0\\ 0&-9&4&0\\ -3&0&0&12\end{pmatrix}.$
We verified in Lemma 5.5 that $Q_{1}$ is equivalent to 9: 1 3 3 0 0 0:91330009:\ 1\ 3\ 3\ 0\ 0\ 09 : 1 3 3 0 0 0.
Let $G_{2}$ be the generalized form $[1,-1-\sqrt{3},2,-1,1,-2]$ and let $Q_{2}$ be the quadratic form associated to $G_{2}$. The matrix of $Q_{2}$ equals
$M_{Q_{2}}=\begin{pmatrix}1&0&0&-3\\ 0&9&-3&-6\\ 0&-3&2&0\\ -3&-6&0&18\end{pmatrix}.$
If we let
$A=\begin{pmatrix}1&-3&0&3\\ 0&-1&0&2\\ 0&-1&1&3\\ 0&-1&0&1\end{pmatrix},$
then $\det(A)=1$, hence $A\in\operatorname{SL}_{4}(\mathbb{Z})$. We have
$A^{\top}M_{Q_{2}}A=\begin{pmatrix}1&-3&0&3\\ 0&-1&0&2\\ 0&-1&1&3\\ 0&-1&0&1\end{pmatrix}^{\top}\begin{pmatrix}1&0&0&-3\\ 0&9&-3&-6\\ 0&-3&2&0\\ -3&-6&0&18\end{pmatrix}\begin{pmatrix}1&-3&0&3\\ 0&-1&0&2\\ 0&-1&1&3\\ 0&-1&0&1\end{pmatrix}=\begin{pmatrix}1&0&0&0\\ 0&2&1&0\\ 0&1&2&0\\ 0&0&0&3\end{pmatrix},$
hence $Q_{2}$ is equivalent to 9: 2 2 3 0 0 $\ 2$\ 3\ 0\ 0\ 29 : 2 2 3 0 0 2.
Let $G_{3}$ be the generalized form $[1,1,1,0,1,1]$ and let $Q_{3}$ be the quadratic form associated to $G_{3}$. The matrix of $Q_{3}$ equals
$M_{Q_{3}}=\begin{pmatrix}2&0&2&0\\ 0&6&0&0\\ 2&0&3&0\\ 0&0&0&3\end{pmatrix}\sim\begin{pmatrix}1&0&0&0\\ 0&2&0&0\\ 0&0&3&0\\ 0&0&0&6\end{pmatrix},$
hence $Q_{3}$ is equivalent to 36: 2 3 6 0 0 0:3623600036:\ 2\ 3\ 6\ 0\ 0\ 036 : 2 3 6 0 0 0. ∎
7.4. Universal quadratic forms associated to classical binary generalized forms for $D=5$
Let $Q$ be a classical quaternary positive definite form over $\mathbb{Z}$ with matrix $M_{Q}$ and suppose that there exists a binary classical generalized form $G$ over $K=\mathbb{Q}(\sqrt{5})$ such that $Q$ is associated to $G$. By Corollary 4.3, $5^{2}\mid\det(M_{Q})$ and by Proposition 4.7, $\operatorname{rank}(M_{Q}\bmod{5})\leq 2$.
In Table 4, we list all universal quaternary quadratic forms $1\oplus L$ with $\det(M_{1\oplus L})$ divisible by $25$ from [Bh, Table 5]. For each of the we compute its rank modulo $5$.
$L$ $\operatorname{rank}(M_{1\oplus L}\bmod{5})$
25: 1 2 13 2 0 0:25121320025:\ 1\ 2\ 13\ 2\ 0\ 025 : 1 2 13 2 0 0 $3$
25: 2 3 5 0 0 2:2523500225:\ 2\ 3\ 5\ 0\ 0\ 225 : 2 3 5 0 0 2 $2$
25: 2 3 5 2 2 0:2523522025:\ 2\ 3\ 5\ 2\ 2\ 025 : 2 3 5 2 2 0 $3$
50: 2 4 7 2 2 0:5024722050:\ 2\ 4\ 7\ 2\ 2\ 050 : 2 4 7 2 2 0 $3$
100: 2 4 13 0 2 0:1002413020100:\ 2\ 4\ 13\ 0\ 2\ 0100 : 2 4 13 0 2 0 $3$
100: 2 4 14 4 2 0:1002414420100:\ 2\ 4\ 14\ 4\ 2\ 0100 : 2 4 14 4 2 0 $3$
100: 2 5 10 0 0 0:1002510000100:\ 2\ 5\ 10\ 0\ 0\ 0100 : 2 5 10 0 0 0 $2$
Table 4. Ternary forms $L$ such that $1\oplus L$ is universal and $\det(M_{1\oplus L})$ is divisible by $25$.
Proposition 7.5.
Let $K=\mathbb{Q}(\sqrt{5})$. If $G$ is a classical $\mathbb{Z}$-universal binary generalized form over $K$ and $Q$ is a quadratic form associated to $G$, then $Q$ is equivalent to one of the following forms:
$\begin{pmatrix}1&0&0&0\\ 0&2&1&0\\ 0&1&3&0\\ 0&0&0&5\end{pmatrix},\ \begin{pmatrix}1&0&0&0\\ 0&2&0&0\\ 0&0&5&0\\ 0&0&0&10\end{pmatrix}.$
Conversely, for each of these forms, there exists an equivalent form $Q$ associated to a classical $\mathbb{Z}$-universal binary generalized form over $K$.
Proof.
Let $M_{Q}$ be the matrix of $Q$. If $Q$ is associated to a classical $\mathbb{Z}$-universal binary form, then $Q$ must be one of the forms listed in Table 4. Moreover, $\operatorname{rank}(M_{Q}\bmod{5})\leq 2$. The only forms $L$ in Table 4 such that $\operatorname{rank}(M_{1\oplus L}\bmod{5})\leq 2$ are 25: 2 3 5 0 0 2:2523500225:\ 2\ 3\ 5\ 0\ 0\ 225 : 2 3 5 0 0 2 and 100: 2 5 10 0 0 0:1002510000100:\ 2\ 5\ 10\ 0\ 0\ 0100 : 2 5 10 0 0 0.
Let $G_{1}=[3-\omega_{5},0,1,-4,0,0]$ and $G_{2}=[2+\omega_{5},0,6-2\omega_{5},-4,0,-8]$ where $\omega_{5}=\frac{1+\sqrt{5}}{2}$ and let $Q_{1}$, $Q_{2}$ be the quadratic forms associated to $G_{1}$ and $G_{2}$, respectively. By Lemma 5.3, the matrices of $Q_{1}$ and $Q_{2}$ are
$M_{Q_{1}}$ $=\begin{pmatrix}1&-2&0&0\\ -2&9&0&0\\ 0&0&2&1\\ 0&0&1&3\end{pmatrix}\sim\begin{pmatrix}1&0&0&0\\ 0&2&1&0\\ 0&1&3&0\\ 0&0&0&5\end{pmatrix},$
$M_{Q_{2}}$ $=\begin{pmatrix}1&3&0&0\\ 3&14&0&0\\ 0&0&2&-4\\ 0&0&-4&18\end{pmatrix}\sim\begin{pmatrix}1&0&0&0\\ 0&2&0&0\\ 0&0&5&0\\ 0&0&0&10\end{pmatrix}.$
(We showed this for $G_{1}$ in Lemma 5.6.) ∎
7.5. Universal quadratic forms associated to binary generalized forms for $D=6,7,10$
Proposition 7.6.
Let $K=\mathbb{Q}(\sqrt{6})$. If $G$ is a $\mathbb{Z}$-universal binary generalized form over $K$ and $Q$ is the quadratic form associated to $G$, then $Q$ is equivalent to the form
$\begin{pmatrix}1&0&0&0\\ 0&2&0&0\\ 0&0&3&0\\ 0&0&0&6\end{pmatrix}.$
Conversely, there exists an equivalent form $Q$ associated to a $\mathbb{Z}$-universal binary generalized form over $K$.
Proof.
Let $M_{Q}$ be the matrix of $Q$. By Corollary 4.3, $6^{2}\mid\det(M_{Q})$ and by Proposition 5.14,
$\frac{\det(M_{Q})}{6^{2}}\equiv 0,1\pmod{4}.$
Moreover, by Proposition 4.7, $\operatorname{rank}(M_{Q}\bmod{2})\leq 2$ and $\operatorname{rank}(M_{Q}\bmod{3})\leq 2$.
All universal classical quaternary quadratic forms with determinant divisible by $36$ were already listed in Table 3. The only form satisfying the additional conditions is 36: 2 3 6 0 0 0:3623600036:\ 2\ 3\ 6\ 0\ 0\ 036 : 2 3 6 0 0 0.
If we let $G=[3,2\sqrt{6},3,-5,\sqrt{6},-1]$, then
$M_{Q}=\begin{pmatrix}1&0&0&6\\ 0&66&18&0\\ 0&18&5&0\\ 6&0&0&42\end{pmatrix}\sim\begin{pmatrix}1&0&0&0\\ 0&2&0&0\\ 0&0&3&0\\ 0&0&0&6\end{pmatrix}.$
(We showed this in Lemma 5.7.) ∎
Proposition 7.7.
Let $K=\mathbb{Q}(\sqrt{7})$. If $G$ is a $\mathbb{Z}$-universal binary generalized form over $K$ and $Q$ is the quadratic form associated to $G$, then $Q$ is equivalent to the form
$\begin{pmatrix}1&0&0&0\\ 0&2&1&0\\ 0&1&4&0\\ 0&0&0&7\end{pmatrix}.$
Conversely, there exists an equivalent form $Q$ associated to a $\mathbb{Z}$-universal binary generalized form over $K$.
Proof.
Let $M_{Q}$ be the matrix of $Q$. By Corollary 4.3, $7^{2}\mid\det(M_{Q})$ and by Proposition 5.14,
$\frac{\det(M_{Q})}{7^{2}}\equiv 0,1\pmod{4}.$
The only forms in [Bh, Table 5] satisfying these conditions are those with determinant equal to $49$, namely
$L_{1}$ =49: 2 3 9 2 2 0,:absent49239220=49:\ 2\ 3\ 9\ 2\ 2\ 0,= 49 : 2 3 9 2 2 0 ,
$L_{2}$ =49: 2 4 7 0 0 2,:absent49247002=49:\ 2\ 4\ 7\ 0\ 0\ 2,= 49 : 2 4 7 0 0 2 ,
$L_{3}$ =49: 2 5 6 0 2 2.:absent49256022=49:\ 2\ 5\ 6\ 0\ 2\ 2.= 49 : 2 5 6 0 2 2 .
By Proposition 4.7, $\operatorname{rank}(M_{Q}\bmod{7})\leq 2$. We have $\operatorname{rank}(1\oplus L_{1}\bmod{7})=3$, $\operatorname{rank}(1\oplus L_{2}\bmod{7})=2$, and $\operatorname{rank}(1\oplus L_{3}\bmod{7})=3$. Thus, the only candidate is 49: 2 4 7 0 0 2:4924700249:\ 2\ 4\ 7\ 0\ 0\ 249 : 2 4 7 0 0 2.
Let $G=[3,1+2\sqrt{7},4,-5,-1+2\sqrt{7},3]$. The matrix $M_{Q}$ of the quadratic form $Q$ associated to $G$ equals
$M_{Q}=\begin{pmatrix}1&0&0&0\\ 0&77&28&14\\ 0&28&11&0\\ 0&14&0&35\end{pmatrix}\sim\begin{pmatrix}1&0&0&0\\ 0&2&1&0\\ 0&1&4&0\\ 0&0&0&7\end{pmatrix}.$
(We showed this in Lemma 5.8.) ∎
Proposition 7.8.
Let $K=\mathbb{Q}(\sqrt{10})$. If $G$ is a $\mathbb{Z}$-universal binary generalized form over $K$ and $Q$ is the quadratic form associated to $G$, then $Q$ is equivalent to the form
$\begin{pmatrix}1&0&0&0\\ 0&2&0&0\\ 0&0&5&0\\ 0&0&0&10\end{pmatrix}.$
Conversely, there exists an equivalent form $Q$ associated to a $\mathbb{Z}$-universal binary generalized form over $K$.
Proof.
Let $M_{Q}$ be the matrix of $Q$. By Corollary 4.3, $10^{2}\mid\det(M_{Q})$. The quadratic forms from [Bh, Table 5] with determinant divisible by $100$ were already listed in Table 4. By Proposition 4.7, $\operatorname{rank}(M_{Q}\bmod{5})\leq 2$. The only form satisfying these criteria is 100: 2 5 10 0 0 0:1002510000100:\ 2\ 5\ 10\ 0\ 0\ 0100 : 2 5 10 0 0 0.
If we let $G=[3,\sqrt{10},5,-5,\sqrt{10},7]$, then
$M_{Q}=\begin{pmatrix}1&0&0&0\\ 0&110&20&0\\ 0&20&17&20\\ 0&0&20&30\end{pmatrix}\sim\begin{pmatrix}1&0&0&0\\ 0&2&0&0\\ 0&0&5&0\\ 0&0&0&10\end{pmatrix}.$
(We showed this in Lemma 5.9.) ∎

### 21 Mar 2026 v2

@@
 11E12, 11E16, 11E20, 11E25, 11E39, 11R11, 11R80
-We acknowledge support by Czech Science Foundation (GAČR) grant 21-00420M and the Charles University Research Centre program UNCE/24/SCI/022.
+We acknowledge support by Czech Science Foundation (GAČR) grant 26-20514S, and Charles University programmes PRIMUS/24/SCI/010 and UNCE/24/SCI/022, and GAUK projects 134824 and 236225.
 1. Introduction
@@
 Diagonal quaternary universal forms were classified by Ramanujan [Ra], who stated that there exist $55$ such forms.
-Dickson [Di] corrected Ramanujan’s statement by pointing out that the form $x^{2}+2y^{2}+5z^{2}+5u^{2}$ does not represent $15$.
+Dickson [Di1] corrected Ramanujan’s statement by pointing out that the form $x^{2}+2y^{2}+5z^{2}+5u^{2}$ does not represent $15$.
 Thus there exist exactly $54$ diagonal universal forms.
 Dickson [Di2] also proved various results on representation by non-diagonal quaternary forms.
-The 90’s saw a proof of the celebrated Conway-Schneeberger 15-Theorem. A quadratic form having an integer matrix is called classical.
+In 1993, John H. Conway and W. A. Schneeberger proved the celebrated 15-Theorem. A quadratic form having an integer matrix is called classical.
 Theorem 1.1 (15-Theorem, Conway-Schneeberger).
 A classical quadratic form $Q$ over $\mathbb{Z}$ is universal if and only if it represents the integers
-[…]
+$1,2,3,5,6,7,10,14,\text{and }15.$
 The proof was simplified by Bhargava [Bh].
-As a corollary, one obtains all classical quaternary forms – there are exactly $207$ of them.
+As a corollary, one obtains all universal classical quaternary forms – there are exactly $207$ of them.
 Bhargava and Hanke [BH] extended the result to non-classical forms by proving the 290-Theorem: If a positive definite quadratic form $Q$ over $\mathbb{Z}$ represents every positive integer $n\leq 290$, then $Q$ is universal.
@@
 By a result of Siegel [Si], if a sum of squares is universal over $K$, then $K=\mathbb{Q}$ or $\mathbb{Q}(\sqrt{5})$.
-In 1993, Chan, Kim, and Raghavan [CKR] proved that a ternary universal quadratic form exists over a real quadratic field $K=\mathbb{Q}(\sqrt{D})$ if and only if $D$ equals $2$, $3$, or $5$.
+In 1993, Chan, Kim, and Raghavan [CKR] proved that a ternary universal quadratic form exists over a real quadratic field $K=\mathbb{Q}(\sqrt{D})$ if and only if $D$ equals $2$, 33, or $5$.
 Moreover, they identified all ternary universal forms over these fields up to equivalence.
@@
 M.
-Kim [Ki] proved that there exist only finitely many $D$’s such that $\mathbb{Q}(\sqrt{D})$ admits a universal form in $7$ variables (see also [KKP2] for an explicit estimate on the size of $D$).
-On the other hand, a $8$-ary universal form exists over each $\mathbb{Q}(\sqrt{n^{2}-1})$ (when $n^{2}-1$ is squarefree) [Ki2].
-Blomer and Kala [BK] showed: For any positive integer $r$, there are infinitely many quadratic fields $\mathbb{Q}(\sqrt{D})$ that do not have a universal lattice of rank $\leq r$.
+Kim [Ki1] proved that there exist only finitely many $D$’s such that $\mathbb{Q}(\sqrt{D})$ admits a universal form in $7$ variables (see also [KKP2] for an explicit estimate on the size of $D$).
+On the other hand, a $8$-ary universal form exists over each $\mathbb{Q}(\sqrt{n^{2}-1})$ whenever $n^{2}-1$ is squarefree [Ki2].
+Blomer and Kala [BK1] showed: For any positive integer $r$, there are infinitely many quadratic fields $\mathbb{Q}(\sqrt{D})$ that do not have a universal lattice of rank $\leq r$.
 Lower bounds for the rank of a universal form depending on $D$ were obtained by Kala and Tinková [KT].
@@
 We call a positive definite Hermitian form over an imaginary quadratic field $\mathbb{Z}$-universal if it represents every $n\in\mathbb{Z}_{\geq 1}$.
-Earnest and Khosravani [EK] proved that universal binary Hermitian forms exist over $K$ for only finitely many $m$, and identified all such forms in the case when $K$ has class number $1$.
+Earnest and Khosravani [EK] proved that universal binary Hermitian forms exist over $K=\mathbb{Q}(\sqrt{m})$ for only finitely many $m$, and identified all such forms in the case when $K$ has class number $1$.
 B.
@@
 Kim, and P.-S.
-Park [KKP] proved an analogy of the 15-Theorem for Hermitian quadratic forms: If a positive definite Hermitian form $H$ represents $1$, $2$, $3$, $5$, $6$, $7$, $10$, $13$, $14$, and $15$, then $H$ is universal.
+Park [KKP1] proved an analogy of the 15-Theorem for Hermitian quadratic forms: If a positive definite Hermitian form $H$ represents $1$, 22, 33, 55, 66, 77, $10$, $13$, $14$, and $15$, then $H$ is universal.
 They also completely determined the minimal rank of a positive Hermitian form over all imaginary quadratic fields.
-The object of study of our paper are generalized quadratic forms introduced recently by Browning, Pierce, and Schindler [BPS, Definition 1.1].
+Our paper concerns generalized quadratic forms introduced recently by Browning, Pierce, and Schindler [BPS, Definition 1.1].
 They can be defined for an arbitrary Galois extension $K$ of $\mathbb{Q}$ but we will consider them only over quadratic fields.
 A quadratic field $K=\mathbb{Q}(\sqrt{D})$ has two $\mathbb{Q}$-automorphisms: the identity and the conjugation map
-$\tau:K\hookrightarrow K,a+b\sqrt{D}\mapsto a-b\sqrt{D}.$
+$\tau:K$ $\hookrightarrow K$
+$a+b\sqrt{D}$ $\mapsto a-b\sqrt{D}.$
 Informally speaking, a generalized quadratic form may contain also the conjugates of the variables. For example,
@@
 If $G(z_{1},z_{2},\dots,z_{n})$ is a generalized form in $n$ variables and
-$G(\alpha_{1},\alpha_{2},\dots,\alpha_{n})\in\mathbb{Z},\quad\alpha_{1},\alpha_{2},\dots,\alpha_{n}\in\mathcal{O}_{K},$
+$G(\alpha_{1},\alpha_{2},\dots,\alpha_{n})\in\mathbb{Z},\quad\forall\alpha_{1},\alpha_{2},\dots,\alpha_{n}\in\mathcal{O}_{K},$
 then we say that $G$ is $\mathbb{Z}$-valued.
@@
 There are only finitely many real quadratic fields which admit a binary or ternary $\mathbb{Z}$-universal generalized quadratic form.
-We will prove these as Theorems 5.10, 5.11 and 5.12 in Section 5.
+We will prove these as Theorems 4.10, 4.11 and 4.12 in Section 4.
 In contrast with the situation for binary and ternary forms, it is not very difficult to construct a generalized form in four variables which is $\mathbb{Z}$-universal over every real quadratic field (and we do this later in Proposition 4.13).
@@
 The rest of the paper is organized as follows.
-In Section 2, we collect the preliminaries about quadratic forms and quadratic fields.
 In Section 2, we find a minimal set of generators of the ring of $\mathbb{Q}$-valued generalized polynomials.
@@
 We prove Theorems 1.2, 1.3 and 1.4 at the end of Section 4 and Theorem 5.1 in Section 5.
-We find all quaternary forms associated to $\mathbb{Z}$-universal binary generalized forms for $D\in\{2,3,6,7,10\}$ (and to classical ones for $D=5$) in Section 7.
+In the appendix, we give the proofs of Lemmata 4.5-4.9 in Section A4 and a more detailed approach to the proof of Theorem 5.1 in Section A5.
+In Section A6 we explicitly show all quaternary forms associated to $\mathbb{Z}$-universal binary generalized forms for $D\in\{2,3,6,7,10\}$ (and to classical ones for $D=5$).
 Acknowledgments
 This research project was accomplished at the Student Number Theory Seminar at Charles University.
-We thank Vítězslav Kala for organizing the seminar.
-We also thank Nicolas Daans for suggesting the proof of Proposition 6.7 and Pavlo Yatsyna for his recommendation of the reference [Ca].
-2. Preliminaries
-Let $K$ be a number field and let $\mathcal{O}_{K}$ denote the ring of integers of $K$. Consider a polynomial $p$ in $n$ variables over $K$, i.e.,
-$p(x_{1},x_{2},\dots,x_{n})=\sum_{\begin{subarray}{c}i_{1},\dots,i_{n}\geq 0\\ i_{1}+\cdots+i_{n}\leq k\end{subarray}}a_{i_{1},i_{2},\dots,i_{n}}x_{1}^{i_{1}}x_{2}^{i_{2}}\cdots x_{n}^{i_{n}}$
-where $k\in\mathbb{Z}_{\geq 0}$ and $a_{i_{1},i_{2},\dots,i_{n}}\in K$ for $i_{1},i_{2},\dots,i_{n}\in\{0,1,\dots,k\}$.
-We will make use of the multi-index notation.
-If $\mathbf{i}=(i_{1},i_{2},\dots,i_{n})\in\mathbb{Z}_{\geq 0}^{n}$ is a multi-index, then we let $|\mathbf{i}|=\sum_{r=1}^{n}i_{r}$ be the sum of its components, and if $\mathbf{x}=(x_{1},x_{2},\dots,x_{n})$, then we let $\mathbf{x}^{\mathbf{i}}=x_{1}^{i_{1}}x_{2}^{i_{2}}\cdots x_{n}^{i_{n}}$.
-The polynomial $p$ can be succinctly written as
-$p(\mathbf{x})=\sum_{|\mathbf{i}|\leq k}a_{\mathbf{i}}\mathbf{x}^{\mathbf{i}}.$
-The lexicographic order on $\mathbb{Z}_{\geq 0}^{n}$ is defined as follows: Let $\mathbf{i},\mathbf{j}\in\mathbb{Z}_{\geq 0}^{n}$, $\mathbf{i}=(i_{1},i_{2},\dots,i_{n})$ and $\mathbf{j}=(j_{1},j_{2},\dots,j_{n})$.
-Then $\mathbf{i}<\mathbf{j}$ if and only if there exists $s\in\{1,2,\dots,n\}$ such that $i_{r}=j_{r}$ for $1\leq r<s$ and $i_{s}<j_{s}$.
-A quadratic form $Q$ in $n$ variables over $K$ is a homogeneous polynomial of degree $2$, i.e.,
-$Q(x_{1},x_{2},\dots,x_{n})=\sum_{1\leq i\leq j\leq n}a_{ij}x_{i}x_{j}$
-where $a_{ij}\in K$. The matrix $M_{Q}$ of the form $Q$ is defined as
-$M_{Q}=\begin{pmatrix}a_{11}&\frac{a_{12}}{2}&\dots&\frac{a_{1n}}{2}\\ \frac{a_{12}}{2}&a_{22}&\dots&\frac{a_{2n}}{2}\\ \vdots&\vdots&\ddots&\vdots\\ \frac{a_{1n}}{2}&\frac{a_{2n}}{2}&\dots&a_{nn}\end{pmatrix}.$
-The rank of $Q$ is the rank of the matrix $M_{Q}$ over $K$.
-If the coefficients $a_{ij}$ belong to $\mathcal{O}_{K}$, then $Q$ is called integral.
-If the matrix $M_{Q}$ has entries in $\mathcal{O}_{K}$, i.e., $a_{ij}\in 2\mathcal{O}_{K}$ for $i\neq j$, then $Q$ is called classical.
-Let $Q$ be a quadratic form over $\mathbb{Q}$.
-We say that $Q$ represents $b\in\mathbb{Z}$ over $\mathbb{Z}$ if there exists $\mathbf{a}\in\mathbb{Z}^{n}$ such that $Q(\mathbf{a})=b$.
-In particular, according to our definition, every quadratic form represents 00.
-If there exists $\mathbf{a}\in\mathbb{Z}^{n}\setminus\{(0,0,\dots,0)\}$ such that $Q(\mathbf{a})=0$, then we say that $Q$ represents 00 non-trivially.
-The form $Q$ is called positive definite if $M_{Q}$ is a positive definite matrix and $Q$ is called universal if it is a positive definite integral form such that it represents every $b\in\mathbb{Z}_{\geq 1}$.
-In the rest of the paper, we consider quadratic forms over $K$ only when $K=\mathbb{Q}$ or $K$ is a quadratic field.
-Let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}\setminus\{0,1\}$ is squarefree.
-If we let
-$\omega_{D}=\begin{cases}\sqrt{D},&\text{if }D\equiv 2,3\pmod{4},\\ \frac{1+\sqrt{D}}{2},&\text{if }D\equiv 1\pmod{4},\end{cases}$
-then $(1,\omega_{D})$ is a basis of $\mathcal{O}_{K}$. The quadratic field $K$ has two embeddings into $\mathbb{C}$: the identity and the conjugation map
-$\tau:K\hookrightarrow\mathbb{C},a+b\sqrt{D}\mapsto a-b\sqrt{D}.$
+We thank Nicolas Daans, Vítězslav Kala and Pavlo Yatsyna for their helpful suggestions and the anonymous referee for their constructive feedback.
+2. Generalized polynomials with rational values
+Throughout the article let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}\setminus\{0,1\}$ is squarefree. If we let
+[…]
+then $(1,\omega_{D})$ is a basis of $\mathcal{O}_{K}$.
 For $\alpha\in K$, we let $\operatorname{Tr}(\alpha)=\alpha+\tau(\alpha)$ and $\operatorname{N}(\alpha)=\alpha\tau(\alpha)$ be the trace and norm of $\alpha$, respectively.
-More generally, if $K$ is a number field with $r$ real and $2s$ complex embeddings
-$\sigma_{1},\sigma_{2},\dots,\sigma_{r}:K\hookrightarrow\mathbb{R},\sigma_{r+1},\bar{\sigma}_{r+1},\dots,\sigma_{r+s},\bar{\sigma}_{r+s}:K\hookrightarrow\mathbb{C},$
-then we let $V=\mathbb{R}^{r}\times\mathbb{C}^{s}$ and call $V$ the Minkowski space. The field $K$ embeds into $V$ via the map
-$\sigma:K\hookrightarrow V,\alpha\mapsto(\sigma_{1}(\alpha),\dots,\sigma_{r+s}(\alpha)).$
-We have $K\otimes_{\mathbb{Q}}\mathbb{R}\simeq V$ (see [Mi, p. 70]). In particular, $\sigma(K)$ is dense in $V$.
-3. Generalized polynomials with rational values
 We begin by defining generalized polynomials over a quadratic field.
@@
 (2) $g(z_{1},z_{2},\dots,z_{n})=\sum_{\begin{subarray}{c}i_{1},j_{1},\dots,i_{n},j_{n}\geq 0\\ i_{1}+j_{1}+\cdots+i_{n}+j_{n}\leq k\end{subarray}}\alpha_{i_{1},j_{1},\dots i_{n},j_{n}}z_{1}^{i_{1}}\tau(z_{1})^{j_{1}}\cdots z_{n}^{i_{n}}\tau(z_{n})^{j_{n}}$
-where $k\in\mathbb{Z}_{\geq 0}$ and $\alpha_{i_{1},j_{1},\dots,i_{n},j_{n}}\in K$ for every $i_{1},j_{1},\dots,i_{n},j_{n}\in\{0,1,\dots,k\}$.
-If $\mathbf{z}=(z_{1},z_{2},\dots,z_{n})$, then we let $\tau(\mathbf{z})=(\tau(z_{1}),\tau(z_{2}),\dots,\tau(z_{n}))$. Using the multi-index notation, we can write $g$ as
+where $k\in\mathbb{Z}_{\geq 0}$, $\alpha_{i_{1},j_{1},\dots,i_{n},j_{n}}\in K$ for every $i_{1},j_{1},\dots,i_{n},j_{n}\in\{0,1,\dots,k\}$ and $\tau:K\hookrightarrow\mathbb{C}$ where $\tau(a+b\sqrt{D})=a-b\sqrt{D}$ is the conjugation map.
+We will often use the multiindex notation, where $\mathbf{i}=(i_{1},i_{2},\dots,i_{n})\in\mathbb{Z}_{\geq 0}^{n}$ is a multiindex, $|\mathbf{i}|=\sum_{r=1}^{n}i_{r}$ is the sum of its components, and for $\mathbf{x}=(x_{1},x_{2},\dots,x_{n})$ we let $\mathbf{x}^{\mathbf{i}}=x_{1}^{i_{1}}x_{2}^{i_{2}}\cdots x_{n}^{i_{n}}$.
+The lexicographic order on $\mathbb{Z}_{\geq 0}^{n}$ is defined as follows: Let $\mathbf{i},\mathbf{j}\in\mathbb{Z}_{\geq 0}^{n}$, $\mathbf{i}=(i_{1},i_{2},\dots,i_{n})$ and $\mathbf{j}=(j_{1},j_{2},\dots,j_{n})$.
+Then $\mathbf{i}<\mathbf{j}$ if and only if there exists $s\in\{1,2,\dots,n\}$ such that $i_{r}=j_{r}$ for $1\leq r<s$ and $i_{s}<j_{s}$.
+Hence if $\mathbf{z}=(z_{1},z_{2},\dots,z_{n})$, then we let $\tau(\mathbf{z})=(\tau(z_{1}),\tau(z_{2}),\dots,\tau(z_{n}))$ and we can write the generalized polynomial $g$ as
 $g(\mathbf{z})=\sum_{|\mathbf{i}|+|\mathbf{j}|\leq k}\alpha_{\mathbf{i}\mathbf{j}}\mathbf{z}^{\mathbf{i}}\tau(\mathbf{z})^{\mathbf{j}}.$
@@
 $g(z_{1},z_{2},\dots,z_{n})=\tilde{g}(z_{1},\tau(z_{1}),\dots,z_{n},\tau(z_{n})).$
-First, suppose that $K$ is real ($r=2$ and $s=0$).
+First, suppose that $K$ is real.
 We know that $\sigma(K)=\{(\alpha,\tau(\alpha)):\;\alpha\in K\}$ is dense in $V=\mathbb{R}^{2}$.
@@
 Because the polynomial $\tilde{g}$ vanishes on $\sigma(K)^{n}$, it vanishes on the whole set $\mathbb{R}^{2n}$, and hence $\tilde{g}$ is the zero polynomial.
-Secondly, suppose that $K$ is imaginary ($r=0$ and $s=1$).
-In this case, $\sigma(K)=K$ is dense in $V=\mathbb{C}$.
-Thus $K^{n}$ is dense in $\mathbb{C}^{n}$ and the polynomial $\tilde{g}$ is zero on the set
+Secondly, suppose that $K$ is imaginary. In this case, $\sigma(K)=K$ is dense in $V=\mathbb{C}$. Thus $K^{n}$ is dense in $\mathbb{C}^{n}$ and the polynomial $\tilde{g}$ is zero on the set
 $\{(z_{1},\overline{z_{1}},\dots,z_{n},\overline{z_{n}}):\;z_{1},\dots,z_{n}\in K\}.$
@@
 If $d=0$, then $g(z)=a\in\mathbb{Q}$ is a constant polynomial.
-Thus we assume that $d\geq 1$ and that the $\mathbb{Q}$-valued generalized polynomials of degree $<d$ are contained in $S_{1}$.
+Thus we assume that $d\geq 1$ and that the $\mathbb{Q}$-valued generalized polynomials of degree <d<d are contained in $S_{1}$.
 We write $g$ as
@@
 $h_{1}(z)=\sum_{\begin{subarray}{c}i,j\geq 1\\ i+j\leq d\end{subarray}}\alpha_{i,j}z^{i-1}\tau(z)^{j-1}$
-is a $\mathbb{Q}$-valued generalized polynomial of degree $<d$. By induction assumption, $h_{1}\in S_{1}$, and because we also know that $z\tau(z)\in S_{1}$, we get $g_{1}\in S_{1}$.
+is a $\mathbb{Q}$-valued generalized polynomial of degree <d<d. By the inductive hypothesis, $h_{1}\in S_{1}$, and because we also know that $z\tau(z)\in S_{1}$, we get $g_{1}\in S_{1}$.
 Next, we have $g_{2}(z)=\alpha_{d,0}z^{d}+\tau(\alpha_{d,0})\tau(z)^{d}+h_{2}(z)$ where
 $h_{2}(z)=\sum_{i=0}^{d-1}\alpha_{i,0}z^{i}+\tau(\alpha_{i,0})\tau(z)^{i}$
-is a $\mathbb{Q}$-valued generalized polynomial of degree $<d$.
-By induction assumption, $h_{2}\in S_{2}$.
+is a $\mathbb{Q}$-valued generalized polynomial of degree <d<d.
+By the inductive hypothesis, $h_{2}\in S_{2}$.
 We show $\alpha_{d,0}z^{d}+\tau(\alpha_{d,0})\tau(z)^{d}\in S_{1}$.
@@
 $r_{1}(z)=\sum_{i=1}^{d-1}\binom{d}{i}z^{i-1}\tau(z)^{d-i-1}$
-is a $\mathbb{Q}$-valued generalized polynomial of degree $<d$, hence $r_{1}\in S_{1}$. Because $(z+\tau(z))^{d}$ and $z\tau(z)$ also belong to $S_{1}$, we get $z^{d}+\tau(z)^{d}\in S_{1}$.
+is a $\mathbb{Q}$-valued generalized polynomial of degree <d<d, hence $r_{1}\in S_{1}$. Because $(z+\tau(z))^{d}$ and $z\tau(z)$ also belong to $S_{1}$, we get $z^{d}+\tau(z)^{d}\in S_{1}$.
 We also have $\sqrt{D}(z^{d}-\tau(z)^{d})=\sqrt{D}(z-\tau(z))\cdot r_{2}(z)$ where
 $r_{2}(z)=\sum_{i=0}^{d-1}z^{d-1-i}\tau(z)^{i}$
-is a $\mathbb{Q}$-valued generalized polynomial of degree $<d$, hence $r_{2}\in S_{1}$.
+is a $\mathbb{Q}$-valued generalized polynomial of degree <d<d, hence $r_{2}\in S_{1}$.
 Because $\sqrt{D}(z-\tau(z))\in S_{1}$, we get $\sqrt{D}(z^{d}-\tau(z)^{d})\in S_{1}$.
@@
 We have $S_{n}\subset\Gamma_{K}^{n}$ because the generators of $S_{n}$ are $\mathbb{Q}$-valued, and we want to prove $\Gamma_{K}^{n}\subset S_{n}$.
-We proceed by induction on $n$. For $n=1$, this is Lemma 3.5. Suppose that $n\geq 2$ and the statement is true for polynomials with $<n$ variables.
+We proceed by induction on $n$. For $n=1$, this is Lemma 2.5. Suppose that $n\geq 2$ and the statement is true for polynomials with <n<n variables.
 Let $g\in\Gamma_{K}^{n}$.
@@
 If $d=0$, then $g(\mathbf{z})=a\in\mathbb{Q}$ is a constant polynomial.
-Thus we assume that $d\geq 1$ and that the $\mathbb{Q}$-valued generalized polynomials of degree $<d$ in $n$ variables are contained in $S_{n}$.
+Thus we assume that $d\geq 1$ and that the $\mathbb{Q}$-valued generalized polynomials of degree <d<d in $n$ variables are contained in $S_{n}$.
 The polynomial $g$ is a $\mathbb{Q}$-linear combination of the elements in the basis from Lemma 2.6.
@@
 $g_{\mathbf{j}\mathbf{i}}(\mathbf{z})$ $=\sqrt{D}\left(\mathbf{z}^{\mathbf{i}}\tau(\mathbf{z})^{\mathbf{j}}-\tau(\mathbf{z})^{\mathbf{i}}\mathbf{z}^{\mathbf{j}}\right)\in S_{n}.$
-If $i_{n}=0$ and $j_{n}=0$, then the variable $z_{n}$ does not appear in $g_{\mathbf{i}\mathbf{j}}$ and $g_{\mathbf{j}\mathbf{i}}$, hence $g_{\mathbf{i}\mathbf{j}},g_{\mathbf{j}\mathbf{i}}\in S_{n}$ by induction assumption.
+If $i_{n}=0$ and $j_{n}=0$, then the variable $z_{n}$ does not appear in $g_{\mathbf{i}\mathbf{j}}$ and $g_{\mathbf{j}\mathbf{i}}$, hence $g_{\mathbf{i}\mathbf{j}},g_{\mathbf{j}\mathbf{i}}\in S_{n}$ by the inductive hypothesis.
 If $i_{n}\geq 1$, then we let
@@
 $g_{\mathbf{j}\mathbf{i}}(\mathbf{z})$ $=\sqrt{D}\left(z_{n}h(\mathbf{z})-\tau(z_{n})\tau(h(\mathbf{z}))\right)=\frac{1}{2}\left(g^{1}_{\mathbf{j}\mathbf{i}}+g^{2}_{\mathbf{j}\mathbf{i}}\right).$
-Because $h(\mathbf{z})$ has degree $d-1$, we can apply induction on $h(\mathbf{z})+\tau(h(\mathbf{z}))$ and $\sqrt{D}(h(\mathbf{z})-\tau(h(\mathbf{z})))$. Since
+Because $h(\mathbf{z})$ has degree d1d-1, we can apply induction on $h(\mathbf{z})+\tau(h(\mathbf{z}))$ and $\sqrt{D}(h(\mathbf{z})-\tau(h(\mathbf{z})))$. Since
 $z_{n}+\tau(z_{n}),\ \sqrt{D}(z_{n}-\tau(z_{n})),\ h(\mathbf{z})+\tau(h(\mathbf{z})),\ \sqrt{D}(h(\mathbf{z})-\tau(h(\mathbf{z})))\in S_{n},$
@@
 Next, we show that if $\mathbf{i}\in\mathbb{Z}_{\geq 0}^{n}$ and $2|\mathbf{i}|=d$, then $g_{\mathbf{i}\mathbf{i}}(\mathbf{z})=\mathbf{z}^{\mathbf{i}}\tau(\mathbf{z})^{\mathbf{i}}\in S_{n}$.
-If $i_{n}=0$, then $\mathbf{g}_{ii}$ is a polynomial in $n-1$ variables, hence $\mathbf{g}_{ii}\in S_{n}$ by induction assumption.
+If $i_{n}=0$, then $\mathbf{g}_{ii}$ is a polynomial in n1n-1 variables, hence $\mathbf{g}_{ii}\in S_{n}$ by the inductive hypothesis.
 If $i_{n}\geq 1$, then we let
@@
 $z_{n}\tau(z_{n})\in\mathbb{Q}\left[z_{n}+\tau(z_{n}),\sqrt{D}\left(z_{n}-\tau(z_{n})\right)\right]\subset S_{n}.$
-The generalized polynomial $h$ has degree $d-2$, hence $h\in S_{n}$ by induction assumption, and $g_{\mathbf{i}\mathbf{i}}\in S_{n}$ follows. ∎
+The generalized polynomial $h$ has degree d2d-2, hence $h\in S_{n}$ by the inductive hypothesis, and $g_{\mathbf{i}\mathbf{i}}\in S_{n}$ follows. ∎
 3. Generalized quadratic forms
@@
 An integral generalized quadratic form will be called classical if $\alpha_{ij},\gamma_{ij}\in 2\mathcal{O}_{K}$ for $i<j$ and $\beta_{ij}\in 2\mathcal{O}_{K}$ for $1\leq i,j\leq n$.
-If $G$ is a generalized quadratic form in $n$ variables over $K$, then we can define a quadratic form $Q$ in $2n$ variables by
+Let us recall that a quadratic form $Q$ in $n$ variables over $K$ is a homogeneous polynomial of degree $2$ over $K$.
+We can also obtain it as a generalized quadratic form with $\beta_{ij},\gamma_{ij}=0$ for all $1\leq i,j\leq n.$
+Then if $G$ is a generalized quadratic form in $n$ variables over $K$, then we can define a quadratic form $Q$ associated to $G$ in $2n$ variables by
 (4) $Q(x_{1},y_{1},\dots,x_{n},y_{n})=G(x_{1}+y_{1}\omega_{D},\dots,x_{n}+y_{n}\omega_{D}).$
 In other words, $Q$ is obtained from $G$ by expressing the variables $z_{1},z_{2},\dots,z_{n}$ in the integral basis $(1,\omega_{D})$.
-We call $Q$ the quadratic form associated to $G$.
 It has coefficients in $K$ and the variables $x_{1},y_{1},\dots,x_{n},y_{n}$ take values in $\mathbb{Q}$.
@@
 $G(\mathbf{z})=(\mathbf{z},\tau(\mathbf{z}))M_{G}\begin{pmatrix}\mathbf{z}\\ \tau(\mathbf{z})\end{pmatrix}$
-for $\mathbf{z}\in K^{n}$.
+for $\mathbf{z}\in K^{n}$. The rank of $G$ is the rank of the matrix $M_{G}$ over $K$.
 The $2n\times 2n$ matrix
@@
 determines the change of variables $(x_{1},y_{1},\dots,x_{n},y_{n})\mapsto(z_{1},\dots,z_{n},\tau(z_{1}),\dots,\tau(z_{n}))$.
-Thus, if $M_{Q}$ is the matrix of the associated quadratic form $Q$ defined by (4), then
+Thus, if is the quadratic form $Q$ is associated to the quadratic form $G$ defined by (4), and $M_{Q},M_{G}$ are their respective matrices, then
 $M_{Q}=T^{\top}M_{G}T.$
+If $G$ is a generalized quadratic form over $\mathbb{Q}$ then we say that $G$ represents $b\in\mathbb{Z}$ over $\mathbb{Z}$ if there exists $\mathbf{a}\in\mathbb{Z}^{n}$ such that $G(\mathbf{a})=b$.
+In particular, according to our definition, every quadratic form represents 0.
+If there exists $\mathbf{a}\in\mathbb{Z}^{n}\setminus\{(0,0,\dots,0)\}$ such that $G(\mathbf{a})=0$, then we say that $G$ represents 0 non-trivially.
+The form $G$ is called positive definite if $M_{G}$ is a positive definite matrix and $G$ is called universal if it is a positive definite integral form such that it represents every $b\in\mathbb{Z}_{\geq 1}$.
 Lemma 3.2.
@@
 If $Q$ is the quadratic form in $2n$ variables associated to $G$ and $M_{Q}$ is the matrix of $Q$, then
-$\det(M_{Q})=D^{n}\cdot\begin{cases}\det(2M_{G}),&\text{if }D\equiv 2,3\pmod{4},\\ \det(M_{G}),&\text{if }D\equiv 1\pmod{4}.\end{cases}$
+[…]
 Proof.
@@
 Thus
-$\det T=(-1)^{\frac{n(n-1)}{2}}(\det M)^{n}=\begin{cases}(-1)^{\frac{n(n+1)}{2}}D^{\frac{n}{2}}2^{n},&\text{if }D\equiv 2,3\pmod{4},\\ (-1)^{\frac{n(n+1)}{2}}D^{\frac{n}{2}},&\text{if }D\equiv 1\pmod{4}.\end{cases}$
+[…]
 In the case $D\equiv 2,3\pmod{4}$, we get
@@
 Let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}\setminus\{0,1\}$ is squarefree and let $G$ be a generalized form in $n$ variables given by (3).
-Assume that $G(\mathbf{a})\in\mathbb{Q}$ for every $\mathbf{a}\in K^{2}$.
+Assume that $G(\mathbf{a})\in\mathbb{Q}$ for every $\mathbf{a}\in K^{n}$.
 If $1\leq i\leq j\leq n$, then we let $\alpha_{ij}=r_{ij}+s_{ij}\omega_{D}$ and $\beta_{ij}=t_{ij}+u_{ij}\omega_{D}$ where $r_{ij},s_{ij},t_{ij},u_{ij}\in\mathbb{Q}$.
@@
 Proof.
-We introduce the following notation: if $M$ is a $2n\times 2n$ matrix and $1\leq k\leq 2n$, let $[M]_{k}$ denote the $k$-th row of $M$.
+We introduce the following notation: if $M$ is a $2n\times 2n$ matrix and $1\leq k\leq 2n$, let $M]_{k}$ denote the $k$-th row of $M$.
 If $1\leq i\leq j\leq n$, then we let $\alpha_{ij}=r_{ij}+s_{ij}\omega_{D}$ and $\beta_{ij}=t_{ij}+u_{ij}\omega_{D}$ where $r_{ij},s_{ij},t_{ij},u_{ij}\in\mathbb{Q}$.
 First assume that $D\equiv 2,3\pmod{4}$. If $1\leq k\leq n$, then the $2k$-th row of $M_{Q}$ equals
-$[M_{Q}]_{2k}=\begin{pmatrix}\frac{b_{1k}}{2}&\frac{c_{1k}}{2}&\dots&\frac{b_{kk}}{2}&c_{kk}&\dots&\frac{b_{nk}}{2}&\frac{c_{kn}}{2}\end{pmatrix}.$
+$M_{Q}]_{2k}=\begin{pmatrix}\frac{b_{1k}}{2}&\frac{c_{1k}}{2}&\dots&\frac{b_{kk}}{2}&c_{kk}&\dots&\frac{b_{nk}}{2}&\frac{c_{kn}}{2}\end{pmatrix}$.
 By Lemma 3.6, if $1\leq i<j\leq n$, then
@@
 Since we assume that $G$ is integral, $r_{ij},s_{ij},t_{ij},u_{ij}\in\mathbb{Z}$.
-Thus $D$ divides every element of $[M_{Q}]_{2k}$.
+Thus $D$ divides every element of $M_{Q}]_{2k}$.
 It follows that $M_{Q}\bmod{p}$ contains $n$ zero rows, hence $\operatorname{rank}(M_{Q}\bmod{p})\leq n$.
 Secondly, assume that $D\equiv 1\pmod{4}$.
-Let $N_{Q}$ be the matrix obtained from $M_{Q}$ by subtracting two times the $2k$-th row of $M_{Q}$ from the $(2k-1)$-th row of $M_{Q}$ for every $1\leq k\leq n$.
-The $(2k-1)$-th row of $N_{Q}$ equals
-$[M_{Q}]_{2k-1}-2[M_{Q}]_{2k}=\begin{pmatrix}\frac{a_{1k}}{2}-b_{1k}&\frac{b_{k1}}{2}-c_{1k}&\dots&a_{kk}-b_{kk}&\frac{b_{kk}}{2}-2c_{kk}&\dots&\frac{a_{kn}}{2}-b_{nk}&\frac{b_{kn}}{2}-c_{kn}\end{pmatrix}.$
+Let $N_{Q}$ be the matrix obtained from $M_{Q}$ by subtracting two times the $2k$-th row of $M_{Q}$ from the (2k1)(2k-1)-th row of $M_{Q}$ for every $1\leq k\leq n$.
+The (2k1)(2k-1)-th row of $N_{Q}$ equals
+$M_{Q}]_{2k-1}-2[M_{Q}]_{2k}=\begin{pmatrix}\frac{a_{1k}}{2}-b_{1k}&\frac{b_{k1}}{2}-c_{1k}&\dots&a_{kk}-b_{kk}&\frac{b_{kk}}{2}-2c_{kk}&\dots&\frac{a_{kn}}{2}-b_{nk}&\frac{b_{kn}}{2}-c_{kn}\end{pmatrix}$.
 By Lemma 3.6, if $1\leq i<j\leq n$, then
@@
 Since we assume $D\equiv 1\pmod{4}$, $p$ is an odd prime and $2$ is invertible modulo $p$.
-We see that every element of $[N_{Q}]_{2k-1}$ vanishes modulo $p$, hence $N_{Q}\bmod{p}$ contains $n$ zero rows and $\operatorname{rank}(M_{Q}\bmod{p})=\operatorname{rank}(N_{Q}\bmod{p})\leq n$.
+We see that every element of $N_{Q}]_{2k-1}$ vanishes modulo $p$, hence $N_{Q}\bmod{p}$ contains $n$ zero rows and $\operatorname{rank}(M_{Q}\bmod{p})=\operatorname{rank}(N_{Q}\bmod{p})\leq n$.
 ∎
@@
 $c_{22}$ $=\frac{1+D}{2}r_{22}+\frac{1+3D}{4}s_{22}+\frac{1-D}{4}t_{22}=\frac{1+D}{2}c_{1}+\frac{1+3D}{4}c_{2}+\frac{1-D}{4}f.\qed$
-In Lemmas 5.4 to 5.9, we find an example of a $\mathbb{Z}$-universal generalized form over $\mathbb{Q}(\sqrt{D})$ for $D\in\{2,3,6,7,10\}$ and of a classical $\mathbb{Z}$-universal generalized form over $\mathbb{Q}(\sqrt{5})$.
+In Lemmas 4.4 to 4.9, we find an example of a $\mathbb{Z}$-universal generalized form over $\mathbb{Q}(\sqrt{D})$ for $D\in\{2,3,6,7,10\}$ and of a classical $\mathbb{Z}$-universal generalized form over $\mathbb{Q}(\sqrt{5})$.
+We prove only Lemma 4.4, the rest is stated and proved in the Appendix in Section A4.
 Lemma 4.4.
@@
 Proof.
-The coefficients of $G_{2}$ are $a=1$, $b=-\sqrt{2}$, $c=1$, $d=-1$, $e=0$, $f=-1$.
+The coefficients of $G_{2}$ are $a=1$, $b=-\sqrt{2}$, $c=1$, d=1d=-1, $e=0$, f=1f=-1.
 Let $Q_{2}$ be the quaternary quadratic form $Q_{2}(x_{1},y_{1},x_{2},y_{2})=G_{2}(x_{1}+y_{1}\sqrt{2},x_{2}+y_{2}\sqrt{2})$ and let $M_{Q_{2}}$ be the matrix of $Q_{2}$.
@@
 The generalized form $G_{3}$ is $\mathbb{Z}$-universal over $K$.
-Proof.
-The coefficients of $G_{3}$ are $a=2$, $b=-2\sqrt{3}$, $c=2$, $d=-3$, $e=-\sqrt{3}$, $f=0$.
-Let $Q_{3}$ be the quaternary quadratic form $Q_{3}(x_{1},y_{1},x_{2},y_{2})=G_{3}(x_{1}+y_{1}\sqrt{3},x_{2}+y_{2}\sqrt{3})$.
-By Lemma 5.3 with $D=3$,
-$M_{Q_{3}}=\begin{pmatrix}2a_{1}+d&6a_{2}&b_{1}+e_{1}&3(b_{2}-e_{2})\\ 6a_{2}&3(2a_{1}-d)&3(b_{2}+e_{2})&3(b_{1}-e_{1})\\ b_{1}+e_{1}&3(b_{2}-e_{2})&2c_{1}+f&6c_{2}\\ 3(b_{2}-e_{2})&3(b_{1}-e_{1})&6c_{2}&3(2c_{1}-f)\end{pmatrix}=\begin{pmatrix}1&0&0&-3\\ 0&21&-9&0\\ 0&-9&4&0\\ -3&0&0&12\end{pmatrix}.$
-We get
-$M_{Q_{3}}\sim\begin{pmatrix}1&0&0&0\\ 0&21&-9&0\\ 0&-9&4&0\\ 0&0&0&3\end{pmatrix}\sim\begin{pmatrix}1&0&0&0\\ 0&1&-1&0\\ 0&-1&4&0\\ 0&0&0&3\end{pmatrix}\sim\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&3&0\\ 0&0&0&3\end{pmatrix},$
-The diagonal quadratic form $x^{2}+y^{2}+3z^{2}+3w^{2}$ is positive definite and universal by the 15-Theorem, hence $G_{3}$ is also positive definite and $\mathbb{Z}$-universal. ∎
 Lemma 4.6.
@@
 where $\omega_{5}=\frac{1+\sqrt{5}}{2}$. The generalized form $G_{5}$ is classical and $\mathbb{Z}$-universal over $K$.
-Proof.
-The coefficients of $G_{5}$ are $a=3-\omega_{5}$, $b=0$, $c=1$, $d=-4$, $e=0$, and $f=0$.
-Since the coefficients of the off-diagonal terms are even, $G_{5}$ is classical.
-Let $Q_{5}$ be the quaternary quadratic form $Q_{5}(x_{1},y_{1},x_{2},y_{2})=G_{5}(x_{1}+y_{1}\omega_{5},x_{2}+y_{2}\omega_{5})$.
-By Lemma 5.3 with $D=5$,
-$M_{Q_{5}}=\begin{pmatrix}M_{Q_{5},1}&M_{Q_{5},2}\\ M_{Q_{5},2}^{\top}&M_{Q_{5},3}\end{pmatrix}$
-where
-$M_{Q_{5},1}$ $=\begin{pmatrix}2a_{1}+a_{2}+d&a_{1}+3a_{2}+\frac{d}{2}\\ a_{1}+3a_{2}+\frac{d}{2}&3a_{1}+4a_{2}-d\end{pmatrix}=\begin{pmatrix}1&-2\\ -2&9\end{pmatrix},$
-$M_{Q_{5},2}$ $=\begin{pmatrix}b_{1}+\frac{b_{2}}{2}+e_{1}+\frac{e_{2}}{2}&\frac{b_{1}}{2}+\frac{3}{2}b_{2}+\frac{e_{1}}{2}-e_{2}\\ \frac{b_{1}}{2}+\frac{3}{2}b_{2}+\frac{e_{1}}{2}+\frac{3}{2}e_{2}&\frac{3}{2}b_{1}+2b_{2}-e_{1}-\frac{1}{2}e_{2}\end{pmatrix}=\begin{pmatrix}0&0\\ 0&0\end{pmatrix},$
-$M_{Q_{5},3}$ $=\begin{pmatrix}2c_{1}+c_{2}+f&c_{1}+3c_{2}+\frac{f}{2}\\ c_{1}+3c_{2}+\frac{f}{2}&3c_{1}+4c_{2}-f\end{pmatrix}=\begin{pmatrix}2&1\\ 1&3\end{pmatrix},$
-thus
-$M_{Q_{5}}=\begin{pmatrix}1&-2&0&0\\ -2&9&0&0\\ 0&0&2&1\\ 0&0&1&3\end{pmatrix}\sim\begin{pmatrix}1&0&0&0\\ 0&5&0&0\\ 0&0&2&1\\ 0&0&1&3\end{pmatrix}.$
-This quadratic form is positive definite and universal by the 15-Theorem. Alternatively, it is possible to consult [Bh, Table 5], which contains the equivalent form
-$\begin{pmatrix}1&0&0&0\\ 0&2&1&0\\ 0&1&3&0\\ 0&0&0&5\end{pmatrix}$
-encoded as “25: 2 3 5 0 0 2:2523500225:\ 2\ 3\ 5\ 0\ 0\ 225 : 2 3 5 0 0 2”. ∎
 Lemma 4.7.
@@
 The generalized form $G_{6}$ is $\mathbb{Z}$-universal over $K$.
-Proof.
-The coefficients of $G_{6}$ are $a=3$, $b=2\sqrt{6}$, $c=3$, $d=-5$, $e=\sqrt{6}$, and $f=-1$.
-Let $Q_{6}$ be the quaternary quadratic form $M_{Q_{6}}(x_{1},y_{1},x_{2},y_{2})=G_{6}(x_{1}+y_{1}\sqrt{6},x_{2}+y_{2}\sqrt{6})$.
-By Lemma 5.3 with $D=6$,
-$M_{Q_{6}}=\begin{pmatrix}2a_{1}+d&12a_{2}&b_{1}+e_{1}&6(b_{2}-e_{2})\\ 12a_{2}&6(2a_{1}-d)&6(b_{2}+e_{2})&6(b_{1}-e_{1})\\ b_{1}+e_{1}&6(b_{2}+e_{2})&2c_{1}+f&12c_{2}\\ 6(b_{2}-e_{2})&6(b_{1}-e_{1})&12c_{2}&6(2c_{1}-f)\end{pmatrix}=\begin{pmatrix}1&0&0&6\\ 0&66&18&0\\ 0&18&5&0\\ 6&0&0&42\end{pmatrix}.$
-We get
-$M_{Q_{6}}\sim\begin{pmatrix}1&0&0&0\\ 0&66&18&0\\ 0&18&5&0\\ 0&0&0&6\end{pmatrix}\sim\begin{pmatrix}1&0&0&0\\ 0&3&3&0\\ 0&3&5&0\\ 0&0&0&6\end{pmatrix}\sim\begin{pmatrix}1&0&0&0\\ 0&3&0&0\\ 0&0&2&0\\ 0&0&0&6\end{pmatrix}.$
-The last quadratic form is positive definite and universal by the 15-Theorem. ∎
 Lemma 4.8.
@@
 The generalized form $G_{7}$ is $\mathbb{Z}$-universal over $K$.
-Proof.
-The coefficients of $G_{7}$ are $a=3$, $b=1+2\sqrt{7}$, $c=4$, $d=-5$, $e=-1+2\sqrt{7}$, and $f=3$. By Lemma 5.3 with $D=7$,
-$M_{Q_{7}}=\begin{pmatrix}2a_{1}+d&14a_{2}&b_{1}+e_{1}&7(b_{2}-e_{2})\\ 14a_{2}&7(2a_{1}-d)&7(b_{2}+e_{2})&7(b_{1}-e_{1})\\ b_{1}+e_{1}&7(b_{2}+e_{2})&2c_{1}+f&14c_{2}\\ 7(b_{2}-e_{2})&7(b_{1}-e_{1})&14c_{2}&7(2c_{1}-f)\end{pmatrix}=\begin{pmatrix}1&0&0&0\\ 0&77&28&14\\ 0&28&11&0\\ 0&14&0&35\end{pmatrix}.$
-Let
-$A_{7}=\begin{pmatrix}1&0&0&0\\ 0&2&-3&3\\ 0&-5&8&-7\\ 0&-1&1&-1\end{pmatrix}.$
-It can be verified that $\det(A_{7})=1$, hence $A_{7}\in\operatorname{SL}_{4}(\mathbb{Z})$. The equivalent form with the matrix
-$A_{7}^{\top}M_{Q_{7}}A_{7}=\begin{pmatrix}1&0&0&0\\ 0&2&-3&3\\ 0&-5&8&-7\\ 0&-1&1&-1\end{pmatrix}^{\top}\begin{pmatrix}1&0&0&0\\ 0&77&28&14\\ 0&28&11&0\\ 0&14&0&35\end{pmatrix}\begin{pmatrix}1&0&0&0\\ 0&2&-3&3\\ 0&-5&8&-7\\ 0&-1&1&-1\end{pmatrix}=\begin{pmatrix}1&0&0&0\\ 0&2&1&0\\ 0&1&4&0\\ 0&0&0&7\end{pmatrix}$
-is positive definite and universal by the 15-Theorem. Alternatively, it can be found in [Bh, Table 5] as “49: 2 4 7 0 0 2”. ∎
 Lemma 4.9.
@@
 then $G_{10}$ is $\mathbb{Z}$-universal over $K$.
-Proof.
-The coefficients of $G_{10}$ are $a=3$, $b=\sqrt{10}$, $c=5$, $d=-5$, $e=\sqrt{10}$, and $f=7$.
-Let $Q_{10}$ be the quaternary quadratic form $Q_{10}(x_{1},y_{1},x_{2},y_{2})=G_{10}(x_{1}+y_{1}\sqrt{10},x_{2}+y_{2}\sqrt{10})$.
-By Lemma 5.3 with $D=10$,
-$M_{Q_{10}}=\begin{pmatrix}2a_{1}+d&20a_{2}&b_{1}+e_{1}&10(b_{2}-e_{2})\\ 20a_{2}&10(2a_{1}-d)&10(b_{2}+e_{2})&10(b_{1}-e_{1})\\ b_{1}+e_{1}&10(b_{2}+e_{2})&2c_{1}+f&20c_{2}\\ 10(b_{2}-e_{2})&10(b_{1}-e_{1})&20c_{2}&10(2c_{1}-f)\end{pmatrix}=\begin{pmatrix}1&0&0&0\\ 0&110&20&0\\ 0&20&17&20\\ 0&0&20&30\end{pmatrix}.$
-Let
-$A_{10}=\begin{pmatrix}1&0&0&0\\ 0&1&1&2\\ 0&-6&-5&-10\\ 0&4&3&7\end{pmatrix}.$
-It can be verified that $\det(A_{10})=1$, hence $A\in\operatorname{SL}_{4}(\mathbb{Z})$. The equivalent form with the matrix
-$A_{10}^{\top}M_{Q_{10}}A_{10}=\begin{pmatrix}1&0&0&0\\ 0&1&1&2\\ 0&-6&-5&-10\\ 0&4&3&7\end{pmatrix}^{\top}\begin{pmatrix}1&0&0&0\\ 0&110&20&0\\ 0&20&17&20\\ 0&0&20&30\end{pmatrix}\begin{pmatrix}1&0&0&0\\ 0&1&1&2\\ 0&-6&-5&-10\\ 0&4&3&7\end{pmatrix}=\begin{pmatrix}1&0&0&0\\ 0&2&0&0\\ 0&0&5&0\\ 0&0&0&10\end{pmatrix}$
-is positive definite and universal by the 15-Theorem. ∎
 Now we are ready to prove Theorem 1.2, which we restate here for the convenience of the reader.
@@
 If a binary $\mathbb{Z}$-universal generalized form exists over $K=\mathbb{Q}(\sqrt{D})$, then $D\in\{2,3,6,7,10\}$ by Proposition 4.2 (i).
-Conversely, in Lemmas 5.4, 5.5, 5.7, 5.8 and 5.9 we found an example of a generalized form with the required properties over each of these fields.
+Conversely, in Lemmas 4.4, 4.5, 4.7, 4.8 and 4.9 we found an example of a generalized form with the required properties over each of these fields.
 Next, we prove (ii).
@@
 Every squarefree $D\in\{1,\dots,62\}$ satisfies this condition.
-Additionally, the squarefree determinants greater than $62$ on the list are $66,70,74,77,78,82,85,86,87,93,94,95,102,106,110$.
+Additionally, the squarefree determinants greater than $62$ on the list are 66,70,74,77,78,82,85,86,87,93,94,95,102,106,11066,70,74,77,78,82,85,86,87,93,94,95,102,106,110.
 ∎
@@
 We show that there exists a positive constant $C$ such that $D\leq C$ for every such field $K=\mathbb{Q}(\sqrt{D})$.
-If $D\equiv 2,3\pmod{4}$, then the statement follows from Theorems 5.10 and 5.11.
+If $D\equiv 2,3\pmod{4}$, then the statement follows from Theorems 4.10 and 4.11.
 Hence, we can assume that $D\equiv 1\pmod{4}$.
@@
 This quadratic form is universal by Lagrange’s Four Square Theorem. ∎
-We end this section with a proposition that will be used in Section 7.
+We end this section with a proposition that is used in Section A6.
 Proposition 4.14.
@@
 where $a,b,c,e\in\mathcal{O}_{K}$ and $d,f\in\mathbb{Z}$ because $G$ is integral. Expanding $\det(2M_{G})$ along the first column, we get
-$\det(2M_{G})=2a\begin{vmatrix}2c&\tau(e)&f\\ \tau(e)&2\tau(a)&\tau(b)\\ f&\tau(b)&2\tau(c)\end{vmatrix}-b\begin{vmatrix}b&d&e\\ \tau(e)&2\tau(a)&\tau(b)\\ f&\tau(b)&2\tau(c)\end{vmatrix}+d\begin{vmatrix}b&d&e\\ 2c&\tau(e)&f\\ f&\tau(b)&2\tau(c)\end{vmatrix}-e\begin{vmatrix}b&d&e\\ 2c&\tau(e)&f\\ \tau(e)&2\tau(a)&\tau(b)\end{vmatrix},$
+$\det(2M_{G})=$ $2a\cdot\det\begin{pmatrix}2c&\tau(e)&f\\ \tau(e)&2\tau(a)&\tau(b)\\ f&\tau(b)&2\tau(c)\end{pmatrix}-b\cdot\det\begin{pmatrix}b&d&e\\ \tau(e)&2\tau(a)&\tau(b)\\ f&\tau(b)&2\tau(c)\end{pmatrix}$+
+$+d\cdot\det\begin{pmatrix}b&d&e\\ 2c&\tau(e)&f\\ f&\tau(b)&2\tau(c)\end{pmatrix}-e\cdot\det\begin{pmatrix}b&d&e\\ 2c&\tau(e)&f\\ \tau(e)&2\tau(a)&\tau(b)\end{pmatrix}$,
 hence
-$\det(2M_{G})\equiv$ $2a\left[\tau(e)\tau(b)f+f\tau(e)\tau(b)\right]-b\left[d\tau(b)f+e\tau(e)\tau(b)-2f\tau(a)e-\tau(b)^{2}b-2\tau(c)\tau(e)d\right]$
-$+d\left[2b\tau(e)\tau(c)+df^{2}+2ec\tau(b)-f\tau(e)e-\tau(b)fb\right]$
-$-e\left[b\tau(e)\tau(b)+df\tau(e)-\tau(e)^{2}e-2\tau(a)fb-2\tau(b)cd\right]$
+$\det(2M_{G})\equiv$ $2a\Big(\tau(e)\tau(b)f+f\tau(e)\tau(b)\Big)-b\Big(d\tau(b)f+e\tau(e)\tau(b)-2f\tau(a)e-\tau(b)^{2}b-2\tau(c)\tau(e)d\Big)$
+$+d\Big(2b\tau(e)\tau(c)+df^{2}+2ec\tau(b)-f\tau(e)e-\tau(b)fb\Big)$
+$-e\Big(b\tau(e)\tau(b)+df\tau(e)-\tau(e)^{2}e-2\tau(a)fb-2\tau(b)cd\Big)$
 $\equiv$ $-\operatorname{N}(b)df-\operatorname{N}(b)\operatorname{N}(e)+2\tau(a)bef+\operatorname{N}(b)^{2}+2b\tau(c)d\tau(e)$
@@
 Thus,
-$\frac{\det(M_{Q})}{D^{2}}=\det(2M_{G})\equiv 0,1\pmod{4}.\qed$
+$\frac{\det(M_{Q})}{D^{2}}=\det(2M_{G})\equiv 0,1\pmod{4}$$\qed$
 5. Indefinite generalized quadratic forms
-This section focuses on proving the universality of $G(z,w)=z\tau(z)+w\tau(w)$ over real quadratic fields, where the form is indefinite.
-One of the classical methods for studying representations by quadratic forms involves using the local-global principle.
-We will use the following version of the local-global principle for indefinite forms.
-If $p$ is a prime number, then $\mathbb{Z}_{p}$ denotes the ring of $p$-adic integers.
-Theorem 6.1 ([Ca, Chap. 9, Thm. 1.5]).
-Let $Q$ be a regular indefinite integral form in $n\geq 4$ variables and let $a\neq 0$ be an integer.
-Suppose that $a$ is represented by $Q$ over all $\mathbb{Z}_{p}$.
-Then $a$ is represented by $Q$ over $\mathbb{Z}$.
-Recall that by our definition, every quadratic form $Q$ represents 00 (trivially).
-To represent nonzero $a$, we initially find the representation modulo $p$, followed by lifting it to a representation over $\mathbb{Z}_{p}$ using the following corollaries of Hensel’s lemma.
-The tuple $(x_{1},x_{2},\dots,x_{n})\in\mathbb{Z}_{p}^{n}$ is called primitive if at least one of the $x_{i}$ is invertible in $\mathbb{Z}_{p}$.
-Lemma 6.2 ([Se, Chap. 2, Cor. 2]).
-Suppose $p\neq 2$.
-Let $Q(x)=\sum a_{ij}x_{i}x_{j}$ with $a_{ij}=a_{ji}$ be a quadratic form with coefficients in $\mathbb{Z}_{p}$ whose determinant $\det(a_{ij})$ is invertible.
-Let $a\in\mathbb{Z}_{p}$.
-Every primitive solution of the equation $Q(x)\equiv a\pmod{p}$ lifts to a true solution.
-Lemma 6.3 ([Se, Chap. 2, Cor. 3]).
-Suppose $p=2$.
-Let $Q(x)=\sum a_{ij}x_{i}x_{j}$ with $a_{ij}=a_{ji}$ be a quadratic form with coefficients in $\mathbb{Z}_{2}$ and let $a\in\mathbb{Z}_{2}$.
-Let $x$ be a primitive solution of $Q(x)\equiv a\pmod{8}$.
-We can lift $x$ to a true solution provided $x$ does not annihilate all the $\frac{\partial Q}{\partial x_{j}}$ modulo $4$; this last condition is fulfilled if $\det(a_{ij})$ is invertible.
-We proceed with proving that $x^{2}+y^{2}-Dz^{2}-Dw^{2}$ represents every $a\in\mathbb{Z}$.
-Lemma 6.4.
-Let $p\neq 2$ be a prime, $D\in\mathbb{Z}$ squarefree, and $a\in\mathbb{Z}_{p}$. The quadratic form $x^{2}+y^{2}-Dz^{2}-Dw^{2}$ represents $a$ over $\mathbb{Z}_{p}$.
-Proof.
-We can assume $a\neq 0$. First, suppose $a=p^{2k}a_{1}$ where $k\in\mathbb{Z}_{\geq 0}$ and $p\nmid a_{1}$. The congruence
-$x^{2}+y^{2}\equiv a_{1}\pmod{p}$
-has a solution, hence by Lemma 6.2, there exist $x_{1},y_{1}\in\mathbb{Z}_{p}$ such that
-$x_{1}^{2}+y_{1}^{2}=a_{1}.$
-The tuple $(x,y,z,w)=(p^{k}x_{1},p^{k}y_{1},0,0)$ is a solution of the equation $x^{2}+y^{2}-Dz^{2}-Dw^{2}=a$.
-Secondly, suppose $a=p^{2k+1}a_{1}$ where $k\in\mathbb{Z}_{\geq 0}$ and $p\nmid a_{1}$. We distinguish two cases.
-If $p\nmid D$, then the congruence
-$x^{2}+y^{2}\equiv pa_{1}+D\pmod{p}$
-has a solution, hence by Lemma 6.2, there exist $x_{1},y_{1}\in\mathbb{Z}_{p}$ such that
-$x_{1}^{2}+y_{1}^{2}=pa_{1}+D.$
-The tuple $(x,y,z,w)=(p^{k}x_{1},p^{k}y_{1},p^{k},0)$ is a solution of the original equation because
-$x^{2}+y^{2}-Dz^{2}-Dw^{2}=p^{2k}(x_{1}^{2}+z_{1}^{2}-D)=p^{2k}\cdot pa_{1}=a.$
-If $p\mid D$, let $D_{p}=\frac{D}{p}$. We have $p\nmid D_{p}$ because $D$ is squarefree. The congruence
-$-D_{p}(z^{2}+w^{2})\equiv a_{1}\pmod{p}$
-has a solution, hence by Lemma 6.2, there exist $z_{1},w_{1}\in\mathbb{Z}_{p}$ such that
-$-D_{p}(z_{1}^{2}+w_{1}^{2})=a_{1}.$
-The tuple $(x,y,z,w)=(0,0,p^{k}z_{1},p^{k}w_{1})$ is a solution because
-$-Dz^{2}-Dw^{2}=p^{2k+1}(-D_{p})(z_{1}^{2}+w_{1}^{2})=p^{2k+1}a_{1}=a.\qed$
-Lemma 6.5.
-Let $D\in\mathbb{Z}$ be squarefree and $a\in\mathbb{Z}_{2}$. The quadratic form $x^{2}+y^{2}-Dz^{2}-Dw^{2}$ represents $a$ over $\mathbb{Z}_{2}$.
-Proof.
-We can assume $a\neq 0$. First, suppose that $4\nmid a$, hence $a\equiv 1,2,3,5,6,7\pmod{8}$. We find a solution to the congruence
-$x^{2}+y^{2}-Dz^{2}-Dw^{2}\equiv a\pmod{8}$
-where $x$ is odd, hence $\frac{\partial Q}{\partial x}=2x\neq 0\pmod{4}$. This solution lifts to a true solution in $\mathbb{Z}_{2}$ by Lemma 6.3.
-Let
-$(x,y)=\begin{cases}(1,0),&\text{if }a\equiv 1\pmod{8},\\ (1,1),&\text{if }a\equiv 2\pmod{8},\\ (1,2),&\text{if }a\equiv 5\pmod{8}.\end{cases}$
-We have $x^{2}+y^{2}\equiv a\pmod{8}$ in each case.
-If $a\equiv 3\pmod{8}$, let
-$(z,w)=\begin{cases}(1,1),&\text{if }D\equiv 1,3,5\pmod{8},\\ (1,0),&\text{if }D\equiv 2,6,7\pmod{8},\end{cases}$
-if $a\equiv 6\pmod{8}$, let
-$(z,w)=\begin{cases}(2,0),&\text{if }D\equiv 1,3,5,7\pmod{8},\\ (1,1),&\text{if }D\equiv 2,6\pmod{8},\end{cases}$
-and if $a\equiv 7\pmod{8}$, let
-$(z,w)=\begin{cases}(1,1),&\text{if }D\equiv 1,3,5,7\pmod{8},\\ (1,0),&\text{if }D\equiv 2,6\pmod{8}.\end{cases}$
-In each of the above cases, $a+Dz^{2}+Dw^{2}\pmod{8}\in\{1,2,5\}$, hence we can find $(x,y)=(1,y)$ such that
-$x^{2}+y^{2}\equiv a+Dz^{2}+Dw^{2}\pmod{8}.$
-Secondly, assume $4\mid a$. Let us write $a=4^{k}a_{1}$ where $4\nmid a_{1}$. By the first part of the proof, there exists a tuple $(x_{1},y_{1},z_{1},w_{1})\in\mathbb{Z}_{p}^{4}$ such that
-$x_{1}^{2}+y_{1}^{2}+Dz_{1}^{2}+Dw_{1}^{2}=a_{1}$
-and we set $(x,y,z,w)=(2^{k}x_{1},2^{k}y_{1},2^{k}z_{1},2^{k}w_{1})$. ∎
+We end with a short section providing a shortened proof of the universality of $G(z,w)=z\tau(z)+w\tau(w)$ over real quadratic fields, where the form is indefinite.
+For a more detailed approach, please consult Section A5 of the Appendix.
 Theorem 5.1.
@@
 $Q(x_{1},y_{1},x_{2},y_{2})=(x_{1}+y_{1}\sqrt{D})(x_{1}-y_{1}\sqrt{D})+(x_{2}+y_{2}\sqrt{D})(x_{2}-y_{2}\sqrt{D})=x_{1}^{2}+x_{2}^{2}-Dy_{1}^{2}-Dy_{2}^{2}.$
-The form $Q$ represents $a$ over $\mathbb{Z}_{p}$ for $p$ odd by Lemma 6.4 and for $p=2$ by Lemma 6.5. By Theorem 6.1, it represents $a$ over $\mathbb{Z}$. ∎
-Proposition 6.7.
-Let $D\in\mathbb{Z}_{\geq 1}$. The quadratic form
-$Q(x,y,z,w)=x^{2}+y^{2}-Dz^{2}-Dw^{2}$
-represents 00 non-trivially if and only if $D$ is a sum of two squares.
-Proof.
-Assume that $D$ is a sum of two squares, say $D=a^{2}+b^{2}$ where $a,b\in\mathbb{Z}$. If $(x,y,z,w)=(a,b,1,0)$, then $Q(x,y,z,w)=0$.
-Now we prove the converse under the assumption that $D$ is squarefree. Let $p\mid D$ be an odd prime and $(x,y,z,w)$ be a primitive solution to
-$x^{2}+y^{2}-Dz^{2}-Dw^{2}=0.$
-In particular, not all of $x$, $y$, $z$, $w$ are divisible by $p$.
-First, suppose that $p\nmid x$ or $p\nmid y$. Without loss of generality, we can assume $p\nmid y$. Reducing the equation modulo $p$, we get $x^{2}+y^{2}\equiv 0\pmod{p}$, hence
-$\left(\frac{x}{y}\right)^{2}\equiv-1\pmod{p}.$
-This shows that $-1$ is a square modulo $p$, hence $p\equiv 1\pmod{4}$.
-Secondly, suppose that $p\mid x$ and $p\mid y$.
-From the equation, we have $p^{2}\mid D(z^{2}+w^{2})$, and because $D$ is squarefree, $p\mid z^{2}+w^{2}$.
-Since $(x,y,z,w)$ is primitive, we get $p\nmid z$ or $p\nmid w$, and the congruence $z^{2}+w^{2}\equiv 0\pmod{p}$ implies $p\equiv 1\pmod{4}$.
-We proved that every odd prime $p\mid D$ satisfies $p\equiv 1\pmod{4}$, hence $D$ is a sum of two squares.
-To prove the converse for general $D$, write $D=k^{2}d$ where $k,d\in\mathbb{Z}$ and $d$ is squarefree. If $(x_{0},y_{0},z_{0},w_{0})$ is a solution to
-$x^{2}+y^{2}-k^{2}dz^{2}-k^{2}dw^{2}=0,$
-then $(x_{0},y_{0},kz_{0},kw_{0})$ is a solution to
-$x^{2}+y^{2}-dz^{2}-dw^{2}=0.$
-Thus, if $Q$ represents 00, then $x^{2}+y^{2}-dz^{2}-dw^{2}$ also represents 00. By the first part of the proof, $d$ is a sum of two squares, hence $D$ is also a sum of two squares. ∎
-7. Examples of universal binary generalized forms
-7.1. Preliminaries
-Let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}_{\geq 2}$ is squarefree, $G$ be a $\mathbb{Z}$-valued generalized form over $K$ in $n$ variables, and $Q$ be the quadratic form in $2n$ variables associated to $G$.
-By Lemma 4.5, if one of the following conditions is satisfied:
-  1. i)
-    $D\equiv 2,3\pmod{4}$ and $G$ is integral,
-  2. ii)
-    $D\equiv 1\pmod{4}$ and $G$ is classical,
-then $Q$ is classical. By Theorem 1.2, a $\mathbb{Z}$-universal binary generalized form satisfying the conditions above exists over $K$ if and only if $D\in\{2,3,5,6,7,10\}$.
-A complete list of universal classical positive definite quaternary quadratic forms (up to equivalence) is available [Bh, Table 5].
-In this section, we answer the following question: Which of these quadratic forms are equivalent to a quadratic form $Q$ associated to $G$ satisfying the conditions above?
-For each $D\in\{2,3,5,6,7,10\}$ we identify all such forms.
-We use Bhargava’s notation $D:\ a\ b\ c\ d\ e\ f$ for a ternary quadratic form with matrix
-$M_{L}=\begin{pmatrix}a&\frac{f}{2}&\frac{e}{2}\\ \frac{f}{2}&b&\frac{d}{2}\\ \frac{e}{2}&\frac{d}{2}&c\end{pmatrix}$
-of determinant $D$. If $L=D:\ a\ b\ c\ d\ e\ f$, then $1\oplus L$ denotes the quaternary form with matrix
-$M_{1\oplus L}=\begin{pmatrix}1&0&0&0\\ 0&a&\frac{f}{2}&\frac{e}{2}\\ 0&\frac{f}{2}&b&\frac{d}{2}\\ 0&\frac{e}{2}&\frac{d}{2}&c\end{pmatrix}.$
-We also need a concise way of representing a generalized binary form. Thus we let $[a,b,c,d,e,f]$ represent the binary generalized form with matrix
-$\begin{pmatrix}a&\frac{b}{2}&\frac{d}{2}&\frac{e}{2}\\ \frac{b}{2}&c&\frac{\tau(e)}{2}&\frac{f}{2}\\ \frac{d}{2}&\frac{\tau(e)}{2}&\tau(a)&\frac{\tau(b)}{2}\\ \frac{e}{2}&\frac{f}{2}&\frac{\tau(b)}{2}&\tau(c)\end{pmatrix}$
-where $a,b,c,e\in\mathcal{O}_{K}$ and $d,f\in\mathbb{Z}$. We also use $[a_{11},a_{12},a_{13},a_{14},a_{22},a_{23},a_{24},a_{33},a_{34},a_{44}]$ to represent the quaternary quadratic form with matrix
-$\begin{pmatrix}a_{11}&\frac{a_{12}}{2}&\frac{a_{13}}{2}&\frac{a_{14}}{2}\\ \frac{a_{12}}{2}&a_{22}&\frac{a_{23}}{2}&\frac{a_{24}}{2}\\ \frac{a_{13}}{2}&\frac{a_{23}}{2}&a_{33}&\frac{a_{34}}{2}\\ \frac{a_{14}}{2}&\frac{a_{24}}{2}&\frac{a_{34}}{2}&a_{44}\end{pmatrix}.$
-7.2. Universal quadratic forms associated to binary generalized forms for $D=2$
-Let $G$ be a $\mathbb{Z}$-universal binary generalized form over $\mathbb{Q}(\sqrt{2})$, $Q$ be the associated quaternary quadratic form, and $M_{Q}$ be the matrix of $Q$.
-We have $4\mid\det(M_{Q})$ by Lemma 4.2 and $\frac{\det(M_{Q})}{4}\equiv 0,1\pmod{4}$ by Proposition 5.14.
-From the table of universal classical quaternary forms [Bh, Table 5], we select all entries satisfying these two conditions.
-We also know that $\operatorname{rank}(M_{Q}\bmod 2)\leq 2$ by Proposition 4.7, hence we compute the rank of each of them.
-The results are in Table 1.
-We see that we are left with $23$ possible candidates.
-$L$ $\operatorname{rank}(M_{1\oplus L}\bmod{2})$
-4: 1 1 4 0 0 0:41140004:\ 1\ 1\ 4\ 0\ 0\ 04 : 1 1 4 0 0 0 $3$
-4: 1 2 2 0 0 0:41220004:\ 1\ 2\ 2\ 0\ 0\ 04 : 1 2 2 0 0 0 $2$
-4: 2 2 2 2 2 0:42222204:\ 2\ 2\ 2\ 2\ 2\ 04 : 2 2 2 2 2 0 $3$
-16: 1 2 8 0 0 0:1612800016:\ 1\ 2\ 8\ 0\ 0\ 016 : 1 2 8 0 0 0 $2$
-16: 2 2 4 0 0 0:1622400016:\ 2\ 2\ 4\ 0\ 0\ 016 : 2 2 4 0 0 0 $1$
-16: 2 3 3 2 0 0:1623320016:\ 2\ 3\ 3\ 2\ 0\ 016 : 2 3 3 2 0 0 $2$
-20: 1 2 10 0 0 0:20121000020:\ 1\ 2\ 10\ 0\ 0\ 020 : 1 2 10 0 0 0 $2$
-20: 2 2 5 0 0 0:$\ 2$\ 5\ 0\ 0\ 020 : 2 2 5 0 0 0 $2$
-20: 2 2 6 2 2 0:$\ 2$\ 6\ 2\ 2\ 020 : 2 2 6 2 2 0 $3$
-20: 2 3 4 0 0 2:2023400220:\ 2\ 3\ 4\ 0\ 0\ 220 : 2 3 4 0 0 2 $3$
-20: 2 4 4 4 2 0:2024442020:\ 2\ 4\ 4\ 4\ 2\ 020 : 2 4 4 4 2 0 $3$
-32: 2 4 4 0 0 0:3224400032:\ 2\ 4\ 4\ 0\ 0\ 032 : 2 4 4 0 0 0 $1$
-32: 2 4 5 4 0 0:3224540032:\ 2\ 4\ 5\ 4\ 0\ 032 : 2 4 5 4 0 0 $2$
-36: 2 3 6 0 0 0:3623600036:\ 2\ 3\ 6\ 0\ 0\ 036 : 2 3 6 0 0 0 $2$
-36: 2 4 5 0 2 0:3624502036:\ 2\ 4\ 5\ 0\ 2\ 036 : 2 4 5 0 2 0 $3$
-36: 2 4 6 4 2 0:3624642036:\ 2\ 4\ 6\ 4\ 2\ 036 : 2 4 6 4 2 0 $3$
-36: 2 5 5 4 2 2:3625542236:\ 2\ 5\ 5\ 4\ 2\ 236 : 2 5 5 4 2 2 $3$
-48: 2 3 8 0 0 0:4823800048:\ 2\ 3\ 8\ 0\ 0\ 048 : 2 3 8 0 0 0 $2$
-48: 2 4 6 0 0 0:4824600048:\ 2\ 4\ 6\ 0\ 0\ 048 : 2 4 6 0 0 0 $1$
-48: 2 5 5 2 0 0:4825520048:\ 2\ 5\ 5\ 2\ 0\ 048 : 2 5 5 2 0 0 $2$
-52: 2 3 9 2 0 0:5223920052:\ 2\ 3\ 9\ 2\ 0\ 052 : 2 3 9 2 0 0 $2$
-52: 2 5 6 2 0 2:5225620252:\ 2\ 5\ 6\ 2\ 0\ 252 : 2 5 6 2 0 2 $3$
-52: 2 5 6 4 0 0:5225640052:\ 2\ 5\ 6\ 4\ 0\ 052 : 2 5 6 4 0 0 $2$
-64: 2 4 8 0 0 0:6424800064:\ 2\ 4\ 8\ 0\ 0\ 064 : 2 4 8 0 0 0 $1$
-68: 2 4 9 0 2 0:6824902068:\ 2\ 4\ 9\ 0\ 2\ 068 : 2 4 9 0 2 0 $3$
-68: 2 4 10 4 2 0:68241042068:\ 2\ 4\ 10\ 4\ 2\ 068 : 2 4 10 4 2 0 $3$
-68: 2 5 7 2 0 0:6825720068:\ 2\ 5\ 7\ 2\ 0\ 068 : 2 5 7 2 0 0 $2$
-80: 2 4 10 0 0 0:80241000080:\ 2\ 4\ 10\ 0\ 0\ 080 : 2 4 10 0 0 0 $1$
-80: 2 4 11 4 0 0:80241140080:\ 2\ 4\ 11\ 4\ 0\ 080 : 2 4 11 4 0 0 $2$
-80: 2 5 8 0 0 0:8025800080:\ 2\ 5\ 8\ 0\ 0\ 080 : 2 5 8 0 0 0 $2$
-96: 2 4 12 0 0 0:96241200096:\ 2\ 4\ 12\ 0\ 0\ 096 : 2 4 12 0 0 0 $1$
-96: 2 4 13 4 0 0:96241340096:\ 2\ 4\ 13\ 4\ 0\ 096 : 2 4 13 4 0 0 $2$
-100: 2 4 13 0 2 0:1002413020100:\ 2\ 4\ 13\ 0\ 2\ 0100 : 2 4 13 0 2 0 $3$
-100: 2 4 14 4 2 0:1002414420100:\ 2\ 4\ 14\ 4\ 2\ 0100 : 2 4 14 4 2 0 $3$
-100: 2 5 10 0 0 0:1002510000100:\ 2\ 5\ 10\ 0\ 0\ 0100 : 2 5 10 0 0 0 $2$
-112: 2 4 14 0 0 0:1122414000112:\ 2\ 4\ 14\ 0\ 0\ 0112 : 2 4 14 0 0 0 $1$
-Table 1. Ternary forms $L$ such that $1\oplus L$ is universal, $\det(M_{1\oplus L})$ is divisible by $4$, and $\frac{\det(M_{1\oplus L})}{4}\equiv 0,1\pmod{4}$.
-To determine which of these candidates are equivalent to a quadratic form associated to some $G$, we use the $2$-adic normal form.
-Let $\operatorname{NF}(M)$ denote the normal form of a symmetric matrix $M$ over $\mathbb{Z}_{2}$.
-The precise definition is very technical and we do not give it here.
-It is sufficient for us to be able to compute $\operatorname{NF}(M)$, for which we use SageMath.
-Details on the implementation of the $p$-adic normal form can be found in the reference manual [Sa].
-The important property of the normal form is uniqueness: Let $Q_{1}$ and $Q_{2}$ be two quadratic forms in $n$ variables over $\mathbb{Z}_{2}$ with matrices $M_{Q_{1}}$ and $M_{Q_{2}}$, respectively.
-Then $Q_{1}$ and $Q_{2}$ are equivalent over $\mathbb{Z}_{2}$ if and only if $\operatorname{NF}(M_{Q_{1}})=\operatorname{NF}(M_{Q_{2}})$.
-Lemma 7.1.
-Let $K=\mathbb{Q}(\sqrt{2})$ and let $G$ be an integral binary generalized form over $K$ such that $G(\mathbf{a})\in\mathbb{Q}$ for every $\mathbf{a}\in K^{2}$.
-Let $Q$ be the quaternary quadratic form associated to $G$ and $M_{Q}$ be the matrix of $Q$.
-We have $\operatorname{NF}(M_{Q})\bmod{8}=\operatorname{NF}(M)\bmod{8}$ for some matrix $M$ such that
-(6) $M=\begin{pmatrix}A_{1}&4a_{2}&B_{1}&2E_{1}\\ 4a_{2}&2A_{2}&2E_{2}&2B_{2}\\ B_{1}&2E_{2}&C_{1}&4c_{2}\\ 2E_{1}&2B_{2}&4c_{2}&2C_{2}\end{pmatrix}$
-where $A_{1},B_{1},C_{1}\in\{0,\dots,7\}$, $a_{2},c_{2}\in\{0,1\}$, $A_{2},C_{2},B_{2},E_{1},E_{2}\in\{0,\dots,3\}$, $A_{1}+A_{2}\equiv 0\pmod{4}$, $C_{1}+C_{2}\equiv 0\pmod{4}$, $B_{1}+B_{2}\equiv 0\pmod{2}$, and $E_{1}+E_{2}\equiv 0\pmod{2}$.
-Proof.
-By Lemma 5.3 with $D=2$, the matrix of $Q$ equals
-$M_{Q}=\begin{pmatrix}2a_{1}+d&4a_{2}&b_{1}+e_{1}&2(b_{2}-e_{2})\\ 4a_{2}&2(2a_{1}-d)&2(b_{2}+e_{2})&2(b_{1}-e_{1})\\ b_{1}+e_{1}&2(b_{2}+e_{2})&2c_{1}+f&4c_{2}\\ 2(b_{2}-e_{2})&2(b_{1}-e_{1})&4c_{2}&2(2c_{1}-f)\end{pmatrix}.$
-Let $A_{1}=2a_{1}+d$, $A_{2}=2a_{1}-d$, $C_{1}=2c_{1}+f$, $C_{2}=2c_{1}-f$, $B_{1}=b_{1}+e_{1}$, $B_{2}=b_{1}-e_{1}$, $E_{1}=b_{2}-e_{2}$, and $E_{2}=b_{2}+e_{2}$.
-These parameters clearly satisfy the stated congruences.
-Since $\operatorname{NF}(M_{Q})\bmod{8}$ depends only on $M_{Q}\bmod{8}$, we can reduce $M_{Q}$ modulo $8$ to obtain one of the matrices (6).
-∎
-We compute $\operatorname{NF}(M_{1\oplus L})\bmod{8}$ for each of the forms in Table 1 satisfying $\operatorname{rank}(1\oplus L\bmod{2})\leq 2$ and check whether it equals $\operatorname{NF}(M)\bmod{8}$ for one of the finitely many matrices $M$ in Lemma 7.1.
-If not, then we know that $1\oplus L$ is not equivalent to a quadratic form associated to some $G$ and we can discard it.
-The SageMath code of this computation is available at: https://github.com/zinmik/Generalized-forms.
-We are left with $15$ quadratic forms $1\oplus L$ and for each one, we directly find $G$ such that the quadratic form associated to $G$ is equivalent to $1\oplus L$, see Table 2.
-$L$ $G$ $Q$
-4: 1 2 2 0 0 0:41220004:\ 1\ 2\ 2\ 0\ 0\ 04 : 1 2 2 0 0 0 $[1,-\sqrt{2},1,-1,0,-1]$ $[1,0,0,-4,6,-4,0,1,0,6]$
-16: 2 2 4 0 0 0:1622400016:\ 2\ 2\ 4\ 0\ 0\ 016 : 2 2 4 0 0 0 $[1,-1,1,-1,1,0]$ $[1,0,0,0,6,0,-8,2,0,4]$
-20: 1 2 10 0 0 0:20121000020:\ 1\ 2\ 10\ 0\ 0\ 020 : 1 2 10 0 0 0 $[1,-\sqrt{2},2,-1,-\sqrt{2},-1]$ $[1,0,0,0,6,-8,0,3,0,10]$
-20: 2 2 5 0 0 0:$\ 2$\ 5\ 0\ 0\ 020 : 2 2 5 0 0 0 $[1,-\sqrt{2},2,1,\sqrt{2},1]$ $[3,0,0,-8,2,0,0,5,0,6]$
-32: 2 4 4 0 0 0:3224400032:\ 2\ 4\ 4\ 0\ 0\ 032 : 2 4 4 0 0 0 $[1,-1-\sqrt{2},2,-1,1,-2]$ $[1,0,0,-4,6,-4,-8,2,0,12]$
-36: 2 3 6 0 0 0:3623600036:\ 2\ 3\ 6\ 0\ 0\ 036 : 2 3 6 0 0 0 $[1,0,1,-1,0,1]$ $[1,0,0,0,6,0,0,3,0,2]$
-48: 2 4 6 0 0 0:4824600048:\ 2\ 4\ 6\ 0\ 0\ 048 : 2 4 6 0 0 0 $[1,0,1,-1,0,0]$ $[1,0,0,0,6,0,0,2,0,4]$
-52: 2 3 9 2 0 0:5223920052:\ 2\ 3\ 9\ 2\ 0\ 052 : 2 3 9 2 0 0 $[1,-\sqrt{2},2,1,0,-1]$ $[3,0,0,-4,2,-4,0,3,0,10]$
-52: 2 5 6 4 0 0:5225640052:\ 2\ 5\ 6\ 4\ 0\ 052 : 2 5 6 4 0 0 $[1,-\sqrt{2},2,-1,0,1]$ $[1,0,0,-4,6,-4,0,5,0,6]$
-64: 2 4 8 0 0 0:6424800064:\ 2\ 4\ 8\ 0\ 0\ 064 : 2 4 8 0 0 0 $[1,-\sqrt{2},2,-1,0,-2]$ $[1,0,0,-4,6,-4,0,2,0,12]$
-68: 2 5 7 2 0 0:6825720068:\ 2\ 5\ 7\ 2\ 0\ 068 : 2 5 7 2 0 0 $[2,-2\sqrt{2},2,-1,0,1]$ $[3,0,0,-8,10,-8,0,5,0,6]$
-80: 2 4 10 0 0 0:80241000080:\ 2\ 4\ 10\ 0\ 0\ 080 : 2 4 10 0 0 0 $[1,-1,2,0,-1,-1]$ $[2,0,-4,0,4,0,0,3,0,10]$
-96: 2 4 12 0 0 0:96241200096:\ 2\ 4\ 12\ 0\ 0\ 096 : 2 4 12 0 0 0 $[1,-1+\sqrt{2},3-\sqrt{2},-1,1+\sqrt{2},-2]$ $[1,0,0,0,6,8,-8,4,-8,16]$
-100: 2 5 10 0 0 0:1002510000100:\ 2\ 5\ 10\ 0\ 0\ 0100 : 2 5 10 0 0 0 $[2,-\sqrt{2},2,-1,\sqrt{2},1]$ $[3,0,0,-8,10,0,0,5,0,6]$
-112: 2 4 14 0 0 0:1122414000112:\ 2\ 4\ 14\ 0\ 0\ 0112 : 2 4 14 0 0 0 $[2,0,1,-3,0,0]$ $[1,0,0,0,14,0,0,2,0,4]$
-Table 2. Quadratic forms associated to $\mathbb{Z}$-universal binary generalized forms over $\mathbb{Q}(\sqrt{2})$.
-We proved the following proposition.
-Proposition 7.2.
-Let $K=\mathbb{Q}(\sqrt{2})$ and let $G$ be a $\mathbb{Z}$-universal binary generalized form over $K$.
-If $Q$ is the quadratic form associated to $G$, then $Q$ is equivalent to $1\oplus L$ where $L$ is one of the $15$ ternary forms in Table 2.
-Conversely, for each of these forms, there exists an equivalent quadratic form associated to a $\mathbb{Z}$-universal binary form over $K$.
-7.3. Universal quadratic forms associated to binary generalized forms for $D=3$
-Lemma 7.3.
-Let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}\setminus\{0,1\}$ is squarefree.
-Let $G$ be an integral binary generalized form over $K$ and $Q$ be the quadratic form associated to $G$ with matrix $M_{Q}$.
-Assume that $G(\mathbf{a})\in\mathbb{Q}$ for every $\mathbf{a}\in K^{2}$.
-If $D\equiv 3\pmod{4}$, then $\operatorname{rank}(M_{Q}\bmod{2})\in\{0,2,4\}$.
-Proof.
-Let the matrix of $G$ be given by (5). By Lemma 5.3,
-$M_{Q}=\begin{pmatrix}2a_{1}+d&2Da_{2}&b_{1}+e_{1}&D(b_{2}-e_{2})\\ 2Da_{2}&D(2a_{1}-d)&D(b_{2}+e_{2})&D(b_{1}-e_{1})\\ b_{1}+e_{1}&D(b_{2}+e_{2})&2c_{1}+f&2Dc_{2}\\ D(b_{2}-e_{2})&D(b_{1}-e_{1})&2Dc_{2}&D(2c_{1}-f)\end{pmatrix}.$
-Let $B=b_{1}+e_{1}$ and $E=b_{2}+e_{2}$. Since $D\equiv 1\pmod{2}$, we get
-$M_{Q}\bmod{2}=\begin{pmatrix}d&0&B&E\\ 0&d&E&B\\ B&E&f&0\\ E&B&0&f\end{pmatrix}.$
-There are $16$ choices for $(d,B,E,f)$ modulo $2$ and it can be checked that $\operatorname{rank}(M_{Q}\bmod{2})\in\{0,2,4\}$ for each of them. ∎
-Let $Q$ be a classical quaternary positive definite form over $\mathbb{Z}$ with matrix $M_{Q}$ and suppose that there exists a binary integral generalized form $G$ over $K=\mathbb{Q}(\sqrt{3})$ such that $Q$ is associated to $G$.
-We know from Corollary 4.3 that $3^{2}\mid\det(M_{Q})$.
-Moreover, by Proposition 5.14,
-$\frac{\det(M_{Q})}{3^{2}}\equiv 0,1\pmod{4}.$
-By Proposition 4.7, $\operatorname{rank}(M_{Q}\bmod{3})\leq 2$ and by Lemma 7.3, $\operatorname{rank}(M_{Q}\bmod{2})$ is even.
-In Table 3, we list all universal classical quaternary quadratic forms $1\oplus L$ with $\det(M_{1\oplus L})$ divisible by $9$ and such that $\frac{\det(M_{1\oplus L})}{9}\equiv 0,1\pmod{4}$ from [Bh, Table 5].
-For each of them we compute its rank modulo $2$ and $3$.
-$L$ $\operatorname{rank}(M_{1\oplus L}\bmod{2})$ $\operatorname{rank}(M_{1\oplus L}\bmod{3})$
-9: 1 2 5 2 0 0:91252009:\ 1\ 2\ 5\ 2\ 0\ 09 : 1 2 5 2 0 0 $4$ $3$
-9: 1 3 3 0 0 0:91330009:\ 1\ 3\ 3\ 0\ 0\ 09 : 1 3 3 0 0 0 $4$ $2$
-9: 2 2 3 0 0 $\ 2$\ 3\ 0\ 0\ 29 : 2 2 3 0 0 2 $4$ $2$
-36: 2 3 6 0 0 0:3623600036:\ 2\ 3\ 6\ 0\ 0\ 036 : 2 3 6 0 0 0 $2$ $2$
-36: 2 4 5 0 2 0:3624502036:\ 2\ 4\ 5\ 0\ 2\ 036 : 2 4 5 0 2 0 $3$ $3$
-36: 2 4 6 4 2 0:3624642036:\ 2\ 4\ 6\ 4\ 2\ 036 : 2 4 6 4 2 0 $3$ $3$
-36: 2 5 5 4 2 2:3625542236:\ 2\ 5\ 5\ 4\ 2\ 236 : 2 5 5 4 2 2 $3$ $2$
-45: 2 4 7 0 2 2:4524702245:\ 2\ 4\ 7\ 0\ 2\ 245 : 2 4 7 0 2 2 $4$ $3$
-45: 2 5 5 0 2 0:4525502045:\ 2\ 5\ 5\ 0\ 2\ 045 : 2 5 5 0 2 0 $4$ $3$
-45: 2 5 6 4 2 2:4525642245:\ 2\ 5\ 6\ 4\ 2\ 245 : 2 5 6 4 2 2 $4$ $3$
-72: 2 4 9 0 0 0:7224900072:\ 2\ 4\ 9\ 0\ 0\ 072 : 2 4 9 0 0 0 $2$ $3$
-72: 2 4 10 4 0 0:72241040072:\ 2\ 4\ 10\ 4\ 0\ 072 : 2 4 10 4 0 0 $1$ $3$
-72: 2 5 8 4 0 0:7225840072:\ 2\ 5\ 8\ 4\ 0\ 072 : 2 5 8 4 0 0 $2$ $3$
-90: 2 4 12 2 2 0:90241222090:\ 2\ 4\ 12\ 2\ 2\ 090 : 2 4 12 2 2 0 $3$ $3$
-90: 2 5 9 0 0 0:9025900090:\ 2\ 5\ 9\ 0\ 0\ 090 : 2 5 9 0 0 0 $3$ $3$
-108: 2 4 14 0 2 0:1082414020108:\ 2\ 4\ 14\ 0\ 2\ 0108 : 2 4 14 0 2 0 $3$ $3$
-Table 3. Ternary forms $L$ such that $1\oplus L$ is universal, $\det(M_{1\oplus L})$ is divisible by $9$ and $\frac{\det(M_{1\oplus L})}{9}\equiv 0,1\pmod{4}$.
-Proposition 7.4.
-Let $K=\mathbb{Q}(\sqrt{3})$. If $G$ is a $\mathbb{Z}$-universal binary generalized form over $K$ and $Q$ is the quadratic form associated to $G$, then $Q$ is equivalent to one of the following forms:
-$\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&3&0\\ 0&0&0&3\end{pmatrix},\ \begin{pmatrix}1&0&0&0\\ 0&2&1&0\\ 0&1&2&0\\ 0&0&0&3\end{pmatrix},\ \begin{pmatrix}1&0&0&0\\ 0&2&0&0\\ 0&0&3&0\\ 0&0&0&6\end{pmatrix}.$
-Conversely, for each of these forms, there exists an equivalent form $Q$ associated to a $\mathbb{Z}$-universal binary generalized form over $K$.
-Proof.
-Let $M_{Q}$ be the matrix of $Q$.
-If $Q$ is associated to a $\mathbb{Z}$-universal binary generalized form, then $Q$ must be equivalent to one of the forms in Table 3 by the discussion before the proposition.
-Moreover, $\operatorname{rank}(M_{Q}\bmod{3})\leq 2$ and $\operatorname{rank}(M_{Q}\bmod{2})$ is even.
-The only forms $L$ in Table 3 such that $\operatorname{rank}(M_{1\oplus L}\bmod{2})$ is even and $\operatorname{rank}(M_{1\oplus L}\bmod{3})\leq 2$ are 9: 1 3 3 0 0 0:91330009:\ 1\ 3\ 3\ 0\ 0\ 09 : 1 3 3 0 0 0, 9: 2 2 3 0 0 $\ 2$\ 3\ 0\ 0\ 29 : 2 2 3 0 0 2, and 36: 2 3 6 0 0 0:3623600036:\ 2\ 3\ 6\ 0\ 0\ 036 : 2 3 6 0 0 0.
-Now we prove the converse. Let $G_{1}$ be the generalized form $[2,-2\sqrt{3},2,-3,-\sqrt{3},0]$ and let $Q_{1}$ be the quadratic form associated to $G_{1}$. By Lemma 5.3, the matrix of $Q_{1}$ equals
-$M_{Q_{1}}=\begin{pmatrix}1&0&0&-3\\ 0&21&-9&0\\ 0&-9&4&0\\ -3&0&0&12\end{pmatrix}.$
-We verified in Lemma 5.5 that $Q_{1}$ is equivalent to 9: 1 3 3 0 0 0:91330009:\ 1\ 3\ 3\ 0\ 0\ 09 : 1 3 3 0 0 0.
-Let $G_{2}$ be the generalized form $[1,-1-\sqrt{3},2,-1,1,-2]$ and let $Q_{2}$ be the quadratic form associated to $G_{2}$. The matrix of $Q_{2}$ equals
-$M_{Q_{2}}=\begin{pmatrix}1&0&0&-3\\ 0&9&-3&-6\\ 0&-3&2&0\\ -3&-6&0&18\end{pmatrix}.$
-If we let
-$A=\begin{pmatrix}1&-3&0&3\\ 0&-1&0&2\\ 0&-1&1&3\\ 0&-1&0&1\end{pmatrix},$
-then $\det(A)=1$, hence $A\in\operatorname{SL}_{4}(\mathbb{Z})$. We have
-$A^{\top}M_{Q_{2}}A=\begin{pmatrix}1&-3&0&3\\ 0&-1&0&2\\ 0&-1&1&3\\ 0&-1&0&1\end{pmatrix}^{\top}\begin{pmatrix}1&0&0&-3\\ 0&9&-3&-6\\ 0&-3&2&0\\ -3&-6&0&18\end{pmatrix}\begin{pmatrix}1&-3&0&3\\ 0&-1&0&2\\ 0&-1&1&3\\ 0&-1&0&1\end{pmatrix}=\begin{pmatrix}1&0&0&0\\ 0&2&1&0\\ 0&1&2&0\\ 0&0&0&3\end{pmatrix},$
-hence $Q_{2}$ is equivalent to 9: 2 2 3 0 0 $\ 2$\ 3\ 0\ 0\ 29 : 2 2 3 0 0 2.
-Let $G_{3}$ be the generalized form $[1,1,1,0,1,1]$ and let $Q_{3}$ be the quadratic form associated to $G_{3}$. The matrix of $Q_{3}$ equals
-$M_{Q_{3}}=\begin{pmatrix}2&0&2&0\\ 0&6&0&0\\ 2&0&3&0\\ 0&0&0&3\end{pmatrix}\sim\begin{pmatrix}1&0&0&0\\ 0&2&0&0\\ 0&0&3&0\\ 0&0&0&6\end{pmatrix},$
-hence $Q_{3}$ is equivalent to 36: 2 3 6 0 0 0:3623600036:\ 2\ 3\ 6\ 0\ 0\ 036 : 2 3 6 0 0 0. ∎
-7.4. Universal quadratic forms associated to classical binary generalized forms for $D=5$
-Let $Q$ be a classical quaternary positive definite form over $\mathbb{Z}$ with matrix $M_{Q}$ and suppose that there exists a binary classical generalized form $G$ over $K=\mathbb{Q}(\sqrt{5})$ such that $Q$ is associated to $G$.
-By Corollary 4.3, $5^{2}\mid\det(M_{Q})$ and by Proposition 4.7, $\operatorname{rank}(M_{Q}\bmod{5})\leq 2$.
-In Table 4, we list all universal quaternary quadratic forms $1\oplus L$ with $\det(M_{1\oplus L})$ divisible by $25$ from [Bh, Table 5]. For each of the we compute its rank modulo $5$.
-$L$ $\operatorname{rank}(M_{1\oplus L}\bmod{5})$
-25: 1 2 13 2 0 0:25121320025:\ 1\ 2\ 13\ 2\ 0\ 025 : 1 2 13 2 0 0 $3$
-25: 2 3 5 0 0 2:2523500225:\ 2\ 3\ 5\ 0\ 0\ 225 : 2 3 5 0 0 2 $2$
-25: 2 3 5 2 2 0:2523522025:\ 2\ 3\ 5\ 2\ 2\ 025 : 2 3 5 2 2 0 $3$
-50: 2 4 7 2 2 0:5024722050:\ 2\ 4\ 7\ 2\ 2\ 050 : 2 4 7 2 2 0 $3$
-100: 2 4 13 0 2 0:1002413020100:\ 2\ 4\ 13\ 0\ 2\ 0100 : 2 4 13 0 2 0 $3$
-100: 2 4 14 4 2 0:1002414420100:\ 2\ 4\ 14\ 4\ 2\ 0100 : 2 4 14 4 2 0 $3$
-100: 2 5 10 0 0 0:1002510000100:\ 2\ 5\ 10\ 0\ 0\ 0100 : 2 5 10 0 0 0 $2$
-Table 4. Ternary forms $L$ such that $1\oplus L$ is universal and $\det(M_{1\oplus L})$ is divisible by $25$.
-Proposition 7.5.
-Let $K=\mathbb{Q}(\sqrt{5})$.
-If $G$ is a classical $\mathbb{Z}$-universal binary generalized form over $K$ and $Q$ is a quadratic form associated to $G$, then $Q$ is equivalent to one of the following forms:
-$\begin{pmatrix}1&0&0&0\\ 0&2&1&0\\ 0&1&3&0\\ 0&0&0&5\end{pmatrix},\ \begin{pmatrix}1&0&0&0\\ 0&2&0&0\\ 0&0&5&0\\ 0&0&0&10\end{pmatrix}.$
-Conversely, for each of these forms, there exists an equivalent form $Q$ associated to a classical $\mathbb{Z}$-universal binary generalized form over $K$.
-Proof.
-Let $M_{Q}$ be the matrix of $Q$.
-If $Q$ is associated to a classical $\mathbb{Z}$-universal binary form, then $Q$ must be one of the forms listed in Table 4.
-Moreover, $\operatorname{rank}(M_{Q}\bmod{5})\leq 2$.
-The only forms $L$ in Table 4 such that $\operatorname{rank}(M_{1\oplus L}\bmod{5})\leq 2$ are 25: 2 3 5 0 0 2:2523500225:\ 2\ 3\ 5\ 0\ 0\ 225 : 2 3 5 0 0 2 and 100: 2 5 10 0 0 0:1002510000100:\ 2\ 5\ 10\ 0\ 0\ 0100 : 2 5 10 0 0 0.
-Let $G_{1}=[3-\omega_{5},0,1,-4,0,0]$ and $G_{2}=[2+\omega_{5},0,6-2\omega_{5},-4,0,-8]$ where $\omega_{5}=\frac{1+\sqrt{5}}{2}$ and let $Q_{1}$, $Q_{2}$ be the quadratic forms associated to $G_{1}$ and $G_{2}$, respectively.
-By Lemma 5.3, the matrices of $Q_{1}$ and $Q_{2}$ are
-$M_{Q_{1}}$ $=\begin{pmatrix}1&-2&0&0\\ -2&9&0&0\\ 0&0&2&1\\ 0&0&1&3\end{pmatrix}\sim\begin{pmatrix}1&0&0&0\\ 0&2&1&0\\ 0&1&3&0\\ 0&0&0&5\end{pmatrix},$
-$M_{Q_{2}}$ $=\begin{pmatrix}1&3&0&0\\ 3&14&0&0\\ 0&0&2&-4\\ 0&0&-4&18\end{pmatrix}\sim\begin{pmatrix}1&0&0&0\\ 0&2&0&0\\ 0&0&5&0\\ 0&0&0&10\end{pmatrix}.$
-(We showed this for $G_{1}$ in Lemma 5.6.) ∎
-7.5. Universal quadratic forms associated to binary generalized forms for $D=6,7,10$
-Proposition 7.6.
-Let $K=\mathbb{Q}(\sqrt{6})$. If $G$ is a $\mathbb{Z}$-universal binary generalized form over $K$ and $Q$ is the quadratic form associated to $G$, then $Q$ is equivalent to the form
-$\begin{pmatrix}1&0&0&0\\ 0&2&0&0\\ 0&0&3&0\\ 0&0&0&6\end{pmatrix}.$
-Conversely, there exists an equivalent form $Q$ associated to a $\mathbb{Z}$-universal binary generalized form over $K$.
-Proof.
-Let $M_{Q}$ be the matrix of $Q$. By Corollary 4.3, $6^{2}\mid\det(M_{Q})$ and by Proposition 5.14,
-$\frac{\det(M_{Q})}{6^{2}}\equiv 0,1\pmod{4}.$
-Moreover, by Proposition 4.7, $\operatorname{rank}(M_{Q}\bmod{2})\leq 2$ and $\operatorname{rank}(M_{Q}\bmod{3})\leq 2$.
-All universal classical quaternary quadratic forms with determinant divisible by $36$ were already listed in Table 3.
-The only form satisfying the additional conditions is 36: 2 3 6 0 0 0:3623600036:\ 2\ 3\ 6\ 0\ 0\ 036 : 2 3 6 0 0 0.
-If we let $G=[3,2\sqrt{6},3,-5,\sqrt{6},-1]$, then
-$M_{Q}=\begin{pmatrix}1&0&0&6\\ 0&66&18&0\\ 0&18&5&0\\ 6&0&0&42\end{pmatrix}\sim\begin{pmatrix}1&0&0&0\\ 0&2&0&0\\ 0&0&3&0\\ 0&0&0&6\end{pmatrix}.$
-(We showed this in Lemma 5.7.) ∎
-Proposition 7.7.
-Let $K=\mathbb{Q}(\sqrt{7})$. If $G$ is a $\mathbb{Z}$-universal binary generalized form over $K$ and $Q$ is the quadratic form associated to $G$, then $Q$ is equivalent to the form
-$\begin{pmatrix}1&0&0&0\\ 0&2&1&0\\ 0&1&4&0\\ 0&0&0&7\end{pmatrix}.$
-Conversely, there exists an equivalent form $Q$ associated to a $\mathbb{Z}$-universal binary generalized form over $K$.
-Proof.
-Let $M_{Q}$ be the matrix of $Q$. By Corollary 4.3, $7^{2}\mid\det(M_{Q})$ and by Proposition 5.14,
-$\frac{\det(M_{Q})}{7^{2}}\equiv 0,1\pmod{4}.$
-The only forms in [Bh, Table 5] satisfying these conditions are those with determinant equal to $49$, namely
-$L_{1}$ =49: 2 3 9 2 2 0,:absent49239220=49:\ 2\ 3\ 9\ 2\ 2\ 0,= 49 : 2 3 9 2 2 0 ,
-$L_{2}$ =49: 2 4 7 0 0 2,:absent49247002=49:\ 2\ 4\ 7\ 0\ 0\ 2,= 49 : 2 4 7 0 0 2 ,
-$L_{3}$ =49: 2 5 6 0 2 2.:absent49256022=49:\ 2\ 5\ 6\ 0\ 2\ 2.= 49 : 2 5 6 0 2 2 .
-By Proposition 4.7, $\operatorname{rank}(M_{Q}\bmod{7})\leq 2$.
-We have $\operatorname{rank}(1\oplus L_{1}\bmod{7})=3$, $\operatorname{rank}(1\oplus L_{2}\bmod{7})=2$, and $\operatorname{rank}(1\oplus L_{3}\bmod{7})=3$.
-Thus, the only candidate is 49: 2 4 7 0 0 2:4924700249:\ 2\ 4\ 7\ 0\ 0\ 249 : 2 4 7 0 0 2.
-Let $G=[3,1+2\sqrt{7},4,-5,-1+2\sqrt{7},3]$. The matrix $M_{Q}$ of the quadratic form $Q$ associated to $G$ equals
-$M_{Q}=\begin{pmatrix}1&0&0&0\\ 0&77&28&14\\ 0&28&11&0\\ 0&14&0&35\end{pmatrix}\sim\begin{pmatrix}1&0&0&0\\ 0&2&1&0\\ 0&1&4&0\\ 0&0&0&7\end{pmatrix}.$
-(We showed this in Lemma 5.8.) ∎
-Proposition 7.8.
-Let $K=\mathbb{Q}(\sqrt{10})$. If $G$ is a $\mathbb{Z}$-universal binary generalized form over $K$ and $Q$ is the quadratic form associated to $G$, then $Q$ is equivalent to the form
-$\begin{pmatrix}1&0&0&0\\ 0&2&0&0\\ 0&0&5&0\\ 0&0&0&10\end{pmatrix}.$
-Conversely, there exists an equivalent form $Q$ associated to a $\mathbb{Z}$-universal binary generalized form over $K$.
-Proof.
-Let $M_{Q}$ be the matrix of $Q$.
-By Corollary 4.3, $10^{2}\mid\det(M_{Q})$.
-The quadratic forms from [Bh, Table 5] with determinant divisible by $100$ were already listed in Table 4.
-By Proposition 4.7, $\operatorname{rank}(M_{Q}\bmod{5})\leq 2$.
-The only form satisfying these criteria is 100: 2 5 10 0 0 0:1002510000100:\ 2\ 5\ 10\ 0\ 0\ 0100 : 2 5 10 0 0 0.
-If we let $G=[3,\sqrt{10},5,-5,\sqrt{10},7]$, then
-$M_{Q}=\begin{pmatrix}1&0&0&0\\ 0&110&20&0\\ 0&20&17&20\\ 0&0&20&30\end{pmatrix}\sim\begin{pmatrix}1&0&0&0\\ 0&2&0&0\\ 0&0&5&0\\ 0&0&0&10\end{pmatrix}.$
-(We showed this in Lemma 5.9.) ∎
+By the local-global principle for representations over $\mathbb{Z}$ by indefinite forms (see, e.g., [Ca, Chap.
+9, Thm.
+1.5]), it suffices to verify that each $a\in\mathbb{Z}$ is represented over $\mathbb{Z}_{p}$ for all primes pp.
+This is indeed straightforward to verify; for details see Section A5 of the Appendix.∎
+It is natural to be interested in the ways a quadratic form represents 0 non-trivially.
+In Proposition A5.6, we show that the quadratic form $Q(x,y,z,w)=x^{2}+y^{2}-Dz^{2}-Dw^{2}$ represents 0 non-trivially if and only if $D$ is a sum of two squares.
+The proof of this proposition can be found in the appendix.